### v7

# Probability Distribution Learning: A theoretical framework for Deep Learning

Abstract
This paper introduces probability distribution learning (PD learning), a novel theoretical learning framework. Departing from the traditional statistical learning framework, PD learning focuses on learning the underlying probability distribution, which is modeled as a random variable within the probability simplex. Within this framework, the learning error is decomposed into uncertainty, estimation error, and the model’s fitting error. Subsequently, we present the methodology for calculating uncertainty, along with optimization strategies for both estimation error and fitting error. Given that minimizing the fitting error typically constitutes a non-convex optimization problem, we introduce a standard loss function and the gradient structural control (GSC) algorithm, and demonstrate that by employing this function, the optima of fitting error minimization can be approached by reducing the gradient norm and structural error. Furthermore, we apply the PD learning framework to deep learning, elucidating the mechanisms by which techniques such as random parameter initialization, over-parameterization, bias-variance trade-off, and dropout influence deep model training. Finally, experimental results on various models validate the effectiveness of the proposed framework.
Keywords: probability distribution, learning theory, deep learning, loss function, non-convex optimization
1 Introduction
Along with the remarkable practical achievements of deep learning, several fundamental questions on deep learning remain unanswered within the classical learning theory framework. These include:
  • Why can deep neural networks (DNNs) avoid overfitting despite their immense capacity, and how do they surpass shallow architectures?
  • How do DNNs overcome the curse of dimensionality, and why do optimization algorithms often succeed in finding satisfactory solutions despite non-convex non-linear, and often non-smooth objectives?
  • Which aspects of an architecture affect the performance of the associated models and the mechanisms behind these influences?
Delving into the reasons and factors that govern the performance of DNNs not only enhances the interpretability of DNNs but also facilitates the development of more principled and reliable architecture designs.
Extensive empirical evidence indicates that, in addition to expressivity (Leshno et al., 1993; Barron, 1993), the generalization ability of DNNs significantly influences prediction performance (Kawaguchi et al., 2017). Generalization ability, which reflects a model’s predictive capacity on unseen data, is a crucial aspect of learning models and is currently at the core of theoretical research in deep learning. Some classical theory work attributes generalization ability to the utilization of a low-complexity class of hypotheses and proposes various complexity measures to control generalization error. These include the Vapnik-Chervonenkis (VC) dimension (Wu, 2021), Rademacher complexity (Bartlett and Mendelson, 2003), and Algorithmic Stability (Bousquet and Elisseeff, 2002; Oneto et al., 2015; Maurer, 2017; Elisseeff et al., 2005; Mukherjee et al., 2006; Shalev-Shwartz et al., 2010). However, recent studies by (Zhang et al., 2016, 2021; He et al., 2015a; Belkin et al., 2018a) demonstrate that high-capacity DNNs can still generalize well with genuine labels, yet generalize poorly when trained with random labels. This apparent contradiction with the traditional understanding of overfitting in complex models highlights the need for further investigation (Vapnik and Alexey, 1971; Wu, 2021; Bartlett and Mendelson, 2003). Apart from inconsistencies with experiments, under the existing theoretical framework, expressivity and generalization ability may not solely determine model performance. A typical example is in highly imbalanced binary classification, such as when the ratio of positive to negative samples is 99:1. Suppose a model predicts all inputs as positive; in this case, the overall accuracy of the model is 99%, and the generalization error is 0, but clearly, the model is ineffective. Within the current statistical machine learning framework, the results learned by the model are not unique and depend on the artificially chosen loss function, rather than being entirely determined by the data. This raises deeper questions: If the learning outcome depends on the artificial selection of the loss function, what is the physical significance of the learning outcomes?
Distribution estimation, a cornerstone in various fields such as machine learning, theoretical and applied statistics, information theory, and communication applications, has a rich history. The distribution, described by a probability density/mass function (PDF/PMF), completely characterizes the ”behavior” of a random variable, making distribution estimation fundamental for investigating the properties of a given dataset and for efficient data mining. The goal of distribution estimation is to estimate the true underlying distribution from a given sample set (Aldrich, 1997; Härdle et al., 2004; McLachlan et al., 2019; Meeker et al., 2022). The majority of machine learning tasks can be framed within the distribution estimation context, with both classification and regression problems mathematically represented by learning the conditional probability distributions between features and labels Ng and Jordan (2001). Generative learning corresponds to estimating the underlying joint distribution of features to generate new data. Consequently, approaching machine learning and deep learning from the perspective of distribution estimation is a logical and coherent choice, given its inherent suitability for these tasks.
In this paper, we introduce a novel framework, probability distribution learning (PD learning), which can be regarded as a distribution estimation based on machine learning models and objective optimization strategies. Departing from the traditional statistical learning framework, PD learning focuses on learning the true underlying probability distribution, which is modeled as a random variable within the probability simplex. To achieve this objective, we introduce the concept of learning risk probability as the optimization objective, which is constrained by uncertainty, estimation error and the model’s fitting error.
  1. 1.
    Uncertainty, which is independent of both the model and optimization algorithm, relies solely on prior knowledge and sampling data, constituting the optimal bound of the learning error. By this concept, we theoretically substantiate the assertion that the quantity and quality of data determines the upper limit of machine learning performance, whereas models and algorithms merely serve to approximate this limit. In this paper, we also provide its computational formula and discuss its properties.
  2. 2.
    Estimation error measures the discrepancy between the optimal estimate of the latent distribution and the actual estimation results. This paper proposes a method for calculating the optimal estimate. We also demonstrate that reducing both the fitting error and estimation error can decrease the upper bound of the learning error, while reducing their discrepancy can lower the lower bound of the learning error. Therefore, we present a new bias-variance trade-off: decreasing the fitting error may increase the learning error; however, reducing both fitting error and estimation error while ensuring their equality is the preferred approach.
  3. 3.
    The fitting error is utilized to measure the disparity between the model’s output and the model’s target fit, and minimizing the model’s fitting error poses a typical non-convex optimization challenge. To address the challenge of non-convexity in optimizing fitting error, we introduce a standard loss function and propose the gradient structural control (GSC) algorithm. The proposed standard loss function is a type of loss function that has a unique optimal solution and sampling stability. We have shown that under these two conditions, it has a unique mathematical expression and can be represented as a Young–Fenchel loss function, which includes various common losses. The GSC algorithm is designed to minimize both the gradient norm and the structural error. We demonstrate that by employing this loss function, global optimal solutions in non-convex optimization can be approximated by the GSC algorithm.
We then apply this framework to deep learning, further demonstrating that PD learning offers a novel perspective on the mechanisms by which techniques such as random parameter initialization, over-parameterization, bias-variance trade-off, and dropout influence deep model training. Lastly, the efficacy of the proposed framework is experimentally substantiated.
The article is organized as follows: In Section 2, we delve into related work, reviewing learning theory, modern deep learning analysis, and distribution estimation methods. In Section 3, we define the fundamental concepts and objects utilized in this paper. In Section 4, we delineate the problem that PD learning is intended to address, and provide an exposition of its algorithmic principles and constituent components. In Section 5, we illustrate how PD learning addresses the questions that remain unanswered within the classical learning theory framework. In Section 6, we validate the central conclusions of our proposed framework via experimentation to affirm the efficacy of the PD learning framework. In Section 7, we summarize our findings and discuss their broader implications for the field. All the proofs are deferred to the Appendix.
2 Related work
In this section, we present an overview of learning theory, modern methods for analyzing deep learning, and distribution learning methods.
2.1 Learning theory
We commence by offering a concise summary of the classical mathematical and statistical theory underlying machine learning tasks and algorithms, which, in their most general form, can be formulated as follows (Berner et al., 2021).
Definition 2.1 (Learning task)
In a learning task, one is provided with a loss function $\ell:\mathcal{M}(\mathcal{X},\mathcal{Y})\times\mathcal{Z}\to\mathbb{R}$ and training data $s^{n}=\{z^{(i)}\}_{i=1}^{n}$ which is generated by independent and identically distributed (i.i.d.) sampling according to the unknown true distribution $Z\sim q$, where $Z$ are random variables that take values in $\mathcal{Z}$. The objective is to construct a learning algorithm $\mathcal{A}$ that utilizes training data $s$ to find a model $f_{s}$ that performs well on the entire data, where $f_{s}=\mathcal{A}(s)\in\mathcal{F}$, the hypothesis set $\mathcal{F}\subset\mathcal{M}(\mathcal{X},\mathcal{Y})$, and the mapping $\mathcal{A}:s^{n}\to\mathcal{F}$. Here, performance is measured via $\mathbb{E}_{Z}[\ell(f_{s},z)]$.
Under the definition of this learning task, traditional statistical learning introduces the empirical risk ^$\widehat{\mathcal{R}}_{s}(f):=\frac{1}{n}\sum_{i=1}^{n}\ell(f,z^{(i)})$ and the expected risk $\mathcal{R}(f):=\mathbb{E}_{Z\sim q}[\ell(f,Z)]$. The Bayes-optimal function $f^{\star}\in\mathcal{M}(\mathcal{X},\mathcal{Y})$, which achieves the minimum Bayes risk, has a Bayes risk of $\mathcal{R}^{\star}:=\inf_{f\in\mathcal{M}(\mathcal{X},\mathcal{Y})}\mathcal{R}(f)$. Considering the impact of the algorithm, let $f_{s}=\mathcal{A}(s)$; then, the expected risk of the model $f_{s}$ is given by $\mathcal{R}(f_{s})=\mathbb{E}_{Z\sim q}[\ell(f_{s},Z)]$. Given a hypothesis set, a widely utilized learning algorithm is empirical risk minimization (ERM), denoted as $\mathcal{A}^{erm}$, which minimizes the average loss on the given training data. The function selected by $\mathcal{A}^{erm}$ we denote as $\widehat{f}_{s}\in\mathcal{F}$. Let $f^{\star}_{\mathcal{F}}\in\arg\min_{f\in\mathcal{F}}\mathcal{R}(f)$ represent the best approximation in $\mathcal{F}$, the optimization error denoted by $e_{\rm opt}$, generalization error denoted by $e_{\rm gen}$ and approximation error denoted by $e_{\rm approx}$ are defined as follows (Berner et al., 2021):
$e_{\rm opt}(\mathcal{A},\mathcal{F}):=\widehat{\mathcal{R}}_{s}(f_{s})-\widehat{\mathcal{R}}_{s}(\widehat{f}_{s})\geq\widehat{\mathcal{R}}_{s}(f_{s})-\widehat{\mathcal{R}}_{s}(f^{\star}_{\mathcal{F}}),$ (1)
$e_{\rm gen}(\mathcal{F}):=\sup_{f\in\mathcal{F}}|\mathcal{R}(f)-\widehat{\mathcal{R}}_{s}(f)|\geq\max\{|\mathcal{R}(f_{s})-\widehat{\mathcal{R}}_{s}(f_{s})|,|\mathcal{R}(f_{\mathcal{F}}^{\star})-\widehat{\mathcal{R}}_{s}(f^{\star}_{\mathcal{F}})|\},$
$e_{\rm approx}(\mathcal{F}):=\mathcal{R}(f^{\star}_{\mathcal{F}})-\mathcal{R}^{\star}.$
The error $\mathcal{R}(f_{s})-\mathcal{R}^{\star}$ can then be bounded by:
$\mathcal{R}(f_{s})-\mathcal{R}^{\star}$ $=\mathcal{R}(f_{s})-\widehat{\mathcal{R}}_{s}(f_{s})+\widehat{\mathcal{R}}_{s}(f_{s})-\widehat{\mathcal{R}}_{s}(f^{\star}_{\mathcal{F}})+\widehat{\mathcal{R}}_{s}(f^{\star}_{\mathcal{F}})-\mathcal{R}_{s}(f^{\star}_{\mathcal{F}})+\mathcal{R}_{s}(f^{\star}_{\mathcal{F}})-\mathcal{R}^{\star}$ (2)
$\leq e_{\rm opt}(\mathcal{A},\mathcal{F})+2e_{\rm gen}(\mathcal{F})+e_{\rm approx}(\mathcal{F}).$
The optimization error $e_{\rm opt}$ is primarily influenced by the algorithm $\mathcal{A}$ that is used to find the model $f_{s}$ in the hypothesis set $\mathcal{F}$ for given training data $s$. By far, the gradient-based optimization methods are the most prevalent method to address this issue, leveraging their ability to precisely and efficiently compute the derivatives by automatic differentiation in a lot of widely used models. Owing to the hierarchical structure of neural networks (NNs), the back-propagation algorithm has been proposed to efficiently calculate the gradient with respect to the loss function (Kelley, 1960; Dreyfus, 1962; Rumelhart et al., 1986; Griewank and Walther, 2000). Stochastic gradient descent (SGD), introduced by Robbins (1951), is a typical gradient-based optimization method used to identify the unique minimum of a convex function (Kiefer and Wolfowitz, 1952). In scenarios where the objective function is non-convex, SGD generally converges to a stationary point, which may be either a local minimum, a local maximum, or a saddle point. Nonetheless, there are indeed conditions under which convergence to local minima is guaranteed (Nemirovski et al., 2008; Ge et al., 2015; Lee et al., 2016; Jentzen et al., 2018).
The universal approximation theorem (Lewicki and Marino, 2003; ichi Funahashi, 1989; Hornik et al., 1989; Leshno et al., 1993) states that neural networks with only 2 layers can approximate any continuous function defined on a compact set up to arbitrary precision. This implies that the approximation error, $e_{\rm approx}$, decreases as the model complexity increases.
The generalization error, $e_{\rm gen}(\mathcal{F})$, is defined as the difference between the true risk and the empirical risk. To bound the generalization error, Hoeffding’s inequality (Hoeffding, 1994) alongside the union bound directly provides the PAC (Probably Approximately Correct) bound (Valiant, 1984). The PAC bound is described as follows:
$\Pr[e_{\rm gen}(\mathcal{F})\leq\varepsilon(\frac{1}{\delta},\frac{1}{n})]\geq 1-\delta.$ (3)
In simpler terms, when the sample $s^{n}$ is chosen randomly, good generalization is ensured with a high probability, at least $1-\delta$, by minimizing $\varepsilon(\frac{1}{\delta},\frac{1}{n})$. The function $\varepsilon$ depends mainly on the number of samples $n$ and the confidence parameter $\delta$. The uniform convergence bound represents a classic form of generalization bound, which relies on metrics of complexity for the set $\mathcal{F}$, including the VC-dimension (Wu, 2021) and the Rademacher complexity (Bartlett and Mendelson, 2003). Such bounds are applicable to all the hypotheses within $\mathcal{F}$ and are expressed mathematically as follows:
$\Pr[e_{\rm gen}(\mathcal{F})\leq\varepsilon(\frac{1}{\delta},\frac{1}{n},\mathcal{F})]\geq 1-\delta.$ (4)
A potential drawback of the uniform convergence bounds is that they are independent of the learning algorithm, i.e., they do not take into account the way the hypothesis space is explored. To tackle this issue, algorithm-dependent bounds have been proposed to take advantage of some particularities of the learning algorithm, such as its uniform stability (Bousquet and Elisseeff, 2002) or robustness (Xu and Mannor, 2010). In such instances, the bounds obtained hold for the hypothesis set $\mathcal{A}(s)\in\mathcal{F}$, which is learned with the algorithm $\mathcal{A}$ using the sample $s$. The form of such bounds is represented as:
$\Pr[e_{\rm gen}(\mathcal{A},\mathcal{F})\leq\varepsilon(\frac{1}{\delta},\frac{1}{n},\mathcal{A})]\geq 1-\delta.$ (5)
For instance, the work presented by Hardt et al. (2015) utilizes this method to derive generalization bounds for hypotheses learned by stochastic gradient descent.
However, to the best of our knowledge, these bounds are insufficient in explicating over-parameterization and non-convex optimization mechanisms in DNNs. The available body of evidence suggests that classical tools from statistical learning theory alone, such as uniform convergence, algorithmic stability, or bias-variance trade-off may be unable to fully capture the behavior of DNNs (Zhang et al., 2016; Nagarajan and Kolter, 2019). The classical trade-off highlights the need to balance model complexity and generalization. While simpler models may suffer from high bias, overly complex ones can overfit noise (high variance). However, reconciling the classical bias-variance trade-off with the empirical observation of good generalization, despite achieving zero empirical risk on noisy data using DNNs with a regression loss, presents a significant challenge (Belkin et al., 2018b). Generally speaking, in deep learning, over-parameterization and perfectly fitting noisy data do not exclude good generalization performance. Moreover, empirical evidence across various modeling techniques, including DNNs, decision trees, random features, and linear models, indicates that test error can decrease even below the sweet-spot in the U-shaped bias-variance curve when further increasing the number of parameters (Belkin et al., 2018a; Geiger et al., 2019; Nakkiran et al., 2019).
Apart from inconsistencies with experiments, under the existing theoretical framework (2), the upper bound of the error $\mathcal{R}(f_{s})-\mathcal{R}^{\star}$, composed of $e_{\rm gen}(\mathcal{F})$, $e_{\rm approx}(\mathcal{F})$ and $e_{\rm opt}(\mathcal{A},\mathcal{F})$, may be too loose. In other words, empirical risk and generalization ability may not completely determine model performance. A typical example is in highly imbalanced binary classification, such as when the ratio of positive to negative samples is 99:1. Suppose a model predicts all inputs as positive; in this case, the overall accuracy of the model is 99%, and the generalization error is 0, but clearly, the model is ineffective. Within the machine learning framework 2.1, the result learned by the model $f_{s}$ is not completely deterministic and unique; it depends on the artificially chosen loss function, rather than being entirely determined by the data. This raises deeper questions: If the learning outcome depends on the artificial selection of the loss function, to what objects in the real physical world does the learned result correspond?
2.2 Deep learning theory
Deep learning poses several mysteries within the conventional learning theory framework, including the outstanding performance of overparameterized neural networks, the impact of depth in deep architectures, the apparent absence of the curse of dimensionality, the surprisingly successful optimization performance despite the non-convexity of the problem, understanding what features are learned, why deep architectures perform exceptionally well in solving physical problems, and which fine aspects of an architecture affect the behavior of a learning task in which way. Theoretical insights into these problems from prior research have predominantly followed two main avenues: expressivity (fitting ability) and generalization ability.
Essentially, expressivity is characterized by the capability of DNNs to approximate any function. The universal approximation theorem Hornik et al. (1989); Cybenko (1989) states that a feedforward network with any ”squashing” activation function $\sigma(\cdot)$, such as the logistic sigmoid function, can approximate any Borel measurable function $f(x)$ with any desired non-zero amount of error, provided that the network is given enough hidden layer size $n$. Universal approximation theorems have also been proved for a wider class of activation functions, which includes the now commonly used ReLU Leshno et al. (1993). Generally speaking, the greater the number of parameters a model possesses, the stronger its fitting capability. Moreover, research has demonstrated that extremely narrow neural networks can retain universal approximativity with an appropriate choice of depth Hanin and Sellke (2017); Kidger and Lyons (2019). More precisely, enhancing the depth of a neural network architecture exponentially increases its expressiveness (Poole et al., 2016; Zhang et al., 2023).
Generalization ability represents a model’s predictive ability on unseen data. Increasing the depth of a neural network architecture naturally yields a highly overparameterized model, whose loss landscape is known for a proliferation of local minima and saddle points (Dauphin et al., 2014). The fact that sometimes (for some specific models and datasets), the generalization ability of a model paradoxically increases with the number of parameters, even surpassing the interpolation threshold—where the model fits the training data perfectly. This observation is highly counterintuitive since this regimen usually corresponds to over-fitting. DNNs show empirically that after the threshold, sometimes the generalization error tends to descend again, showing the so called Double-Descent generalization curve, and reaching a better minimum. In order to address this question, researchers have tried to focus on different methods and studied the phenomena under different names, e.g., over-parametrization (Du et al., 2018b; Yun et al., 2018; Chizat et al., 2018; Arjevani and Field, 2022), double descent Poggio et al. (2019); d’Ascoli et al. (2020); Adlam and Pennington (2020); Liu et al. (2021), sharp/flat minima Hochreiter and Schmidhuber (1997); Dinh et al. (2017); He et al. (2019); Petzka et al. (2020); Kaddour et al. (2022); Wu et al. (2022), and benign overfit Koehler et al. (2021); Wang et al. (2021b); Chatterji and Long (2021); Chen et al. (2022). In this paper, we review the related literature and organize it into the following six categories.
  1. 1.
    Complexity-based methods. Conventional statistical learning theory has established a series of upper bounds on the generalization error (generalization bounds) based on the complexity of the hypothesis space, such as VC-dimension Vapnik (2006); Wu (2021) and Rademacher complexity (Bartlett and Mendelson, 2003). Usually, these generalization bounds suggest that controlling the model size can help models generalize better. However, overparameterized DNNs make the generalization bounds vacuous.
  2. 2.
    Algorithmic Stability. Algorithmic Stability (e.g., uniform stability, hypothesis stability) is straightforward: if an algorithm fits the dataset without being overly dependent on particular data points, then achieving small empirical error implies better generalization Bousquet and Elisseeff (2002); Oneto et al. (2015); Maurer (2017); Elisseeff et al. (2005); Mukherjee et al. (2006); Shalev-Shwartz et al. (2010). Algorithmic Stability simply cares about the fact that the algorithm choice should not change too much and is stable enough if we slightly change the dataset.
  3. 3.
    PAC-Bayesian. PAC-Bayesian (or simply ”PAC-Bayes”) theory, which integrates Bayesian statistic theories and PAC theory for stochastic classifiers, has emerged as a method for obtaining some of the tightest generalization bounds Oneto (2020); Xie and Sun (2017). It emerged in the late 1990s, catalyzed by an early paper of Shawe-Taylor and Williamson (1997), and shepherded forward by McAllester McAllester (1998, 1999). McAllester’s PAC-Bayesian bounds are empirical bounds, in the sense that the upper bound only depends on known computable quantities linked to the data. Suppose the prior and the posterior of model $f$ are $\mathcal{P}$ and $\mathcal{Q}$; then the mathematical formulation of the classic McAllester’s PAC-Bayesian bound is as follows.
    $\Pr(\mathbb{E}_{f}[\mathcal{R}(f)]\leq\mathbb{E}_{f}[\widehat{\mathcal{R}}_{s}(f)]+\sqrt{\frac{D_{KL}(\mathcal{Q}\|\mathcal{P})+\log(n/\delta))}{2(n-1)}})\geq 1-\epsilon.$ (6)
    Over the years, many PAC-Bayes bounds were developed in many ever tighter variants, such as the ones for Bernoulli losses Langford and Caruana (2001); Seeger (2003); Maurer (2004), exhibiting fast-rates given small loss variances Tolstikhin and Seldin (2013); Mhammedi et al. (2019), data-dependent priors Rivasplata et al. (2020), and so on. We refer the reader to Alquier (2021) for excellent surveys. Whereas complexity-based bounds typically rely on some notion of the complexity of the hypothesis class, PAC-Bayesian bounds depend on the distance between $\mathcal{Q}$ and a prior distribution $\mathcal{P}$. The choice of $\mathcal{P}$ deeply influences the ability to assess the generalization performance of the final model.
  4. 4.
    The stochasticity of stochastic optimization algorithms as an implicit regularization. Due to the non-linear dependencies within the loss function relative to the parameters, the training of DNNs is a quintessential non-convex optimization problem, fraught with the complexities of numerous local minima. First-order optimization methods, such as SGD, have become the workhorse behind the training of DNNs. Despite its simplicity, SGD also enables high efficiency in complex and non-convex optimization problems Bottou et al. (2016). It is widely believed that the inherent stochasticity of stochastic optimization algorithms performs implicit regularization, guiding parameters towards flat minima with lower generalization errors (Keskar et al., 2016; Chaudhari and Soatto, 2017; Dauphin et al., 2014; Lee et al., 2019). To identify the flat minima, numerous mathematical definitions of flatness exist, with a prevalent one being the largest eigenvalue of the Hessian matrix of the loss function with respect to the model parameters on the training set, denoted as $\lambda_{\max}$ (Jastrzebski et al., 2017; Lewkowycz et al., 2020; Dinh et al., 2017).
  5. 5.
    The role of over-parameterization in non-convex optimization. As for the impact of parameter quantity on optimization algorithms, it is already known in the literature that DNNs in the infinite width limit are equivalent to a Gaussian process Neal and Neal (1996); Williams (1996); Winther (2000); Neal (2012); Lee et al. (2017). The work of Jacot et al. (2018) elucidates that the evolution of the trainable parameters in continuous-width DNNs during training can be captured by the neural tangent kernel (NTK). Recent work has shown that with a specialized scaling and random initialization, the parameters of continuous width two-layer DNNs are restricted in an infinitesimal region around the initialization and can be regarded as a linear model with infinite dimensional features Du et al. (2018b); Li and Liang (2018); Arora et al. (2019); Du et al. (2018a). Since the system becomes linear, the dynamics of gradient descent (GD) within this region can be tracked via properties of the associated NTK and the convergence to the global optima with a linear rate can be proved. Later, the NTK analysis of global convergence is extended to multi-layer neural nets Allen-Zhu et al. (2018a); Zou et al. (2018); Allen-Zhu et al. (2018b). An alternative research direction attempts to examine the infinite-width neural network from a mean-field perspective (Sirignano and Spiliopoulos, 2018; Mei et al., 2018; Chizat and Bach, 2018). The key idea is to characterize the learning dynamics of noisy SGD as the gradient flows over the space of probability distributions of neural network parameters. When the time goes to infinity, the noisy SGD converges to the unique global optimum of the convex objective function for the two-layer neural network with continuous width. NTK and mean-field methods provide a fundamental understanding on training dynamics and convergence properties of non-convex optimization in infinite-width neural networks.
  6. 6.
    Information-based methods. Information-theoretic generalization bounds have been empirically demonstrated to be effective in capturing the generalization behavior of Deep Neural Networks (DNNs) (Negrea et al., 2019; Hellström and Durisi, 2022a). The original information-theoretic bound introduced by Xu and Raginsky (2017) has been extended or improved in various ways, such as the chaining method (Asadi et al., 2018; Zhou et al., 2022; Clerico et al., 2022), the random subset and individual technique Negrea et al. (2019); Bu et al. (2019); Haghifam et al. (2020); G’alvez et al. (2021); Zhou et al. (2020). Information-theoretic generalization bounds have also been explored for some specific algorithms. For example, the weight-based information-theoretic bounds have been successfully applied to characterize the generalization properties of the stochastic gradient Langevin dynamics (SGLD) algorithm Pensia et al. (2018); Bu et al. (2019); Negrea et al. (2019); Haghifam et al. (2020); G’alvez et al. (2021); Wang et al. (2021a). The information bottleneck principle Shwartz-Ziv and Tishby (2017); Ahuja et al. (2021); Wongso et al. (2023), from the perspective of the information change in neural networks, posits that for better prediction, the neural network learns useful information and filters out redundant information. Beyond supervised learning, information-theoretic bounds are also useful in a variety of learning scenarios, including meta-learning Hellström and Durisi (2022b), semi-supervised learning He et al. (2021), transfer learning Wu et al. (2020), among others.
While there is yet no clarity in understanding why, for some algorithms and datasets, increasing the number of parameters actually improves generalization Oneto et al. (2023). The accumulating evidence and existing theories indicate that the exceptional performance of DNNs is attributed to a confluence of factors, including their architectural design, optimization algorithms, and even training data. Elucidating the mechanisms underlying deep learning remains an unresolved issue at present.
2.3 Distribution estimation
In statistical data analysis, determining the probability density function (PDF) or probability mass function (PMF) from a limited number of random samples is a common task, especially when dealing with complex systems where accurate modeling is challenging. Generally, there are three main methods for this problem: parametric, semi-parametric, and non-parametric models.
Parametric models assume that the data follows a specific distribution with unknown parameters. Examples include the Gaussian model, Gaussian Mixture models (GMMs), and distributions from the generalized exponential family. A major limitation of these methods is that they make quite simple parametric assumptions, which may not be sufficient or verifiable in practice. For instance, GMMs are based on the first few moments, and thus provide a poor fit for some distributions, including heavy-tailed processes (Rahman et al., 2008).
Non-parametric models estimate the PDF/PMF based on data points without making assumptions about the underlying distribution (Härdle et al., 2004). Examples include the histogram (Sturges, 1926; Scott, 1979; Freedman and Diaconis, 1981) and kernel density estimation (KDE) (Parzen, 1962; Silverman, 2018; Hall et al., 2004). While these methods are more flexible, they may suffer from overfitting or poor performance in high-dimensional spaces due to the curse of dimensionality.
Semi-parametric models (McLachlan et al., 2019) strike a balance between parametric and non-parametric methods. They do not assume the a prior-shape of the PDF/PMF to estimate as non-parametric techniques. However, unlike the non-parametric methods, the complexity of the model is fixed in advance, in order to avoid a prohibitive increase of the number of parameters with the size of the dataset, and to limit the risk of overfitting. Finite mixture models are commonly used to serve this purpose.
Statistical methods for determining the parameters of a model conditional on the data, including the standard likelihood method (Aldrich, 1997; Hald, 1999; Meeker et al., 2022), M-estimators (Huber, 1964, 1973), and Bayesian methods (MacKay, 2003; Sivia and Skilling, 2006; Murphy, 2022), have been widely employed. The Maximum Likelihood (MLE), introduced by Fisher (Aldrich, 1997), is commonly used in data regression, with earlier ideas discussed by Daniel Bernoulli and and Gauss (Hald, 1999). In 1964, Huber (1964) proposed a generalization of the maximum likelihood objective function. In 1973 Huber (1973) extended the idea of using M-estimators for the solution of regression problems. However, traditional estimators struggle in high-dimensional space: because the required sample size or computational complexity often explodes geometrically with increasing dimensionality, they have almost no practical value for high-dimensional data.
Recently, neural network-based methods have been proposed for PDF/PMF estimation and have shown promising results in problems with high-dimensional data points such as images. There are mainly two families of such neural density estimators: auto-regressive models (Larochelle and Murray, 2011; Papamakarios et al., 2017) and normalizing flows (Rezende and Mohamed, 2015; Dinh et al., 2016). Autoregression-based neural density estimators decompose the density into the product of conditional densities based on probability chain rule $p(x)=\prod_{i}p(x_{i}|x_{1:i-1})$. Each conditional probability $p(x_{i}|x_{1:i-1})$ is modeled by a parametric density (e.g., Gaussian or mixture of Gaussian), and the parameters are learned by neural networks. Density estimators based on normalizing flows represent $x$ as an invertible transformation of a latent variable $z$ with a known density, where the invertible transformation is a composition of a series of simple functions whose Jacobian is easy to compute. The parameters of these component functions are then learned by neural networks.
3 Preliminaries
3.1 Notation
  • The set of real numbers is denoted by $\mathbb{R}$, and the cardinality of set $\mathcal{Z}$ is denoted by $|\mathcal{Z}|$.
  • The $|\mathcal{Z}|$-dimensional probability simplex is denoted by $\Delta^{|\mathcal{Z}|}:=\{p\in\mathbb{R}^{|\mathcal{Z}|}\colon\|p\|_{1}=1\text{ and }p_{i}>0\,\ \forall i\}$.
  • Random variables are denoted using upper case letters such as $X,Y$ which take the values in set $\mathcal{X}$, and $\mathcal{Y}$, respectively.
  • The expectation of a random variable $X$ is denoted by $\mathbb{E}{[X]}$ or $\mathbb{E}_{X}[X]$, with $\mathbb{E}_{X\sim p}[X]$ when the distribution of $X$ is specified. And we define $\mathbb{E}_{X\sim p}[f(X)]=\sum_{x\in\mathcal{X}}p(X=x)f(x)$, where $\mathcal{X}$ is the value space of $X$.
  • The space of fitting models, which is a set of functions endowed with some structure, is represented by $\mathcal{F}_{\Theta}=\{f_{\theta}:\theta\in\Theta\}$, where $\Theta$ denotes the parameter space.
  • The sub-differential is defined as $\partial f(x)=\{s\in\mathbb{R}^{n}|f(y)-f(x)-\langle s,y-x\rangle\geq 0,\forall y\in\mathrm{dom}(f)\}$ (see e.g., (Li et al., 2020),Page 27). If $f(x)$ is smooth and convex, the sub-differential $\partial f(x)$ reduces to the usual gradient, i.e., $\partial f(x)=\{\nabla f(x)\}$.
  • The Legendre-Fenchel conjugate of a function $\Omega$ is denoted by $\Omega^{*}(f):=\sup_{p\in\mathrm{dom}(\Omega)}\langle p,f\rangle-\Omega(p)$. By default, $\Omega$ is a continuous strictly convex function, and its gradient with respect to $p$ is denoted by $p_{\Omega}^{*}$. When $\Omega(\cdot)=\frac{1}{2}\|\cdot\|_{2}^{2}$, we have $p=p_{\Omega}^{*}$.
  • The Fenchel-Young loss $d_{\Omega}\colon\mathrm{dom}(\Omega)\times\mathrm{dom}(\Omega^{*})\to\mathbb{R}^{+}$ generated by $\Omega$ is defined as:
    $d_{\Omega}(q;f):=\Omega(q)+\Omega^{*}(f)-\langle q,f\rangle,$ (7)
    where $\Omega^{*}$ denotes the conjugate of $\Omega$.
  • The Kullback-Leibler (KL) divergence is denoted by $D_{KL}(p\|q):=\sum_{i}p_{i}\log\frac{p_{i}}{q_{i}}$, where $q,p\in\Delta$.
  • The maximum and minimum eigenvalues of matrix $A$ are denoted by $\lambda_{\max}(A)$ and $\lambda_{\min}(A)$, respectively.
  • We denote the Fenchel-Young loss as $d_{\Omega}(q,f)$, and the Hessian matrix of $d_{\Omega}(q,f)$ with respect to $\theta$ is denoted as $\nabla^{2}_{\theta}d_{\Omega}(q,f)=\left[\frac{\partial^{2}d_{\Omega}(q,f_{\theta})}{\partial\theta_{i}\partial\theta_{j}}\right]_{1\leq i,j\leq d}\in\mathbb{R}^{d\times d}\enspace$.
  • To improve clarity and conciseness, we provide the following definitions of symbol:
    $q_{Y|x}(y):=q_{Y|X}(y|x),$ (8)
    $q_{Y|x}:=(q_{Y|x}(y^{(1)}),\cdots,q_{Y|x}(y^{(|\mathcal{Y}|)}))^{\top},$
    $\bar{q}:=\mathbb{E}_{q\sim\mathcal{Q}}[q],$
    $\bar{q}_{X}:=\mathbb{E}_{q_{X}\sim\mathcal{Q}_{X}}[q_{X}],$
    $\bar{q}_{Y|x}:=\mathbb{E}_{q_{Y|x}\sim\mathcal{Q}_{Y|x}}[q_{Y|x}],$
    $\mathrm{Var}(q_{Y|x}):=\mathbb{E}_{q_{Y|x}\sim\mathcal{Q}_{Y|x}}[\|q_{Y|x}-\bar{q}_{Y|x}\|_{2}^{2}],$
    where $q_{X}(x)$ and $q_{Y|X}(y|x)$ represent the marginal and conditional PMFs, respectively, the conditional distribution $q_{Y|x}$ is also modeled as a random variable.
3.2 Lemmas
The theorems in the subsequent sections rely on the following lemmas.
Based on the Fenchel-Young inequality, it is straightforward to obtain the Fenchel-Young loss which unifies many existing losses, including the hinge, logistic, and sparse-max losses, as demonstrated by Blondel et al. (2019).
Lemma 3.1 (Properties of Fenchel-Young losses)
Blondel et al. (2019) The following are the properties of Fenchel-Young losses.
  1. 1.
    $d_{\Omega}(q,f)\geq 0$ for any $q\in\mathrm{dom}(\Omega)$ and $f\in\mathrm{dom}(\Omega^{*})$. If $\Omega$ is a lower semi-continuous proper convex function, then the loss is zero iff $f\in\partial\Omega(q)$. Furthermore, when $\Omega$ is strictly convex, the loss is zero iff $f=q_{\Omega}^{*}$.
  2. 2.
    $d_{\Omega}(q,f)$ is convex with respect to $f$, and its subgradients include the residual vectors: $f_{\Omega^{*}}^{*}-q\in\partial_{f}d_{\Omega}(q;f)$. If $\Omega$ is strictly convex, then $d_{\Omega}(q,f)$ is differentiable and $\nabla_{f}d_{\Omega}(q,f)=f_{\Omega^{*}}^{*}-q$. If $\Omega$ is strongly convex, then $d_{\Omega}(q,f)$ is smooth, i.e., $\nabla_{f}d_{\Omega}(q,f)$ is Lipschitz continuous.
Throughout this paper, in line with practical applications and without affecting the conclusions, we assume that $\Omega$ is a proper, lower semi-continuous (l.s.c.), and strictly convex function or functional. Consequently, the Fenchel-Young loss attains zero if and only if $q=f_{\Omega^{*}}^{*}$.
Lemma 3.2 (Sanov’s theorem (Cover, 1999))
$\lim_{n\to\infty}-\frac{1}{n}\log\Pr(\hat{q}|q)=D_{KL}(\hat{q}\|q).$ (9)
Lemma 3.3 (Donsker-Varadhan representation)
Dupuis and Ellis (2011) Let $p$ and $q$ be distributions on a common space $\mathcal{X}$. Then
$D_{\mathrm{kl}}(p\|q)=\sup_{g}\left\{\mathbb{E}_{X\sim p}[g(X)]-\log\mathbb{E}_{X\sim q}\left[e^{g(X)}\right]\right\},$
where the supremum is taken over measurable functions $g:\mathcal{X}\rightarrow\mathbb{R}$ such that $\mathbb{E}_{X\sim q}\left[e^{g(X)}\right]<\infty$.
Lemma 3.4 (Properties of randomly sampled points on the unit ball)
Suppose that we sample $n$ points $x^{(1)},\ldots,x^{(n)}$ uniformly from the unit ball $\mathcal{B}^{m}_{1}:\{x\in\mathbb{R}^{m},\|x\|_{2}\leq 1\}$. Then with probability $1-O(1/n)$ the following holds (Blum et al., 2020, Theorem 2.8):
$\left\|x^{(i)}\right\|_{2}\geq 1-\frac{2\log n}{m},\text{ for }i=1,\ldots,n.$ (10)
$|\langle x^{(i)},x^{(j)}\rangle|\leq\frac{\sqrt{6\log n}}{\sqrt{m-1}}\text{ for all }i,j=1,\ldots,i\neq j.$
Lemma 3.5 (Gershgorin’s circle theorem)
Let $A$ be a real symmetric $n\times n$ matrix, with entries $a_{ij}$. For $i=\{1,\cdots,n\}$ let $R_{i}$ be the sum of the absolute value of the non-diagonal entries in the $i$-th row: $R_{i}=\sum_{1\leq j\leq n,j\neq i}|a_{ij}|$. For any eigenvalue $\lambda$ of $A$, there exists $i\in\{1,\cdots,n\}$ such that
$|\lambda-a_{ii}|\leq R_{i}.$ (11)
Lemma 3.6 (Markov inequality)
Suppose a random variable $Z\geq 0$, then for all $t\geq 0$, $\operatorname{Pr}(Z\geq t)\leq\mathbb{E}[Z]/t$.
Lemma 3.7 (Paley-Zygmund Inequality)
Let $X$ be a non-negative random variable and suppose that $0<\mathbb{E}[X^{2}]<\infty$. Then for any $t\in[0,1]$,
$\Pr(X>t\mathbb{E}[X])\geq(1-t)^{2}\frac{(\mathbb{E}[X])^{2}}{\mathbb{E}[X^{2}]}.$ (12)
3.3 Setting and basics
The random pair $Z=(X,Y)$ follows the distribution $q$ and takes values in the product space $\mathcal{Z}=\mathcal{X}\times\mathcal{Y}$, where $\mathcal{X}$ represents the input feature space and $\mathcal{Y}$ denotes a finite set of labels. The training data, denoted by an $n$-tuple $s^{n}:=\{z^{(i)}\}_{i=1}^{n}=\{(x^{(i)},y^{(i)})\}_{i=1}^{n}$, is composed of i.i.d. samples drawn from the unknown true distribution $q$. Given the available information, the true underlying distribution is undetermined. Consequently, we model $q$ as a random variable in the space $\Delta$. Let $\mathcal{Q}^{(s^{n},\mathcal{P})}$ (abbreviated as $\mathcal{Q}$) be the posterior distribution of $q$, conditional on the i.i.d. sample set $s^{n}$ and the prior distribution $\mathcal{P}$. Let $\tilde{q}$ denote an unbiased and consistent estimator of $q$, which derives the estimate of $q$ from the sample set $s^{n}$ and prior information $\mathcal{P}$. In the absence of prior knowledge about $q$, the Monte Carlo method (MC) is commonly utilized. This method estimates $q$ based on the event frequencies, denoted by $\hat{q}(s^{n}):=\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\{z\}}(z^{(i)})$. In this article, we refer to $\hat{q}$ as the frequency distribution estimator to distinguish it from the general estimator $\tilde{q}$. We also assume that $\tilde{q}$ is a better estimator relative to $\hat{q}$. Real-world distributions are usually smooth, which inspires the common practice of smoothing $\hat{q}$ to construct $\tilde{q}$ to approximate the true distribution $q$. These smoothing techniques effectively integrate prior knowledge and are often advantageous, as exemplified by classical Laplacian smoothing. For simplicity, when the sample set is given, we abbreviate the results estimated by the estimator $\tilde{q}(s^{n})$ and $\hat{q}(s^{n})$ as $\tilde{q}$ and $\hat{q}$, respectively.
It is important to note that, given the sample set $s^{n}$ and the prior distribution $\mathcal{P}$, all previously mentioned calculations are feasible. In this article, we utilize the notation $f_{\theta}$ (hereinafter abbreviated as $f$) to denote the model characterized by the parameter vector $\theta$. Additionally, $f_{\theta,x}$ (abbreviated as $f_{x}$) represents the model with a parameter vector $\theta$ and a specific input $x$. We treat the feature $x$ as a parameter of the fitting model. Unlike other parameters, $x$ is static and specific to a particular input. When utilizing an embedding module to process input features, $x$ transitions from a static input to trainable parameters. To characterize the structural properties of the model, we introduce the following definition:
Definition 3.1 (Structural matrix)
We define $A:=\nabla_{\theta}f^{\top}\nabla_{\theta}f$ and $A_{x}:=\nabla_{\theta}f_{x}^{\top}\nabla_{\theta}f_{x},$ as the structural matrices corresponding to the model $f$ and $f_{x}$.
For ease of reference, we list the symbol variables defined above in the table below.
Table 1: Definitions and abbreviations of key variables.
$q$ the true underlying distribution which is modeled as a random variable in the space $\Delta$.
$\mathcal{P}$ the prior distribution of $q$.
$s^{n}$($s$) i.i.d. samples drawn from the unknown true distribution $q$.
$\mathcal{Q}^{(s^{n},\mathcal{P})}$ ($\mathcal{Q}$) the posterior distribution of $q$, conditional on the i.i.d. sample set $s^{n}$ and the prior distribution $\mathcal{P}$.
$\hat{q}(s^{n})$($\hat{q}$) the frequency distribution estimator.
$\tilde{q}(s^{n})$( $\tilde{q}$) an unbiased and consistent estimator of $q$.
$f_{\theta}$($f$) the model characterized by the parameter vector $\theta$.
4 PD learning framework
4.1 Problem formulation
In this subsection, we provide the definition of PD learning and compare the PD learning framework with the traditional machine/deep learning framework to clarify the similarities and differences.
A significant portion of machine learning tasks can be conceptualized in the area of distribution learning. Both classification and regression tasks are mathematically characterized by the process of learning the conditional probability distribution of labels based on input features. Generative learning is equivalent to learning the underlying joint distribution of features. Consequently, approaching machine learning and deep learning from a perspective of distribution learning is a logical and coherent choice.
PD learning aims to estimate the underlying distribution $q$ or $q_{Y|X}$, which are modeled as random variables, using an i.i.d. sample set $s^{n}$. We refer to tasks with $q$ as the learning goal as native PD learning, while tasks with $q_{Y|X}$ as the learning goal are classified as conditional PD learning. Since the learning target is modeled as a random variable, we define the optimization objective as the probability that the discrepancy between the model’s outcome and the random distribution exceeds a specified threshold, which we call learning risk probability. For native PD learning and conditional PD learning, the learning risk probabilities are formulated as follows:
$\mathcal{R}_{n}(q,p,\epsilon)=\Pr(\|p-q\|\geq\epsilon),$ (13)
$\mathcal{R}_{c}(q,p,\epsilon)=\mathbb{E}_{X\sim\bar{q}_{X}}[\mathcal{R}_{n}(q_{Y|X},p_{Y|X},\epsilon)],$
where ϵ+italic-$\epsilon\in\mathbb{R}^{+}$, and $\mathcal{R}_{c}(q,p,\epsilon)$ the expectation of $\mathcal{R}_{n}(q_{Y|X},p_{Y|X},\epsilon)$ under $X\sim\bar{q}_{X}$. We thus formulate PD learning in two situations when employing the Fenchel-Young loss $d_{\Omega}(\cdot,\cdot)$ as the loss function:
  • Native PD learning
    $\min_{\theta}\mathcal{R}_{n}(q,p,\epsilon)\,\ \mathrm{s.t.}\,\ s=\{z_{i}\}_{i=1}^{n},$ (14)
    where $p=(f_{\theta})_{\Omega^{*}}^{*}$, ϵ+italic-$\epsilon\in\mathbb{R}^{+}$.
  • Conditional PD learning
    $\min_{\theta}\mathcal{R}_{c}(q,p,\epsilon)\,\ \mathrm{s.t.}\,\ s=\{z_{i}\}_{i=1}^{n},$ (15)
    where $p_{Y|x}=(f_{\theta,x})_{\Omega^{*}}^{*}$, ϵ+italic-$\epsilon\in\mathbb{R}^{+}$.
In PD learning, we utilize the Legendre-Fenchel conjugates of $f_{\theta}$ and $f_{\theta,x}$, denoted as $p$ and $p_{Y|x}$, to approximate the $q$ and $q_{Y|x}$. In the subsection 4.4, we will demonstrate that loss functions satisfying the conditions of unique optimum and sampling stability, which we term standard loss functions, can be expressed in the form of Fenchel-Young loss. Additionally, we will illustrate that most currently used loss functions are indeed standard loss functions.
Once the conditional distribution function of the label with respect to a specified feature or the joint distribution function is determined, it is feasible to determine the distribution of the label corresponding to the given feature. Therefore, within the machine learning paradigm, native PD learning aligns with generative learning, while conditional PD learning is inherently associated with discriminative learning. Native PD learning can be considered as a specific case of conditional PD learning when the feature variable $X$ is fixed, i.e., $X\sim\delta_{x}$. Conversely, conditional PD learning can be viewed as a scenario where multiple native PD learning tasks are conducted concurrently. This paper employs the $L_{2}$-norm for quantifying discrepancies; alternative norms are also feasible in light of the equivalence of norms in finite-dimensional spaces.
4.2 Learning risk probability optimization
Direct optimization of learning risk probability is exceedingly challenging. In this subsection, we introduce these components: learning error, uncertainty, fitting error, and estimation error. We demonstrate that optimizing these components can effectively optimize the learning risk probability.
We begin by providing the specific definitions of these concepts as follows:
  1. 1.
    Fitting Error: This error measures the performance of the model in fitting the target $\tilde{q}$, denoted by $\mathrm{Fit}_{m}(\tilde{q},p)=\|p-\tilde{q}\|_{m}$, where $\|\cdot\|_{m}$ denotes the $L_{m}$-norm. In the absence of the index $m$, the default norm is the $L_{2}$ norm.
  2. 2.
    Uncertainty: Since the variances $\mathrm{Var}(q)$ and $\mathrm{Var}(q_{Y|x})$ are independent of the models $f$ and $f_{x}$ and depend solely on the sample set and prior knowledge, we refer to them as the uncertainties of $q$ and $q_{Y|x}$ respectively.
  3. 3.
    Learning Error We define learning error as $\mathrm{Lrn}(q,p)=\mathbb{E}_{q\sim\mathcal{Q}}[\|q-p\|_{2}^{2}]$. According to the bias-variance decomposition, this error equals the sum of the uncertainty of $q$ and the fitting error relative to the expected value $\bar{q}$, denoted as $\mathrm{Lrn}(q,p)=\mathrm{Var}(q)+\mathrm{Fit}(\bar{q},p)^{2}$. We will subsequently prove that the learning error can bound the learning risk probability from both upper and lower bounds.
  4. 4.
    Estimation Error: This error quantifies the distance between $\bar{q}$ and the corresponding provided estimate, given by $\mathrm{Est}_{m}(\tilde{q})=\|\tilde{q}-\bar{q}\|_{m}$. In the absence of the index $m$, the default norm is the $L_{2}$ norm. The estimation error arises from the finite sample size and serves as a model-independent metric that characterizes the estimator. Since in conditional PD learning, the estimation error exists in both the feature set $\{x^{(i)}\}_{i=1}^{n}$ and the label set $\{y^{(i)}\}_{i=1}^{n}$, we formally define the label estimation error as $\mathrm{Est}(\tilde{q}_{Y|x})$, and the feature estimation error as $\mathrm{Est}_{1}(\tilde{q}_{X})$.
To demonstrate that the learning risk probability can be effectively controlled by optimizing the fitting and estimation errors, we begin by presenting the following lemma:
Lemma 4.1
Given a function $\ell:\mathcal{Z}\to[0,1]$, the following inequality holds:
$\sup_{\ell}|\mathbb{E}_{Z\sim\hat{q}}[\ell(z)]-\mathbb{E}_{Z\sim\bar{q}}[\ell(z)]|=\frac{1}{2}\mathrm{Est}_{1}(\hat{q}).$ (16)
Utilizing Markov’s inequality 3.6 and the Paley- Zygmund inequality 3.7, we derive the upper and lower bounds for the learning risk probability in native PD learning, which encompasses both fitting and estimation errors .
Theorem 4.1
The following formulas hold:
$\mathcal{R}_{n}(q,p,\epsilon)\leq\frac{\mathrm{Lrn}(q,p)}{\epsilon^{2}},\text{ for }\epsilon^{2}>\mathrm{Lrn}(q,p),$ (17)
$\mathcal{R}_{n}(q,p,\epsilon)\geq\frac{(\mathrm{Lrn}(q,p)-\epsilon^{2})^{2}}{4},\text{ for }\epsilon^{2}\leq\mathrm{Lrn}(q,p),$
$\mathrm{Lrn}(q,p)=\mathrm{Var}(q)+\mathrm{Fit}(\bar{q},p)^{2},$
$(\mathrm{Fit}(\tilde{q},p)-\mathrm{Est}(\tilde{q}))^{2}\leq\mathrm{Fit}(\bar{q},p)^{2}\leq(\mathrm{Fit}(\tilde{q},p)+\mathrm{Est}(\tilde{q}))^{2}.$
We then derive the upper and lower bounds on the learning risk probability in conditional PD learning.
Theorem 4.2
The following formulas hold:
$|\mathbb{E}_{X\sim\tilde{q}_{X}}[\mathcal{R}_{n}(q_{Y|x},p_{Y|x},\epsilon)]-\mathcal{R}_{c}(q,p,\epsilon)|\leq\frac{1}{2}\mathrm{Est}_{1}(\tilde{q}_{X}),$ (18)
$\mathcal{R}_{n}(q_{Y|x},p_{Y|x},\epsilon)\leq\frac{\mathrm{Lrn}(q_{Y|x},p_{Y|x})}{\epsilon^{2}},\text{ for }\epsilon^{2}\leq\mathrm{Lrn}(q_{Y|x},p_{Y|x}),$
$\mathcal{R}_{n}(q_{Y|x},p_{Y|x},\epsilon)\geq\frac{(\mathrm{Lrn}(q_{Y|x},p_{Y|x})-\epsilon^{2})^{2}}{4},\text{ for }\epsilon^{2}\leq\mathrm{Lrn}(q_{Y|x},p_{Y|x}),$
$\mathrm{Lrn}(q_{Y|x},p_{Y|x})=\mathrm{Var}(q_{Y|x})+\mathrm{Fit}(\bar{q}_{Y|x},p_{Y|x})^{2},$
$|\mathrm{Fit}(\tilde{q}_{Y|x},p_{Y|x})-\mathrm{Est}(\tilde{q}_{Y|x})|\leq\mathrm{Fit}(\bar{q}_{Y|x},p_{Y|x})\leq\mathrm{Fit}(\tilde{q}_{Y|x},p_{Y|x})+\mathrm{Est}(\tilde{q}_{Y|x}).$
The preceding theorems indicate that the learning risk probability can be controlled by the learning error. The learning error is bounded by the estimation error and the fitting error with respect to the estimators $\tilde{q}$, $\tilde{q}_{X}$ and $\tilde{q}_{Y|x}$. Therefore, to minimize the learning risk probability, it is essential to reduce the estimation error and adjust the model to align with $\bar{q}$ and $\bar{q}_{Y|x}$ by fitting $\tilde{q}$ and $\tilde{q}_{Y|x}$. Furthermore, by employing $\tilde{q}$ as the fitting target instead of $\bar{q}$, the following corollary is derived according to Theorems 4.1 and 4.2.
Corollary 4.1
Given that the estimation error is fixed, we have:
  1. 1.
    Reducing the fitting error can decrease the upper bound of the learning error.
  2. 2.
    The optimal lower bound of the learning error is achieved when the fitting error and the estimation error are equal.
  3. 3.
    Uncertainty represents the least upper bound of the learning error.
The corollary suggests that a decrease in the fitting error could potentially increase the magnitude of the learning error, as this may result in an increase in the lower bound of the learning error. In other words, achieving the minimum fitting error does not always lead to the optimal solution. This provides a theoretical explanation for the occurrence of overfitting in machine learning. Since uncertainty represents the minimum upper bound of the learning error, and it is independent of algorithms and models, being solely dependent on data and prior knowledge, this also theoretically substantiates the assertion that the quality of data and features determines the upper limit of machine learning performance, whereas models and algorithms merely serve to approximate this limit.
4.3 Estimation error optimization
In the previous subsection, we showed that reducing the estimation error effectively regulates the learning risk probability. In this subsection, we will examine the properties of the estimation error and the relevant optimization strategies.
According to Theorems 4.1 and 4.2, it is obvious that the following conclusion can be drawn.
Corollary 4.2
For native PD learning, $\mathrm{Est}(\tilde{q})$ attains its minimum value of 0 when $\tilde{q}=\bar{q}$ In the case of conditional PD learning, the optimal solution for $\mathrm{Est}(\tilde{q}_{Y|x})$ is $\tilde{q}_{Y|x}=\bar{q}_{Y|x}$, and the optimal solution for $\mathrm{Est}_{1}(\tilde{q}_{X})$ is $\tilde{q}_{X}=\bar{q}_{X}$, for all $x\in\mathcal{X}$.
Therefore, we refer to $\bar{q}$, $\bar{q}_{Y|x}$ and $\bar{q}_{X}$ as the optimal estimators for native PD learning and conditional PD learning, respectively. The calculation method for $\bar{q}$ is outlined in the following theorem, and this method is equally applicable to $\bar{q}_{Y|x}$.
Theorem 4.3
With the provision of an i.i.d. sample set $s^{n}$ and a prior distribution $\mathcal{P}$, we have
$\bar{q}=\sum_{q_{i}\in\Delta}\mathcal{Q}(q_{i})q_{i},$ (19)
$\mathcal{Q}(q_{0})=\Pr(s^{n},q_{0})/\sum_{q_{i}\in\Delta}\Pr(s^{n},q_{i}),$
where $n(z)=\sum_{i=1}^{n}\mathbf{1}_{\{z\}}(z^{(i)})$, $\Pr(s^{n},q_{i})=\mathcal{P}(q_{i})n!\prod_{z\in\mathcal{Z}}\frac{q_{i}(z)^{n(z)}}{n(z)!}$.
This theorem describes the distribution of the latent distribution of the whole dataset given an i.i.d. sample set and prior knowledge. It also outlines the computational methodology for our proposed optimal estimator $\bar{q}$. In PD learning, the existence and computability of the optimal estimate represent a significant characteristic and benefit, distinguishing it from traditional theoretical frameworks. In the absence of further information regarding the i.i.d. sample $s^{n}$ for $q$, we invoke the equal-probability hypothesis, according to which $\mathcal{P}$ is assumed to be the uniform distribution across the space $\Delta$. The equal-probability hypothesis is one of the fundamental assumptions of statistical physics, positing that all probabilities are uniformly distributed over the space $\Delta$. Assuming the equal-probability hypothesis holds, it follows that
$\mathcal{Q}(q_{0})=(n!\prod_{z\in\mathcal{Z}}\frac{q_{0}(z)^{n(z)}}{n(z)!})/\sum_{q_{i}\in\Delta}(n!\prod_{z\in\mathcal{Z}}\frac{q_{i}(z)^{n(z)}}{n(z)!}).$ (20)
Although, based on Theorem 4.3, $\bar{q}$, $\bar{q}_{Y|x}$ and $\bar{q}_{X}$ are computationally tractable, the computational cost associated with them is typically high, leading to the widespread practice of using simpler unbiased and consistent estimators as substitutes. Since both $\tilde{q}$ and $\hat{q}$ serve as unbiased and consistent estimators of $q$, we can derive the following theorem.
Theorem 4.4
$\lim_{n\to\infty}\hat{q}=\tilde{q}=\bar{q},$ (21)
$\lim_{n\to\infty}\mathrm{Var}(q)=0.$
This theorem suggests that, given a sufficiently large sample size, the estimation error associated with an arbitrary unbiased and consistent estimator is negligible.
Beyond the sample set, the prior distribution $\mathcal{P}$ also exerts a significant influence on the estimation error. For a typical deterministic single-label classification with negligible noise, the label probability distribution conditioned on feature $x$ is given by $q_{Y|x}=\delta_{y}$, where $y$ is the label associated with $x$. Under these conditions, it follows that $\forall x\in\mathcal{X}$, $\mathrm{Est}(\tilde{q}_{Y|x})=0$ and $\mathrm{Var}(q_{Y|x})=0$.
4.4 Fitting error optimization
Minimizing fitting error typically encounters the challenge of non-convex optimization. To address this issue, we introduce the standard loss function and demonstrate that by using a standard loss, an approximate global optimal solution for the fitting error minimization can be achieved by both diminishing $1/\lambda_{\max}(A)$ and $1/\lambda_{\min}(A)$, as well as by reducing the gradient norm of the model. Drawing upon this finding, we introduce the GSC algorithm and provide an upper bound formula for the fitting error in the context of this algorithm.
4.4.1 Standard loss function
Definition 4.1 (Standard loss function)
A loss function $\ell(q,p):\Delta\times\Delta\to\mathbb{R}^{+}$ is regarded as a standard loss if it meets the following conditions:
  • Unique optimum. The loss function is strictly convex, ensuring that it has a unique minimum when predicting the correct probability $q$, i.e., $p=q$.
  • Sampling stability. The expected loss, $\mathbb{E}_{q}\ell(q,p)$, achieves its minimum value when $p=\mathbb{E}_{q}[q]$, expressed as $\mathbb{E}_{q}[q]=\arg\min_{p}\mathbb{E}_{q}[\ell(q,p)]$.
The unique optimum condition guarantees that the standard loss function is effective in directing the model towards the desired fitting target. Furthermore, the sampling stability condition ensures the consistency of optimization results across different optimization methods, such as SGD, batch-GD, and GD. In other words, minimizing the loss function on the entire dataset yields results consistent with those obtained by minimizing the expected value of the loss function under random sampling. Consequently, the definition of the standard loss function turns out to be an intrinsic result when viewed from the perspective of algorithm optimization.
We establish the equivalence of the standard losses to the Fenchel-Young loss in the following theorem.
Theorem 4.5
Given any two vectors $\mu,\nu$ and a standard loss $\ell(\mu,\nu)$, there exists a strictly convex function $\Omega$ such that $\ell(\mu,\nu)-\ell(\mu,\nu)=d_{\Omega}(\mu,\nu_{\Omega}^{*})$, where $d_{\Omega}(\mu,\nu_{\Omega}^{*})$ denotes the Fenchel-Young loss.
In our study, we represent the standard loss as $d_{\Omega}(q,f_{\theta})$, where $f_{\theta}$ denotes the model parameterized by $\theta$. To maintain generality without compromising clarity, we assume that $d_{\Omega}(q,f_{\theta})$ is continuous and differentiable with respect to $\theta$. The standard loss function is a broad concept that encompasses a wide range of commonly used loss functions in statistics and machine learning, as illustrated in Table 2.
Table 2: Examples of common loss functions and their corresponding standard loss forms (Blondel et al., 2019)
Loss $\mathrm{dom}(\Omega)$ $\Omega(\mu)$ $\hat{y}_{\Omega}(\theta)$ $d_{\Omega}(\theta;y)$
Squared $\mathbb{R}^{d}$ $\frac{1}{2}\|\mu\|^{2}$ $\theta$ $\frac{1}{2}\|y-\theta\|^{2}$
Perceptron $\Delta^{|\mathcal{Y}|}$ 00 $\arg\max(\theta)$ $\max_{i}\theta_{i}-\theta_{k}$
Logistic $\Delta^{|\mathcal{Y}|}$ $-H(\mu)$ $\mathrm{softmax}(\theta)$ $\log\sum_{i}\exp\theta_{i}-\theta_{k}$
Hinge $\Delta^{|\mathcal{Y}|}$ $\langle{\mu},{\mathbf{e}_{k}-\mathbf{1}}\rangle$ $\arg\max(\mathbf{1}-\mathbf{e}_{k}+\theta)$ $\max_{i}\leavevmode\nobreak\ [[i\neq k]]+\theta_{i}-\theta_{k}$
Sparsemax $\Delta^{|\mathcal{Y}|}$ $\frac{1}{2}\|\mu\|^{2}$ $\mathrm{sparsemax}(\theta)$ $\frac{1}{2}\|y-\theta\|^{2}-\frac{1}{2}\|\hat{y}_{\Omega}(\theta)-\theta\|^{2}$
Logistic (one-vs-all) $[0,1]^{|\mathcal{Y}|}$ $-\sum_{i}H([\mu_{i},1-\mu_{i}])$ $\mathrm{sigmoid}(\theta)$ $\sum_{i}\log(1+\exp(-(2y_{i}-1)\theta_{i}))$
  • 1
    $\hat{y}_{\Omega}(\theta)\in\arg\max_{p\in\mathrm{dom}(\Omega)}\langle\theta,p\rangle-\Omega(p)$.
  • 2
    For multi-class classification, we assume that $\mathcal{Y}=\{\mathbf{e}_{i}\}_{i=1}^{d}$ and the ground-truth is $y=\mathbf{e}_{k}$, where $\mathbf{e}_{i}$ denotes a standard basis (”one-hot”) vector.
  • 3
    We denote the Shannon entropy by $H(p):=-\sum_{i}p_{i}\log p_{i}$, where $p\in\Delta^{\mathcal{Y}}$.
4.4.2 Bounds on fitting error
We begin by defining a fitting error optimization problem in native PD learning:
$\min_{\theta}\mathrm{Fit}(\tilde{q},p),$ (22)
where $p=(f_{\theta})_{\Omega}^{*}$ is continuously differentiable with respect to $\theta$, $\tilde{q}$ is the constant vector we aim to fit, and $|\tilde{q}|$ represents its dimensionality.
We proceed to present a central theoretical claim:
Theorem 4.6
If $\lambda_{\min}(A)\neq 0$, we have
$\frac{\|\nabla_{\theta}d_{\Omega}(\tilde{q},f)\|_{2}^{2}}{\lambda_{\max}(A)}\leq\mathrm{Fit}^{2}(\tilde{q},p)\leq\frac{\|\nabla_{\theta}d_{\Omega}(\tilde{q},f)\|_{2}^{2}}{\lambda_{\min}(A)}.$ (23)
This theorem leads to the following significant conclusions:
  1. 1.
    Since $\mathrm{Fit}(\tilde{q},p)$ is equivalent to $\|\nabla_{\theta}d_{\Omega}(\tilde{q},f)\|_{2}$ when $\lambda_{\min}(A)\neq 0$, in the context of the aforementioned non-convex optimization problem 22, the global optima can be approximated by concurrently minimizing $1/\lambda_{\min}(A)$, $1/\lambda_{\max}(A)$ and $\|\nabla_{\theta}d_{\Omega}(\tilde{q},f)\|$.
  2. 2.
    The gradient-based iterative algorithms, such as GD and SGD, allow the model parameters to converge to stationary points, where $\|\nabla_{\theta}d_{\Omega}(\tilde{q},f)\|=0$. The above theorem also provides the condition under which the gradient-based iterative algorithms can converge to a globally optimal solution, i.e., $\lambda_{\max}(A)=\lambda_{\min}(A)$ and $\lambda_{\min}(A)\neq 0$.
Based on the established Theorem 4.6, we can derive the upper bound on the fitting error in conditional PD learning.
Corollary 4.3
If $\lambda_{\min}(A_{x})\neq 0$, we have
$\frac{\|\nabla_{\theta}d_{\Omega}(\tilde{q}_{Y|x},f_{x})\|_{2}^{2}}{\lambda_{\max}(A_{x})}\leq\mathrm{Fit}^{2}(\tilde{q}_{Y|x},p)\leq\frac{\|\nabla_{\theta}d_{\Omega}(\tilde{q}_{Y|x},f_{x})\|_{2}^{2}}{\lambda_{\min}(A_{x})}.$ (24)
Similarly, in conditional PD learning, the gradient-based iterative algorithms are able to converge to a globally optimal solution, provided that the following condition is satisfied:
$\lambda_{\max}(A_{x})\approx\lambda_{\min}(A_{x})\neq 0,\forall x\in\mathcal{X}.$ (25)
4.4.3 GSC algorithm
According to Theorem 4.6, a natural progression subsequently guides the development of the GSC algorithm. The GSC algorithm for native PD learning is mathematically formulated as follows:
$\min_{\theta}\{\log\|\nabla_{\theta}d_{\Omega}(\tilde{q},f)\|_{2}^{2}+(\alpha G(A)+\beta U(A)+\gamma L(A))\},$ (26)
where $U(A)=-\log\lambda_{\min}(A)$, $L(A)=-\log\lambda_{\max}(A)$, $G(A)=U(A)-L(A)$, $\alpha$, $\beta$ and $\gamma$ signify the weights associated with $G(A)$, $U(A)$ and $L(A)$, respectively. According to the definition, $U(A)$ and $L(A)$ represent the coefficients of the upper and lower bounds, respectively, while $G(A)$ denotes the gap between these two coefficients.
For convenience in description, we define the structural error as follows:
$S(\alpha,\beta,\gamma,A)=\alpha G(A)+\beta U(A)+\gamma L(A).$ (27)
Thus, the GSC algorithm aims to minimize $\|\nabla_{\theta}d_{\Omega}(\tilde{q},f)\|_{2}^{2}$ and the structural error $S(\alpha,\beta,\gamma,A)$ concurrently. It is evident that the efficacy of the GSC algorithm is contingent upon the non-vanishing nature of $\lambda_{\min}(A)$; specifically, if $\lambda_{\min}(A)=0$, the control over the upper bound of the fitting error via the gradient norm becomes unattainable.
Gradient-based optimization methods, including SGD, Batch-GD, and Newton’s method, primarily focus on minimizing the squared $L_{2}$ norm of the gradient, $\|\nabla_{\theta}d(\tilde{q},f)\|_{2}^{2}$. Alternative methods, such as simulated annealing and genetic algorithms, can also be employed as complementary strategies to achieve this objective. Furthermore, employing parameter initialization with small values and implementing weight normalization, which together control the amplitude of parameters, significantly aids in regulating gradient norms.
Given a constant $\|\nabla_{\theta}d(\tilde{q},f)\|_{2}^{2}$, a smaller structural error implies a smaller fitting error. On the other hand, $U(A)$ and $L(A)$ bound the rates at which the fitting error is mitigated via the reduction of the gradient norm, while $G(A)$ determines the actual range of reduction in the fitting error after gradient descent. This suggests that when applying gradient-based iterative algorithms, larger values of $U(A)$ and $L(A)$ can achieve faster convergence rates of the fitting error. Therefore, by employing parameter initialization techniques and model structure design to enhance $U(A)$ and $L(A)$, the convergence of the fitting error can be further facilitated. As for the optimization of $S(\alpha,\beta,\gamma,A)$, we establish the following reasonable assumption as the foundation for our analysis on the impact of parameter count and model depth: Uniform Sampling Assumption: Each column of $\nabla_{\theta}f_{x}$ is approximated as uniformly sampled from a ball $\mathcal{B}^{|\theta|}_{\epsilon}:\{x\in\mathbb{R}^{\theta},\|x\|_{2}\leq\epsilon\}$.
It should be noted that the aforementioned assumptions are not inherently and unconditionally valid. In section 5, we will explain how numerous existing deep learning techniques contribute to the substantiation of these assumptions. Based on the established Theorem 4.6, we deduce the following theorem to provide a theoretical basis for reducing structural error.
Theorem 4.7
  1. 1.
    If $|\theta|<|\tilde{q}|$, we have
    $\lambda_{\min}(A)$ $=0,$ (28)
    $\frac{\|\nabla_{\theta}d_{\Omega}(\tilde{q},f)\|_{2}^{2}}{\lambda_{\max}(A)}$ $\leq\mathrm{Fit}^{2}(\tilde{q},p).$
  2. 2.
    If $A=aI$, it follows that:
    $\mathrm{Fit}(\tilde{q},p)=\|\nabla_{\theta}d_{\Omega}(\tilde{q},f)\|_{2}^{2}/a.$ (29)
  3. 3.
    If the model satisfies the uniform sampling assumption 4.4.3 and $\tilde{q}\leq|\theta|-1$, then with probability at least $1-O(1/|\tilde{q}|)$ the following holds:
    $\lambda_{\max}(A)$ $\geq(1-\frac{2\log|\tilde{q}|}{|\theta|})^{2}\epsilon^{2},$ (30)
    $\lambda_{\max}(A)-$ $\lambda_{\min}(A)\leq\frac{2|\tilde{q}|\sqrt{6\log|\tilde{q}|}\epsilon^{2}}{\sqrt{|\theta|-1}}.$
The present theorem motivates the implementation of several strategies to regulate the structural error, thereby facilitating the convergence of the model to the optimal price. These strategies encompass:
  1. 1.
    Mitigating the dependencies among parameters to facilitate the validity of the uniform sampling assumption;
  2. 2.
    Increasing the number of the model parameters when the model aligns with the uniform sampling assumption.
  3. 3.
    Utilizing a specialized initialization method or model architecture to ensure that $G(A)\approx 0$ ($A\approx aI$), and this approximation is maintained throughout the training process.
In section 5, we will elucidate that numerous existing deep learning techniques essentially correspond to the above strategies.
Similarly, the GSC algorithm under conditional PD learning is formulated as
$\min_{\theta}\mathbb{E}_{X}[\log\|\nabla_{\theta}d_{\Omega}(\tilde{q}_{Y|X},f_{X})\|_{2}^{2}+S(\alpha,\beta,\gamma,A_{X})].$ (31)
As its processing method is analogous to that in the native PD learning scenario, it will not be further elaborated here.
4.5 Theoretical framework of PD learning
In this subsection, we provide a comprehensive description of the PD learning framework and summarize the algorithm logic and components in PD learning. In addition, we compare PD learning with the traditional deep learning framework to clarify their similarities and differences, highlighting the advantages inherent in the PD learning framework.
Through the integration of Theorem 4.6 with Theorem 4.1, we derive the following theoretical framework of native PD learning:
  1. 1.
    Transform the Learning Probability Risk minimization into $\mathrm{Fit}(\bar{q},p)$ minimization.
    $\mathcal{R}_{n}(q,p,\epsilon)\leq\frac{\mathrm{Lrn}(q,p)}{\epsilon^{2}},\text{ for }\epsilon^{2}>\mathrm{Lrn}(q,p),$ (32)
    $\mathcal{R}_{n}(q,p,\epsilon)\geq\frac{(\mathrm{Lrn}(q,p)-\epsilon^{2})^{2}}{4},\text{ for }\epsilon^{2}\leq\mathrm{Lrn}(q,p),$
    $\mathrm{Lrn}(q,p)=\mathrm{Var}(q)+\mathrm{Fit}(\bar{q},p)^{2}.$
    The learning risk probability $\mathcal{R}_{n}(q,p,\epsilon)$ is bounded by learning error $\mathrm{Lrn}(q,p)$, which can be decomposed into the uncertainty $\mathrm{Var}(q)$ and the fitting error to the optimal estimate $\mathrm{Fit}(\bar{q},p)$. Since uncertainty solely depends on prior knowledge and sampling data, minimizing $\mathcal{R}_{n}(q,p,\epsilon)$ can be achieved by minimizing $\mathrm{Fit}(\bar{q},p)$.
  2. 2.
    Minimize $\mathrm{Fit}(\bar{q},p)$ by controlling $\mathrm{Fit}(\tilde{q},p)$ and $\mathrm{Est}(\tilde{q})$.
    $(\mathrm{Fit}(\tilde{q},p)-\mathrm{Est}(\tilde{q}))^{2}\leq\mathrm{Fit}(\bar{q},p)^{2}\leq(\mathrm{Fit}(\tilde{q},p)+\mathrm{Est}(\tilde{q}))^{2},$ (33)
    where the computation of $\bar{q}$ is detailed in Theorem  4.3. When the calculation of $\bar{q}$ is complex, we use an unbiased and consistent estimator $\tilde{q}$ to approximate it. The term $\mathrm{Fit}(\bar{q},p)$ is controlled by the estimation error of $\tilde{q}$, denoted as $\mathrm{Est}(\tilde{q})$, and the model’s fitting error to $\tilde{q}$, denoted as $\mathrm{Fit}(\tilde{q},p)$. Reducing both $\mathrm{Est}(\tilde{q})$ and $\mathrm{Fit}(\tilde{q},p)$ can decrease the upper bound of $\mathrm{Fit}(\bar{q},p)$, while reducing their discrepancy can lower the lower bound of $\mathrm{Fit}(\tilde{q},p)$. Consequently, when $\mathrm{Est}(\tilde{q})$ is large, decreasing $\mathrm{Fit}(\tilde{q},p)$ may increase $\mathrm{Fit}(\bar{q},p)$.
  3. 3.
    Achieve fitting error minimization by GSC algorithm.
    $\frac{\|\nabla_{\theta}d_{\Omega}(\tilde{q},f)\|_{2}^{2}}{\lambda_{\max}(A)}\leq\mathrm{Fit}^{2}(\tilde{q},p)\leq\frac{\|\nabla_{\theta}d_{\Omega}(\tilde{q},f)\|_{2}^{2}}{\lambda_{\min}(A)},$ (34)
    where $\lambda_{\min}(A)\neq 0$. The model’s fitting error to $\tilde{q}$, $\mathrm{Fit}(\tilde{q},p)$, is governed by $\|\nabla_{\theta}d_{\Omega}(\tilde{q},f)\|_{2}^{2}$, $1/\lambda_{\min}(A)$, and $1/\lambda_{\max}(A)$. Its minimization can be achieved through the GSC algorithm.
  4. 4.
    Increasing the training sample size can reduce the estimation error.
    $\lim_{n\to\infty}\hat{q}=\tilde{q}=\bar{q},$ (35)
    $\lim_{n\to\infty}\mathrm{Est}(\tilde{q})=0.$
    As the number of sampling data approaches infinity, any unbiased and consistent estimate is equal to the optimal estimate.
  5. 5.
    Increasing the training sample size helps reduce the uncertainty.
    $\lim_{n\to\infty}\mathrm{Var}(q)=0.$ (36)
    When the number of sampling samples approaches infinity, the uncertainty tends to zero.
In conditional PD learning, $\tilde{q}_{Y|x}$ is used as the fitting target for the model. Since $\bar{q}_{X}$ is difficult to calculate, we employ $\tilde{q}_{X}$ as an approximation. Then the actual optimization objective is $\mathbb{E}_{X\sim\tilde{q}_{X}}[\mathrm{Fit}(\tilde{q}_{Y|X},p_{Y|X})^{2}]$. Consequently, in contrast to native PD learning, it is necessary to evaluate the error associated with approximating $\bar{q}_{X}$ by $\tilde{q}_{X}$. Based on Theorem 4.2 and Corollary 4.3, we can derive the following theoretical framework of conditional PD learning:
  1. 1.
    The learning risk probability bias introduced by feature $X$ is controlled by $\tilde{q}_{X}$.
    $|\mathbb{E}_{X\sim\tilde{q}_{X}}[\mathcal{R}_{n}(q_{Y|x},p_{Y|x},\epsilon])-\mathcal{R}_{c}(q,p,\epsilon)|\leq\frac{1}{2}\mathrm{Est}_{1}(\tilde{q}_{X}),$ (37)
    where the computation of $\bar{q}_{X}$ is detailed in Theorem  4.3. When the calculation of $\bar{q}_{X}$ is complex, we use $\tilde{q}_{X}$ to approximate it. Consequently, compared to native PD learning, conditional PD learning incurs additional errors due to the approximation of $\bar{q}_{X}$ by $\tilde{q}_{X}$. The conditional learning risk probability $\mathcal{R}_{c}(q,p,\epsilon)$ is bounded by $\mathbb{E}_{X\sim\tilde{q}_{X}}[\mathcal{R}_{n}(q_{Y|x},p_{Y|x},\epsilon)]$ and $\mathrm{Est}_{1}(\tilde{q}_{X})$.
  2. 2.
    Transform the Learning Probability Risk minimization into $\mathrm{Fit}(\bar{q},p)$ minimization.
    $\mathcal{R}_{n}(q_{Y|x},p_{Y|x},\epsilon)\leq\frac{\mathrm{Lrn}(q_{Y|x},p_{Y|x})}{\epsilon^{2}},\text{ for }\epsilon^{2}\leq\mathrm{Lrn}(q_{Y|x},p_{Y|x}),$ (38)
    $\mathcal{R}_{n}(q_{Y|x},p_{Y|x},\epsilon)\geq\frac{(\mathrm{Lrn}(q_{Y|x},p_{Y|x})-\epsilon^{2})^{2}}{4},\text{ for }\epsilon^{2}\leq\mathrm{Lrn}(q_{Y|x},p_{Y|x}),$
    $\mathrm{Lrn}(q_{Y|x},p_{Y|x})=\mathrm{Var}(q_{Y|x})+\mathrm{Fit}(\bar{q}_{Y|x},p_{Y|x})^{2}.$
    The term $\mathbb{E}_{X\sim\tilde{q}_{X}}[\mathcal{R}_{n}(q_{Y|x},p_{Y|x},\epsilon)]$ is bounded by the expected learning error $\mathbb{E}_{X\sim\tilde{q}_{X}}\mathrm{Lrn}(q_{Y|X},p_{Y|X})$, which can be decomposed into the expected uncertainty $\mathbb{E}_{X\sim\tilde{q}_{X}}[\mathrm{Var}(q_{Y|X})]$ and the expected fitting error to the optimal estimate $\mathbb{E}_{X\sim\tilde{q}_{X}}[\mathrm{Fit}(\bar{q}_{Y|X},p_{Y|X})^{2}]$. Since uncertainty depends solely on prior knowledge and sampling data, minimizing $\mathbb{E}_{X\sim\tilde{q}_{X}}[\mathcal{R}_{n}(q_{Y|x},p_{Y|x},\epsilon)]$ can be achieved by minimizing $\mathbb{E}_{X\sim\tilde{q}_{X}}[\mathrm{Fit}(\bar{q}_{Y|X},p_{Y|X})^{2}]$.
  3. 3.
    Minimize $\mathrm{Fit}(\bar{q}_{Y|x},p_{Y|x})$ by controlling $\mathrm{Fit}(\tilde{q}_{Y|x},p_{Y|x})$ and $\mathrm{Est}(\tilde{q}_{Y|x})$
    $|\mathrm{Fit}(\tilde{q}_{Y|x},p_{Y|x})-\mathrm{Est}(\tilde{q}_{Y|x})|\leq\mathrm{Fit}(\bar{q}_{Y|x},p_{Y|x})\leq\mathrm{Fit}(\tilde{q}_{Y|x},p_{Y|x})+\mathrm{Est}(\tilde{q}_{Y|x}),$ (39)
    where the computation of $\bar{q}_{Y|x}$ is detailed in Theorem  4.3. When using $\tilde{q}_{Y|x}$ to approximate $\bar{q}_{Y|x}$, the fitting error $\mathrm{Fit}(\bar{q}_{Y|x},p_{Y|x})$ is controlled by the estimation error of the unbiased estimate $\tilde{q}_{Y|x}$, denoted as $\mathrm{Est}(\tilde{q}_{Y|x})$, and the model’s fitting error to $\tilde{q}_{Y|x}$, denoted as $\mathrm{Fit}(\tilde{q}_{Y|x},p_{Y|x})$. Reducing both $\mathrm{Est}(\tilde{q}_{Y|x})$ and $\mathrm{Fit}(\tilde{q}_{Y|x},p_{Y|x})$ can decrease the upper bound of $\mathrm{Fit}(\bar{q}_{Y|x},p_{Y|x})$, while reducing their discrepancy can lower the lower bound of $\mathrm{Fit}(\bar{q}_{Y|x},p_{Y|x})$. Consequently, when $\mathrm{Est}(\tilde{q}_{Y|x})$ is large, decreasing $\mathrm{Fit}(\tilde{q}_{Y|x},p_{Y|x})$ may increase $\mathrm{Fit}(\bar{q}_{Y|x},p_{Y|x})$. When we substitute $\bar{q}_{X}$ with $\tilde{q}_{X}$, the introduced bias in the expected learning error is controlled by the estimation error of $\tilde{q}_{X}$, denoted as $\mathrm{Est}_{1}(\bar{q}_{X},\tilde{q}_{X})$.
  4. 4.
    Achieve fitting error minimization by GSC algorithm.
    $\frac{\|\nabla_{\theta}d_{\Omega}(\tilde{q}_{Y|x},f_{x})\|_{2}^{2}}{\lambda_{\max}(A_{x})}\leq\mathrm{Fit}^{2}(\tilde{q}_{Y|x},p_{Y|x})\leq\frac{\|\nabla_{\theta}d_{\Omega}(\tilde{q}_{Y|x},f_{x})\|_{2}^{2}}{\lambda_{\min}(A_{x})},$ (40)
    where $\lambda_{\min}(A_{x})\neq 0$. The model’s fitting error to $\tilde{q}_{Y|x}$, $\mathrm{Fit}(\tilde{q}_{Y|x},p_{Y|x})$, is governed by $\|\nabla_{\theta}d_{\Omega}(\tilde{q}_{Y|x},f_{x})\|_{2}^{2}$, $1/\lambda_{\min}(A_{x})$, and $1/\lambda_{\max}(A_{x})$. Its minimization can be achieved through the GSC algorithm.
  5. 5.
    Increasing the training sample size can reduce the estimation error.
    $\lim_{n\to\infty}\hat{q}_{Y|x}=\tilde{q}_{Y|x}=\bar{q}_{Y|x},\forall x\in\mathcal{X},$ (41)
    $\lim_{n\to\infty}\hat{q}_{X}=\tilde{q}_{X}=\bar{q}_{X},$
    $\lim_{n\to\infty}\mathrm{Est}(\tilde{q}_{Y|x})=0,$
    $\lim_{n\to\infty}\mathrm{Est}(\tilde{q}_{X})=0.$
    As the number of sampling data approaches infinity, any unbiased and consistent estimate becomes the optimal estimate.
  6. 6.
    Increasing the training sample size helps reduce the uncertainty.
    $\lim_{n\to\infty}\mathrm{Var}(q_{Y|x})=0,\forall x\in\mathcal{X}.$ (42)
    When the number of sampling samples approaches infinity, the uncertainty tends to zero.
For a typical deterministic single-label classification with negligible noise, $\tilde{q}_{Y|x}=\bar{q}_{Y|x}=\delta_{y}$, where $y$ is the label associated with $x$. The above upper bound can then be simplified to
$\mathcal{R}_{c}(q,p,\epsilon)\leq\frac{\mathbb{E}_{X\sim\bar{q}_{X}}[\mathrm{Fit}(\tilde{q}_{Y|X},p_{Y|X})^{2}]}{\epsilon^{2}}.$ (43)
Similarly, as the number of sampled instances approaches infinity, the learning risk probability depends solely on the fitting error. This implies that the better the model fits, the better its performance will be in this scenario.
Figure 1: Algorithm derivation logic diagram.
We summarize the algorithm logic of native PD learning and conditional PD learning in Figure 1, where the white box denotes the intermediate quantity of reasoning, whereas the gray box signifies the final processable components. A central aspect of the estimation error is the selection of an unbiased and consistent estimator $\tilde{q}$, which serves as a bridge between the optimal estimator $\bar{q}$ and the model output $p$. Although $\hat{q}$ is theoretically feasible, in practical scenarios, especially in the case of insufficient samples, $\hat{q}$ may not be the best choice. Based on the prior knowledge, we can generate more accurate estimators. If the underlying true distribution is known to be smooth, we can construct $\tilde{q}$ by applying Laplacian smoothing to $\hat{q}$, which often yields smaller estimation error and superior performance. In the field of machine learning, techniques such as label smoothing (Szegedy et al., 2015) and synthetic minority oversampling technique (SMOTE) (Chawla et al., 2002), which introduce a degree of smoothness to the existing data, also yield favorable outcomes. After determining $\tilde{q}$, the subsequent step is to approximate it with the model by minimizing the fitting error $\mathrm{Fit}(\tilde{q},p)$. Optimization of the fitting error, given by $\arg\min_{\theta}\mathrm{Fit}(\tilde{q},p)$, is a typical non-convex problem. Simple gradient-based optimization algorithms do not guarantee a global optimal solution. To address this, we introduce the standard loss function and the GSC algorithm. By using the standard loss function, the GSC algorithm facilitates the approximation of global optimal solutions in fitting error minimization by minimizing both $\|\nabla_{\theta}d_{\Omega}(\tilde{q},f)\|$ and the difference between extreme eigenvalues of $A$. Ultimately, based on the aforementioned algorithm design, the upper and lower bounds on the learning error in the native PD learning and conditional PD learning are derived separately.
Corresponding to the algorithm logic above, the PD learning framework encompasses five principal components: standard loss, fitting model, prior knowledge ($\mathcal{P}$), estimator ($\tilde{q}$), and the GSC algorithm (represented by $\mathcal{A}$), as shown in Figure 2. The estimator yields an approximation of $\bar{q}$, denoted as $\tilde{q}$, leveraging both the sample set $s^{n}$ and prior knowledge $\mathcal{P}$. The standard loss function proposed is a type of loss function that possesses a unique optimal solution and sampling stability. We have demonstrated that under these two conditions, it has a unique mathematical expression and can be represented as a Young–Fenchel loss function, which encompasses various common losses. The fitting model in PD learning is utilized to fit the target distribution $\tilde{q}$ via its Legendre-Fenchel conjugate function. The GSC algorithm is designed to minimize both the gradient norm and the structural error. Within the PD learning framework, we demonstrate that the global optimal solution to non-convex optimization problems, such as fitting error minimization, can be approached using the GSC algorithm.
Figure 2: Basic components of PD learning.
To elucidate PD learning more effectively, we compare it with the conventional deep learning theory framework.
  • Learning objective. The underlying true distribution $q$ or $q_{Y|x}$ represents the learning objectives in PD learning. Within the conventional deep learning theoretical framework, the primary objective is to ascertain the unknown mapping between features and labels.
  • Modeling method. In PD learning, the learning objective is conceptualized as a random variable within the probability simplex. Conversely, within the conventional deep learning theoretical framework, the unknown mapping is represented as a function that processes input features to generate corresponding labels within a defined functional space $\mathcal{F}_{\Theta}$.
  • Optimization objective and its reformulation. The optimization objective in PD learning is the minimization of learning risk probability, as defined in (15). This objective is achieved by controlling the estimation error and the fitting error. In contrast, within the conventional deep learning theoretical framework, the optimization objective is the minimization of expected risk. Expected risk minimization is reformulated as Empirical Risk Minimization (ERM) and generalization error minimization, collectively referred to as Structural Risk Minimization (SRM).
  • Model and its fitting method. In PD learning, neural networks or other models that can implement error backpropagation optimization are utilized to fit the target. This is achieved by employing the Legendre-Fenchel conjugate function of the model’s output. Conversely, in the conventional deep learning theoretical framework, neural networks are typically used to directly fit the target without the intermediary step of the conjugate function.
  • Loss function and its role. In PD learning, the standard loss function is used to facilitate the minimization of the gradient norm. In contrast, within the conventional deep learning theoretical framework, a variety of loss functions may be employed. The minimization of the loss function, in conjunction with the regularization term corresponding to SRM, serves as the optimization objective.
  • Optimization algorithm. The GSC algorithm is used to optimize the fitting error, which can ensure obtaining an approximate global optimal solution for fitting error minimization in the PD learning framework. Gradient-based iterative algorithms are employed to approximate stationary points rather than the global optimum. Within the conventional deep learning theoretical framework, gradient-based iterative algorithms, including GD, SGD, and Adaptive Moment Estimation(Adam), are utilized to achieve SRM. While these algorithms strive to identify the global optimal solution, they do not ensure its attainment.
  • Theoretical guarantee of learning reliability. In PD learning, both upper and lower bounds of learning risk probability are provided and can be accurately computed, providing insights into the mechanisms underlying over-parameterization, non-convex optimization, and other related techniques. Within the conventional deep learning theoretical framework, the existing bounds primarily are upper bounds of generalization error, which are insufficient in explicating over-parameterization and non-convex optimization mechanisms in DNNs. Furthermore, these bounds often incorporate unknown elements, such as the underlying true distribution $q$, which makes their precise computation difficult.
5 Understanding deep learning from the perspective of PD learning
In this section, we analyze and discuss various deep learning techniques within the PD learning framework. We propose that many deep learning techniques can be interpreted from the perspective of the three GSC algorithm implementations discussed in subsubsection 4.4.3. These implementations involve enhancing gradient independence to achieve the Uniform Sampling Assumption and reducing the structural error by increasing the number of parameters upon fulfilling the Uniform Sampling Assumption.
5.1 Enhancing the substantiation of the uniform sampling assumption
During model training process, numerous techniques deliberately or inadvertently introduce randomness to reduce the dependency among parameters and the enhance the independence among the columns of $\nabla_{\theta}f_{x}$. Consequently, we can interpret some existing deep learning techniques from the perspective of fostering the fulfillment of the Uniform Sampling Assumption.
Random initialization
During the parameter initialization phase, the random initialization of parameters in DNNs Glorot and Bengio (2010); He et al. (2015b) ensures that the gradient $\nabla_{\theta}f_{x}$ for an untrained model is randomly assigned and helps to reduce the dependence among the elements of $\nabla_{\theta}f_{x}$. Thus, the random initialization of parameters allows the columns of the gradient matrix $\nabla_{\theta}f_{x}$ to be approximately uniformly sampled from a ball.
Gradient-based optimization
During the model training, randomly selecting samples in each iteration of SGD or Batch Gradient Descent (Batch-GD) leads to a random update of gradient direction, which enhances independence between the columns of $\nabla_{\theta}f_{x}$. This, in turn, enhances their mutual orthogonality according to Lemma 3.4. Therefore, we believe that the stochasticity introduced by SGD facilitates the validity of the uniform sampling assumption, thereby enabling the structural error to be controlled by the model’s parameter size.
Depth of model
From a model structure perspective, if the model possesses an extremely large number of layers, it is justifiable to assume that the majority of the gradients exhibit a high degree of approximate independence. As the number of layers between parameters $\theta_{i}$ and $\theta_{j}$ increases, the dependence between $\nabla_{\theta_{i}}f$ and $\nabla_{\theta_{j}}f$ weakens Balduzzi et al. (2017). Thus, an increase in model depth helps to make the Uniform Sampling Assumption 4.4.3 hold.
Dropout
Dropout Krizhevsky et al. (2012) ensures that the loss remains relatively constant even when specific nodes in the DNN are temporarily discarded. This effectively reduces the influence of these nodes on other parameters, thereby weakening the dependence between $\nabla_{\theta}f_{x}$ and $\nabla_{\theta}f_{x}$ and improving the rationality of the Uniform Sampling Assumption 4.4.3.
5.2 Reducing structural error by increasing parameter count
over-parameterization
As analyzed above, the applications of parameter random initialization, stochasticity inherent in gradient-based iterative algorithms, parameter and feature normalization, and dropout, contribute to validating the Uniform Sampling Assumption 4.4.3 and reducing the sampling ball radius . Based on the conclusion of Theorem 4.7, it can be deduced that the following conclusion holds with probability $1-O(1/|\mathcal{Y}|)$:
$\lambda_{\max}(A_{x})$ $\geq(1-t^{2})^{2}\epsilon^{2},$ (44)
$\lambda_{\max}(A_{x})$ $-\lambda_{\min}(A_{x})\leq 2|\mathcal{Y}|\sqrt{3}t\epsilon^{2},$
where $t=\frac{2\log|\mathcal{Y}|}{|\theta|-1}$. Therefore, increasing the number of parameters $|\theta|$ can effectively manage the upper bound of $G(A_{x})$, and be conducive to reducing the value of $L(A_{x})$. In addition, when $|\theta|\gg 4\epsilon^{2}\log|\mathcal{Y}|$, it follows that $G(A_{x})\approx 0$. According to Theorem 4.5, in this case the fitting error is proportional to the norm of gradients with the rate being $L(A_{x})$. Thus increasing the number of parameters $|\theta|$ while decreasing $\|\nabla_{\theta}(f_{x})_{i}\|$ can reduce both the structural error and fitting error. This underlies the ability of the overparameterized model to approximate global optimal solutions for non-convex optimization through a gradient-based optimization algorithm.
6 Experiments
In this section, we aim to validate the effectiveness of the PD learning framework and the GSC algorithm through three experiments. Firstly, we examine the variations in structural error, gradient norm, and fitting error during the training process of the model to confirm that the deep model’s training adheres to the GSC algorithm’s description. Then, we analyze the convergence curves of models with varying structures as the number of layers increases, discussing the impact of model depth on convergence speed and results.
In our experiments, we utilize the MNIST database Yang et al. (2019) as a case study to visualize how the structural error, gradient norm and fitting error evolve as we increase the network depth across different model architectures. This database has 60k training images and 10k testing images in 10 categories, containing 10 handwritten digits. The model architectures and configuration parameters in the experiment are depicted in Figures 3.
Figure 3: Model architectures and configuration parameters.
The experiments were executed using the following computing environment: Python 3.7, Pytorch 2.2.2, and a GeForce RTX2080Ti GPU.
Table 3: The configuration of Training.
loss function cross-entropy loss #epochs 1
batch size 64 learning rate 0.01
momentum 0.9 data size 28*28
Initially, we set $k=1$ for both Model A and Model B and employed the batch-GD algorithm to train the two models. The training parameters are detailed in Table 3, and the changes in key indicators during the training process are illustrated in Figures 4. The fitting errors of the models and their respective bounds are depicted in Figure 5.
Figure 4: Changes in indicators during the training process.
Figure 5: Changes of bounds during the training process.
Based on Figure 4 and Figure 5, we can obtain the following conclusions:
  1. 1.
    As the training progresses, $U(A_{x})$, $L(A_{x})$ and $G(A_{x})$, apart from a brief initial increase, generally exhibit a stable downward trend. In contrast, the gradient norm $\|\nabla_{\theta}d_{\Omega}(\tilde{q}_{Y|x},f_{x})\|_{2}^{2}$ initially rises and then undergoes sharp oscillations before decreasing towards zero. Even when the training reaches a stable state, the gradient norm still exhibits significant fluctuations.
    Figure 6: Model c with increasing layers
    Figure 7: Model d with increasing blocks
    The diminution of structural error throughout the progression of gradient iterations demonstrates that, as we explained in Paragraph 5.1, the stochasticity inherent in the SGD algorithm is conducive to reducing structural error.
  2. 2.
    As the training progresses, we observe that the variation in fitting error increasingly aligns with the trends of its upper and lower bounds. Concurrently, while the bounds themselves decrease, the gap between them narrows. This phenomenon validates our proposed GSC algorithm, demonstrating that the minimization of the fitting error is achieved by controlling its upper and lower bounds. This also empirically validates the validity of the bounds of fitting error in Theorem 4.5 and Theorem 4.5.
  3. 3.
    Upon comparing the fitting errors, gradient norms, and structural errors of various models, it becomes evident that Models c and d, with a larger number of parameters, exhibit superior fitting performance. Furthermore, Models a and b, which have fewer parameters, display significant discrepancies between their fitting error bounds and the actual fitting errors. In contrast, the fitting errors of Models c and d are nearly identical to their respective upper and lower bounds. Consequently, this finding aligns with our theoretical conclusions regarding over-parameterization.
The experimental results regarding the impact of model depth on fitting error optimization are illustrated in Figure 7. As the number of layers $k$ increases from 0 to 3, the model’s convergence to better parameters is facilitated, resulting in a stable reduction in fitting error. However, beyond $k=3$, further increases in layers do not yield positive effects; Model d does not show significant improvements in convergence speed or final values, while Model c experiences gradient explosion at $k=5$, and the fitting error does not converge to a smaller value. This observation is consistent with our expectations: increasing the number of layers in a model can enhance its fitting capability, but the training difficulty of models tends to increase with their depth. Compared to Model c, Model d, which incorporates residual blocks, exhibits more stable convergence. Nevertheless, at the same number of layers, the fitting error values attained by Model d are higher than those of Model c, a phenomenon worthy of attention.
7 Conclusion
This study introduces a novel theoretical learning framework, PD learning, which provides a comprehensive perspective on various machine learning and statistical tasks. PD learning is characterized by its focus on the true underlying distribution, which is modeled as a random variable within the probability simplex. In order to estimate the underlying distribution, we introduce the concept of learning risk probability as the optimization objective which is bounded by estimation error and fitting error. To address the challenge of non-convexity in optimizing fitting error, we also introduce the GSC algorithm for non-convex optimization objectives, which extends gradient-based optimization methods and establishes an effective bound on fitting error under its application. We illustrate that PD learning provides clearer insights into the unresolved issues in deep learning. The consistency of our framework with existing theories and experimental results, as well as the explanatory power of existing theories and phenomena, demonstrate the correctness and effectiveness of our proposed framework.
Acknowledgments and Disclosure of Funding
This work was supported in part by the National Natural Science Foundation of China under Grant 72171172 and 92367101; the Aeronautical Science Foundation of China under Grant 2023Z066038001; the National Natural Science Foundation of China Basic Science Research Center Program under Grant 62088101; Shanghai Municipal Science and Technology Major Project under Grant 2021SHZDZX0100; Chinese Academy of Engineering, Strategic Research and Consulting Program, under Grant 2023-XZ-65.
Appendix A Proofs
In the appendices, we prove the results stated in the paper and provide necessary technical details and discussions.
A.1 Proof of Lemma 4.1
Firstly, let’s derive the upper bound of $\mathrm{Est}_{1}(\hat{q})$ as follows:
$|\mathbb{E}_{\hat{q}}$ $[\ell]-\mathbb{E}_{\bar{q}}[\ell]|=|\sum_{z\in\mathcal{Z}}(\hat{q}(z)-\bar{q}(z))\ell(z)|$ (45)
$=|\sum_{z\in\mathcal{Z}}(\bar{q}(z)-\hat{q}(z))(\ell(z)-\frac{1}{2})+\frac{1}{2}(\bar{q}(z)-\hat{q}(z))|$
$\leq|\sum_{z\in\mathcal{Z}}(\bar{q}(z)-\hat{q}(z))(\ell(z)-\frac{1}{2})|$
$+|\sum_{z\in\mathcal{Z}}\frac{1}{2}(\bar{q}(z)-\hat{q}(z))|$
$\leq\sum_{z\in\mathcal{Z}}|\bar{q}(z)-\hat{q}(z)||\ell(z)-\frac{1}{2}|$
$\leq\frac{1}{2}\|\hat{q}-\bar{q}\|_{1}.$
Define a function
$g(z)=\begin{cases}1,&\text{if }z\in B\\ 0,&\text{if }z\in\mathcal{Z}/B\end{cases}$
, where $B=\{z|\hat{q}(z)\geq\bar{q}(z)\}$. Note that
$\|\hat{q}-\bar{q}\|_{1}$ $=\sum_{z\in\mathcal{Z}}|\hat{q}(z)-\bar{q}(z)|$
$=\sum_{z\in B}(\hat{q}(z)-\bar{q}(z))+\sum_{z\in\mathcal{Z}/B}(\bar{q}(z)-\hat{q}(z))$
$=\sum_{z\in B}\hat{q}(z)-\sum_{z\in B}\bar{q}(z)+1$
$-\sum_{z\in B}\bar{q}(z)-(1-\sum_{z\in B}\hat{q}(z))$
$=2|\mathbb{E}_{Z\sim\hat{q}}[g(z)]-\mathbb{E}_{Z\sim\bar{q}}[g(z)]|$
$\leq 2\sup_{\ell}|\mathbb{E}_{Z\sim\hat{q}}[\ell(z)]-\mathbb{E}_{Z\sim\bar{q}}[\ell(z)]|.$
So, we have
$\|\hat{q}-\bar{q}\|_{1}=2\sup_{\ell}|\mathbb{E}_{Z\sim\hat{q}}[\ell(z)]-\mathbb{E}_{Z\sim\bar{q}}[\ell(z)]|,$ (46)
where $\ell:\mathcal{Z}\to[0,1]$.
A.2 Proof of Theorem 4.1
Applying lemma 3.6, we obtain the following result:
$\Pr(\|p-q\|_{2}\geq\epsilon)$ $\leq\frac{\mathbb{E}_{q\sim\mathcal{Q}}[\|p-q\|_{2}^{2}]}{\epsilon^{2}}$ (47)
$=\frac{\mathrm{Var}(q)+\|\bar{q}-p\|_{2}^{2}}{\epsilon^{2}},$
where $\mathrm{Var}(q)=\mathbb{E}_{q\sim\mathcal{Q}}[\|\bar{q}-q\|_{2}^{2}]$, $\bar{q}=\mathbb{E}_{q\sim\mathcal{Q}}q$. The final equality is derived from the bias-variance decomposition. It is evident that the right-hand side of the aforementioned inequality is less than 1 only when $\epsilon^{2}\geq\mathrm{Var}(q)+\|\bar{q}-p\|_{2}^{2}$, thereby enabling the control of the learning risk probability.
When $\epsilon^{2}\leq\mathrm{Var}(q)+\|\bar{q}-p\|_{2}^{2}$, according to the Paley-Zygmund Inequality 3.7, we also have
$\Pr(\|p-q\|_{2}\geq\epsilon)$ $=\Pr(\|p-q\|_{2}^{2}\geq\epsilon^{2})$ (48)
$\geq\frac{(\mathbb{E}[\|p-q\|^{2}_{2}]-\epsilon^{2})^{2}}{\mathbb{E}[\|p-q\|^{4}_{2}]}$
$\geq\frac{(\mathrm{Var}(q)+\|\bar{q}-p\|_{2}^{2}-\epsilon^{2})^{2}}{4},$
where the last inequality is derived from the fact that $\|p-q\|_{2}^{2}\leq 2$.
Given that $\mathrm{Lrn}(q,p)=\mathrm{Var}(q)+\|\bar{q}-p\|_{2}^{2}$, it follows that
$\mathcal{R}_{n}(q,p,\epsilon)\leq\frac{\mathrm{Lrn}(q,p)}{\epsilon^{2}},\quad\text{for }\epsilon^{2}>\mathrm{Lrn}(q,p),$ (49)
$\mathcal{R}_{n}(q,p,\epsilon)\geq\frac{(\mathrm{Lrn}(q,p)-\epsilon^{2})^{2}}{4},\quad\text{for }\epsilon^{2}\leq\mathrm{Lrn}(q,p).$
According to the triangle inequality of norms, we have
$(\|p-\tilde{q}\|_{2}-\|\tilde{q}-\bar{q}\|_{2})^{2}\leq(\|p-\bar{q}\|^{2}_{2}\leq\|p-\tilde{q}\|_{2}+\|\tilde{q}-\bar{q}\|_{2})^{2}.$ (50)
A.3 Proof of Theorem 4.2
Since $0\leq\mathcal{R}_{n}(q_{Y|x},p_{Y|x},\epsilon)\leq 1$, according to Lemma 4.1, we have
$|\mathbb{E}_{X\sim\tilde{q}_{X}}[\mathcal{R}_{n}(q_{Y|X},p_{Y|X},\epsilon)]-\mathcal{R}_{c}(q,p,\epsilon)|$ (51)
$\leq\frac{1}{2}\mathrm{Est}_{1}(\tilde{q}_{X}).$
According to Theorem 4.1, we obtain the following result:
$\mathcal{R}_{n}(q_{Y|X},p_{Y|X},\epsilon)\leq\frac{\mathrm{Lrn}(q_{Y|x},p_{Y|x})}{\epsilon^{2}},\text{ for }\epsilon^{2}\leq\mathrm{Lrn}(q_{Y|x},p_{Y|x}),$ (52)
$\mathcal{R}_{n}(q_{Y|X},p_{Y|X},\epsilon)\geq\frac{(\mathrm{Lrn}(q_{Y|x},p_{Y|x})-\epsilon^{2})^{2}}{4},\text{ for }\epsilon^{2}\leq\mathrm{Lrn}(q_{Y|x},p_{Y|x}),$
$\mathrm{Lrn}(q_{Y|x},p_{Y|x})=\mathrm{Var}(q_{Y|x})+\mathrm{Fit}(p_{Y|x},\bar{q}_{Y|x})^{2},$
$|\mathrm{Fit}(\tilde{q}_{Y|x},p_{Y|x})-\mathrm{Est}(\tilde{q}_{Y|x})|\leq\mathrm{Fit}(p_{Y|x},\bar{q}_{Y|x})\leq\mathrm{Fit}(\tilde{q}_{Y|x},p_{Y|x})+\mathrm{Est}(\tilde{q}_{Y|x}).$
A.4 Proof of Theorem 4.3
Assuming $q=q_{0}$, the frequency distribution $\hat{q}$, inferred from the sample $s^{n}$, adheres to a multinomial distribution: $n\hat{q}\sim\operatorname{multinomial}(n,q_{0})$. With the provision of i.i.d. sample data $s^{n}$ and a prior distribution such that $q\sim\mathcal{P}$, the application of Bayes’s theorem facilitates the derivation of the posterior distribution for $q$.
$\mathcal{Q}(q_{0})$ $=\Pr(q_{0}|s^{n})$ (53)
$=\frac{\Pr(\hat{q}|q_{0})\Pr(q_{0})}{\Pr(s)}$
$=\frac{\Pr(\hat{q}|q_{0})\mathcal{P}(q_{0})}{\Pr(\hat{q})}$
$=\mathcal{P}(q_{0})(n!\prod_{z\in\mathcal{Z}}\frac{q_{0}(z)^{n(z)}}{n(z)!})/\sum_{q_{i}\in\Delta}(\mathcal{P}(q_{i})n!\prod_{z\in\mathcal{Z}}\frac{q_{i}(z)^{n(z)}}{n(z)!}),$
where $n(z)=\sum_{i=1}^{n}\mathbf{1}_{\{z\}}(z^{(i)})$.
A.5 Proof of Theorem 4.4
According to the Bayesian theorem and Sanov’s theorem 3.2, we have
$\Pr(q_{0}|s)={\Pr(\hat{q}|q_{0})\Pr(q_{0})}/{\Pr(\hat{q})},$ (54)
$\lim_{n\to\infty}-\frac{1}{n}\log\Pr(\hat{q}|q_{0})=D_{KL}(\hat{q}\|q_{0}).$
Therefore, we obtain
$\lim_{n\to\infty}\Pr(q_{0}|s)=\lim_{n\to\infty}e^{-nD_{KL}(\hat{q}\|q_{0})}\Pr(q_{0})/\Pr(\hat{q}).$ (55)
Since $\Pr(q_{0})/\Pr(\hat{q})$ is finite, then we have
$\lim_{n\to\infty}\mathcal{Q}(q_{0})\sim\delta_{\hat{q}}.$ (56)
Since $\hat{q}$ and $\tilde{q}$ are both the unbiased and consistent estimators of $q$, it follows that
$\lim_{n\to\infty}\mathrm{Var}(q)=0,\\ \lim_{n\to\infty}\hat{q}=\tilde{q}=\bar{q}.$ (57)
A.6 Proof of Theorem 4.5
Since $\ell(\mu,\nu)$ is continuous, differentiable, and has a unique minimum value of zero at $\nu=\mu$, where its gradient with respect to $\nu$ vanishes, i.e., $\nabla_{\nu}\ell(\mu,\mu)=0$, we can express this gradient as
$\nabla_{\nu}\ell(\mu,\nu)=(\nu-\mu)^{\top}g(\mu,\nu).$ (58)
Due to the uniqueness of the minimum value, the expected loss, $\mathbb{E}_{\omega\sim Q}[\ell(\omega,\nu)]$, is strictly convex, and at its minimum $\bar{\omega}$, where $\bar{\omega}=\mathbb{E}_{\omega\sim Q}[\omega]$, the gradient with respect to $\nu$ vanishes, i.e., $\nabla_{\nu}\mathbb{E}_{\omega\sim Q}[\ell(\omega,\bar{\omega})]=0$, which implies that $\sum_{\omega\in\mathcal{Q}}Q(\omega)(\bar{\omega}-\omega)^{\top}g(\omega,\bar{\omega})=0$, where $\omega\sim Q$.
Given that $Q(\omega)$ is an arbitrary distribution function, $g(\omega,\nu)$ must be independent of $\omega$, allowing us to replace $g(\omega,\nu)$ by $g(\nu)$. Since $\mathbb{E}_{\omega\sim Q}[\ell(\omega,\nu)]$ is a strictly convex function, its Hessian with respect to $\nu$ can be derived as:
$\nabla^{2}_{\nu}\mathbb{E}_{\omega\sim Q}[\ell(\omega,\nu)]=\sum_{\omega\in\mathcal{Q}}Q(\omega)[g(\nu)+(\nu-\omega)^{\top}\nabla_{\nu}g(\nu)].$ (59)
Simplifying we get:
$\nabla^{2}_{\nu}\mathbb{E}_{\omega\sim Q}[\ell(\omega,\nu)]=g(\nu)+(\nu-\bar{\omega})\nabla_{\nu}g(\nu)>0.$ (60)
The inequality above is derived from the fact that $\nabla^{2}_{\nu}\mathbb{E}_{\omega\sim Q}[\ell(\omega,\nu)]$ is strictly convex.
When $\nu$ is set to $\bar{\omega}$, we have $\nabla^{2}_{\nu}\mathbb{E}_{\omega\sim Q}[\ell(\omega,\nu)]=g(\bar{\omega})>0$. Since $\bar{\omega}$ can take any value, it follows that for any arbitrary $\nu$, we get $g(\nu)>0$.
Let us define a strictly convex function $\Omega(\nu)$ with its Hessian matrix given by $\nabla^{2}_{\nu}\Omega(\nu)=g(\nu)$. The gradient of the loss function $\ell(\mu,\nu)$ with respect to $\nu$ can be expressed as follows:
$\nabla_{\nu}\ell(\mu,\nu)$ $=\nabla^{2}_{\nu}\Omega(\nu)(\nu-\mu)$ (61)
$=\nabla_{\nu}\nu_{\Omega}^{*}\nu-\nabla_{\nu}\nu_{\Omega}^{*}\mu$
$=\nabla_{\nu}(\langle\nu,\nu_{\Omega}^{*}\rangle-\Omega(\nu))-\mu^{\top}\nabla_{\nu}\nu_{\Omega}^{*}$
$=\nabla_{\nu}(\Omega^{*}(\nu_{\Omega}^{*})-\langle\mu,\nu_{\Omega}^{*}\rangle+h(\mu)).$
Since $\nabla_{\nu}\ell(\mu,\mu)=0$, we find that $h(\mu)=\Omega^{*}(\mu_{\Omega}^{*})-\langle\mu,\mu_{\Omega}^{*}\rangle+c=\Omega(\mu)+c$, where $c$ is a constant. This leads to the following expression for the loss function:
$\ell(\mu,\nu)=\Omega^{*}(\nu_{\Omega}^{*})-\langle\mu,\nu_{\Omega}^{*}\rangle+\Omega(\mu)+c,$ (62)
$\ell(\mu,\mu)=c.$
A.7 Proof of Theorem 4.6
According to the given definition, we have
$\|\nabla_{\theta}d_{\Omega}(\tilde{q},f)\|_{2}^{2}=e^{\top}Ae,$ (63)
where $A:=\nabla_{\theta}f^{\top}\nabla_{\theta}f$,$e:=\tilde{q}-f_{\Omega}^{*}$. Let $A=UDU^{\top}$ denote the eigenvalue decomposition of $A$. Consequently, $D$ is a diagonal matrix whose entries are the eigenvalues of $A$. We then have
$\min_{\|e\|_{2}=K}e^{\top}Ae$ $=\min_{\|e\|_{2}^{2}=K}e^{\top}(UDU^{\top})e$ (64)
$=\min_{\|e\|_{2}^{2}=K}(U^{\top}e)^{\top}D(U^{\top}e)$
$=\min_{\|U^{\top}e\|_{2}^{2}=K}(U^{\top}e)^{\top}D(U^{\top}e)$
$=K\min_{x\in\mathcal{S}}\sum_{i}x_{i}D_{ii},$
where $\mathbb{S}=\{(x_{1},\dots,x_{N})\in\mathbb{R}^{N}\mid x_{i}\geq 0,\leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{i}x_{i}=1\}$. The final step is equivalent to
$K\min_{x\in\mathbb{S}}\sum_{i}x_{i}D_{ii}=K\min_{i}D_{ii}=\lambda_{\min}(A)\|e\|_{2}^{2}.$ (65)
Similarly, we obtain
$\max_{\|e\|_{2}=K}e^{\top}Ae=\lambda_{\max}(A)\|e\|_{2}^{2}.$ (66)
It follows that
$\lambda_{\min}(A)\mathrm{Fit}^{2}(\tilde{q},p)\leq\|\nabla_{\theta}d_{\Omega}(\tilde{q},f)\|_{2}^{2}\leq\lambda_{\max}(A)\mathrm{Fit}^{2}(\tilde{q},p).$ (67)
A.8 Proof of Theorem 4.7
When $|\theta|<|\tilde{q}|$, the matrix $A$ is underdetermined, resulting in $\lambda_{\min}(A)=0$. Leveraging Theorem 4.6, we obtain:
$\lambda_{\max}(A)\|e\|_{2}^{2}\geq\|\nabla_{\theta}d_{\Omega}(\tilde{q},f)\|_{2}^{2}\geq 0.$ (68)
Since each column of $\nabla_{\theta}f$ is uniformly from the ball $\mathcal{B}^{|\theta|}_{\epsilon}:\{x\in\mathbb{R}^{|\theta|},\|x\|_{2}\leq\epsilon\}$, according to Theorem 3.4, then with probability $1-O(1/|\tilde{q}|)$ the following holds:
$A_{ii}=\left\|\nabla_{\theta}f_{i}\right\|^{2}_{2}\geq(1-\frac{2\log|\tilde{q}|}{|\theta|})^{2}\epsilon^{2},\text{ for }i=1,\ldots,|h|,$ (69)
$|A_{ij}|=|\langle\nabla_{\theta}f_{i},\nabla_{\theta}f_{j}\rangle|\leq\frac{\sqrt{6\log|\tilde{q}|}}{\sqrt{|\theta|-1}}\epsilon^{2}\quad\text{ for all }i,j=1,\ldots,|h|,i\neq j.$
Because $A=\nabla_{\theta}f^{\top}\nabla_{\theta}f$ is symmetric positive semidefinite matrix, then we have $\lambda_{\min}(A)\leq A_{ii}\leq\lambda_{\max}(A)$. Then the following holds with with probability $1-O(1/|\tilde{q}|)$ at least
$\lambda_{\max}(A)\geq(1-\frac{2\log|\tilde{q}|}{|\theta|})^{2}\epsilon^{2}.$ (70)
According to Lemma 3.5, there exist $k,k^{\prime}\in\{1,\cdots,|\tilde{q}|\}$
$\lambda_{\max}(A)-A_{kk}$ $\leq R_{k},$ (71)
$A_{k^{\prime}k^{\prime}}-\lambda_{\min}(A)$ $\leq R_{k^{\prime}},$
where $R_{k}=\sum_{1\leq j\leq n,j\neq k}|A_{kj}|$, $R_{k^{\prime}}=\sum_{1\leq j\leq n,j\neq k^{\prime}}|A_{k^{\prime}j}|$. Based on the inequalities (69), we obtain
$R_{k}+R_{k^{\prime}}\leq 2(|\tilde{q}|-1)\frac{\sqrt{6\log|\tilde{q}|}}{\sqrt{|\theta|-1}}\epsilon^{2}.$ (72)
Since $A_{kk}\leq\epsilon^{2}$ and $|\tilde{q}|\leq|\theta|-1$, then we have
$\lambda_{\max}(A)-\lambda_{\min}(A)$ $\leq R_{k}+R_{k^{\prime}}+A_{kk}-A_{k^{\prime}k^{\prime}}$ (73)
$\leq R_{k}+R_{k^{\prime}}+\epsilon^{2}-A_{k^{\prime}k^{\prime}}$
$\leq 2(|h|-1)\frac{\sqrt{6\log|\tilde{q}|}}{\sqrt{|\theta|-1}}\epsilon^{2}+\epsilon^{2}-(1-\frac{2\log|\tilde{q}|}{|\theta|})^{2}\epsilon^{2}$
$\leq 2(|\tilde{q}|-1)\epsilon^{2}\frac{\sqrt{6\log|\tilde{q}|}}{\sqrt{|\theta|-1}}+\frac{4\log|\tilde{q}|}{|\theta|}\epsilon^{2}$
$\leq\frac{2|\tilde{q}|\epsilon^{2}\sqrt{6\log|\tilde{q}|}}{\sqrt{|\theta|-1}}.$