### v8

# Distribution Learning and Its Application in Deep Learning

Abstract
This paper introduces a novel theoretical learning framework, termed probability distribution learning (PD learning). Departing from the traditional statistical learning framework, PD learning focuses on learning the underlying probability distribution, which is modeled as a random variable within the probability simplex. Within this framework, the optimization objective is learning error, which quantifies the posterior expected discrepancy between the model’s predicted distribution and the underlying true distribution, given available sample data and prior knowledge. To optimize the learning error, this paper proposes the necessary conditions for loss functions, models, and optimization algorithms, ensuring that these conditions are all satisfied in real-world machine learning scenarios. Based on these conditions, the non-convex optimization mechanism corresponding to model training can be theoretically resolved. Moreover, the paper provides both model-dependent and model-independent bounds on learning error, offering new insights into the model’s fitting ability and generalization capabilities. Furthermore, the paper applies the PD learning framework to elucidate the mechanisms by which various techniques, including random parameter initialization, over-parameterization, and dropout, influence deep model training. Finally, the paper substantiates the key conclusions of the proposed framework through experimental results.
keywords:
probability distribution , learning theory , deep learning , non-convex optimization , generalization
journal: Neurocomputing
[…]
[1]organization=College of Electronics and Information Engineering,addressline=Tongji University, city=Shanghai, postcode=201804, state=Shanghai, country=China
[…]
[2]organization=National Key Laboratory of Autonomous Intelligent Unmanned Systems,addressline=Tongji University, city=Shanghai, postcode=201210, state=Shanghai, country=China
[…]
[3]organization=Shanghai Research Institute for Intelligent Autonomous Systems,addressline=Tongji University, city=Shanghai, postcode=201210, state=Shanghai, country=China
{graphicalabstract}[Uncaptioned image]
Algorithm derivation logic diagram of PD learnig framework.
{highlights}
Introduce a learning framework aimed at learning the latent distribution.
Theoretically addresses the non-convex optimization mechanism of DNNs
The upper bound of the learning error is the mutual information $I(X;Y)$.
The lower bound of the learning error is the uncertainty of target distribution.
1 Introduction
Along with the remarkable practical achievements of deep learning, several fundamental questions on deep learning remain unanswered within the classical learning theory framework. These include:
  • 1.
    Why can deep neural networks (DNNs) avoid overfitting despite their immense capacity, and how do they surpass shallow architectures?
  • 2.
    How do DNNs overcome the curse of dimensionality, and why do optimization algorithms often succeed in finding satisfactory solutions despite non-convex non-linear, and often non-smooth objectives?
  • 3.
    Which aspects of an architecture affect the performance of the associated models and the mechanisms behind these influences?
Delving into the reasons and factors that govern the performance of DNNs not only enhances the interpretability of DNNs but also facilitates the development of more principled and reliable architecture designs.
Extensive empirical evidence indicates that, in addition to expressivity [1, 2], the generalization ability of DNNs significantly influences prediction performance [3]. Generalization ability, which reflects a model’s predictive capacity on unseen data, is a crucial aspect of learning models and is currently at the core of theoretical research in deep learning. Some classical theory work attributes generalization ability to the utilization of a low-complexity class of hypotheses and proposes various complexity measures to control generalization error. These include the Vapnik-Chervonenkis (VC) dimension [4], Rademacher complexity [5], and Algorithmic Stability [6, 7, 8, 9, 10, 11]. However, recent studies by [12, 13, 14, 15] demonstrate that high-capacity DNNs can still generalize well with genuine labels, yet generalize poorly when trained with random labels. This apparent contradiction with the traditional understanding of overfitting in complex models highlights the need for further investigation [16, 4, 5]. Apart from inconsistencies with experiments, under the existing theoretical framework, expressivity and generalization ability may not solely determine model performance. A typical example is in highly imbalanced binary classification, such as when the ratio of positive to negative samples is 99:1. Suppose a model predicts all inputs as positive; in this case, the overall accuracy of the model is 99%, and the generalization error is 0, but clearly, the model is ineffective. Within the current statistical machine learning framework, the results learned by the model are not unique and depend on the artificially chosen loss function, rather than being entirely determined by the data. This raises deeper questions: If the learning outcome depends on the artificial selection of the loss function, what is the physical significance of the learning outcomes?
Distribution estimation, a cornerstone in various fields such as machine learning, theoretical and applied statistics, information theory, and communication applications, has a rich history. The distribution, described by a probability density/mass function (PDF/PMF), completely characterizes the ”behavior” of a random variable, making distribution estimation fundamental for investigating the properties of a given dataset and for efficient data mining. The goal of distribution estimation is to estimate the true underlying distribution from a given sample set [17, 18, 19, 20]. The majority of machine learning tasks can be framed within the distribution estimation context, with both classification and regression problems mathematically represented by learning the conditional probability distributions between features and labels [21]. Generative learning corresponds to estimating the underlying joint distribution of features to generate new data. Consequently, approaching machine learning and deep learning from the perspective of distribution estimation is a logical and coherent choice, given its inherent suitability for these tasks.
In this paper, we introduce a novel framework, probability distribution learning (PD learning), which can be regarded as a distribution estimation based on machine learning models and objective optimization strategies. A significant portion of machine learning tasks can be conceptualized in the area of distribution learning. Both classification and regression tasks are mathematically characterized by the process of learning the conditional probability distribution of labels based on input features. Generative learning is equivalent to learning the underlying joint distribution of features. Consequently, approaching machine learning and deep learning from a perspective of distribution learning is a logical and coherent choice. The contributions of this paper are stated as follows:
  • 1.
    Unlike traditional machine learning tasks defined in 2.1, the learning objective of PD learning is explicitly defined as the true underlying probability distribution.
  • 2.
    The true underlying probability distribution is modeled as a random variable within the probability simplex, facilitating better integration of sampling data and prior knowledge.
  • 3.
    This paper introduces the learning error as the optimization objective, which quantifies the posterior expected discrepancy between the model’s predicted distribution and the underlying true distribution, given the available sample data and prior knowledge.
  • 4.
    To optimize the learning error, which often constitutes a typical non-convex optimization problem, this paper proposes the necessary conditions for loss functions, models, and optimization algorithms:
    1. (a)
      The loss functions must satisfy the conditions of unique optimum and sampling stability 4.
    2. (b)
      The model must possesses gradients or subgradients with respect to parameters for any input.
    3. (c)
      The optimization algorithms utilized must be capable of simultaneously optimizing gradient norm and structural error, referred to as the gradient structure control (GSC) algorithm 7.
  • 5.
    Under these conditions, this paper proves that although model training is a non-convex optimization problem, its global optimum can be approximated. This conclusion elucidates the non-convex optimization mechanism corresponding to model training in deep learning.
  • 6.
    This paper provides both model-dependent and model-independent bounds on learning error. Model-dependent bounds reveal that as the number of sampling samples increases, the learning error converges to zero. Model-independent bounds indicate that, given finite samples, the greater the mutual information between features and labels, the larger the potential learning error. Additionally, irrespective of the model used, there exists a lower bound on learning error that depends solely on data.
  • 7.
    The equivalence between PD learning and the traditional machine learning paradigm that optimizes the sum of loss expectation and $L_{2}$ regularization term is demonstrated.
  • 8.
    The PD learning framework is applied to elucidate the mechanisms by which techniques such as random parameter initialization, over-parameterization, bias-variance trade-off, and dropout influence deep model training.
  • 9.
    The key conclusions of the proposed framework are substantiated through experimental results.
The remaining sections of this paper are organized as follows: The progress and current status in research are reviewed in Section 2. The fundamental concepts and objects utilized in this paper are defined in Section 3. The details of the proposed method are introduced in Section 4.
  • 1.
    Subsection 4.1 introduces the optimization objectives, related definitions, and the necessary conditions that loss functions, models, and optimization algorithms must satisfy within the PD learning framework.
  • 2.
    Subsection 4.2 and Subsection 4.3 theoretically analyze and prove the properties of standard loss functions and the GSC algorithm, respectively.
  • 3.
    Subsection 4.4 demonstrates that, under the conditions 4.1 outlined in Subsection 4.1, the model can converge to an approximate global optimal solution for any fitting target, irrespective of whether the optimization objective is a convex function of the parameters.
  • 4.
    Subsection 4.5 establishes that, employing unbiased and consistent estimators, both the estimation error and uncertainty diminish to zero as the sample size increases.
  • 5.
    Subsection 4.6 provides both model-dependent and model-independent bounds on the learning error, offering a comprehensive understanding of the model’s performance.
  • 6.
    Subsection 4.7 summarizes the algorithmic logic of PD learning and contrasts it with the classical machine learning paradigms.
The application of PD learning in understanding the questions that remain unanswered within deep learning is illustrated in Section 5. The experimental settings and results are presented in Section 6. The work summary and outlook of this article can be found in Section 7. All the proofs are detailed in A.
2 Literature review
In this section, we present an overview of learning theory, modern methods for analyzing deep learning, and distribution learning methods.
2.1 Learning theory
We commence by offering a concise summary of the classical mathematical and statistical theory underlying machine learning tasks and algorithms, which, in their most general form, can be formulated as follows [22].
Definition 2.1 (Learning task)
In a learning task, one is provided with a loss function $\ell:\mathcal{M}(\mathcal{X},\mathcal{Y})\times\mathcal{Z}\to\mathbb{R}$ and training data $s^{n}=\{z^{(i)}\}_{i=1}^{n}$ which is generated by independent and identically distributed (i.i.d.) sampling according to the unknown true distribution $Z\sim q$, where $Z$ are random variables that take values in $\mathcal{Z}$. The objective is to construct a learning algorithm $\mathcal{A}$ that utilizes training data $s$ to find a model $f_{s}$ that performs well on the entire data, where $f_{s}=\mathcal{A}(s)\in\mathcal{F}$, the hypothesis set $\mathcal{F}\subset\mathcal{M}(\mathcal{X},\mathcal{Y})$, and the mapping $\mathcal{A}:s^{n}\to\mathcal{F}$. Here, performance is measured via $\mathbb{E}_{Z}[\ell(f_{s},z)]$.
To analyze the influencing factors of learning, the classical statistical machine learning framework introduces the optimization error ($e_{\text{opt}}$), generalization error ($e_{\text{gen}}$), and approximation error ($e_{\text{approx}}$).
The optimization error $e_{\rm opt}$ is primarily influenced by the algorithm $\mathcal{A}$ that is used to find the model $f_{s}$ in the hypothesis set $\mathcal{F}$ for given training data $s$. By far, the gradient-based optimization methods are the most prevalent method to address this issue, leveraging their ability to precisely and efficiently compute the derivatives by automatic differentiation in a lot of widely used models. Owing to the hierarchical structure of neural networks (NNs), the back-propagation algorithm has been proposed to efficiently calculate the gradient with respect to the loss function [23, 24, 25, 26]. Stochastic gradient descent (SGD) [27], is a typical gradient-based optimization method used to identify the unique minimum of a convex function [28]. In scenarios where the objective function is non-convex, SGD generally converges to a stationary point, which may be either a local minimum, a local maximum, or a saddle point. Nonetheless, there are indeed conditions under which convergence to local minima is guaranteed [29, 30, 31, 32].
The approximation error, $e_{\text{approx}}$, describes the model’s fitting capability. The universal approximation theorem [33, 34, 35, 1] states that neural networks with only 2 layers can approximate any continuous function defined on a compact set up to arbitrary precision. This implies that the $e_{\rm approx}$, decreases as the model complexity increases.
The generalization error, $e_{\rm gen}(\mathcal{F})$, is defined as the difference between the true risk and the empirical risk. To bound the generalization error, Hoeffding’s inequality [36] alongside the union bound directly provides the PAC (Probably Approximately Correct) bound [37]. The uniform convergence bound represents a classic form of PAC generalization bound, which relies on metrics of complexity for the set $\mathcal{F}$, including the VC-dimension [4] and the Rademacher complexity [5]. Such bounds are applicable to all the hypotheses within $\mathcal{F}$. A potential drawback of the uniform convergence bounds is that they are independent of the learning algorithm, i.e., they do not take into account the way the hypothesis space is explored. To tackle this issue, algorithm-dependent bounds have been proposed to take advantage of some particularities of the learning algorithm, such as its uniform stability [6] or robustness [38].
However, the available body of evidence suggests that classical tools from statistical learning theory alone, such as uniform convergence, algorithmic stability, or bias-variance trade-off may be unable to fully capture the behavior of DNNs [12, 39]. The classical trade-off highlights the need to balance model complexity and generalization. While simpler models may suffer from high bias, overly complex ones can overfit noise (high variance). Reconciling the classical bias-variance trade-off with the empirical observation of good generalization, despite achieving zero empirical risk on noisy data using DNNs with a regression loss, presents a significant challenge [40]. Generally speaking, in deep learning, over-parameterization and perfectly fitting noisy data do not exclude good generalization performance. Moreover, empirical evidence across various modeling techniques, including DNNs, decision trees, random features, and linear models, indicates that test error can decrease even below the sweet-spot in the U-shaped bias-variance curve when further increasing the number of parameters [15, 41, 42]. Apart from inconsistencies with experiments, under the existing theoretical framework, the upper bound of the expected risk may be too loose. In other words, empirical risk and generalization ability may not completely determine model performance. A typical example is in highly imbalanced binary classification, such as when the ratio of positive to negative samples is 99:1. Suppose a model predicts all inputs as positive; in this case, the overall accuracy of the model is 99%, and the generalization error is 0, but clearly, the model is ineffective. Within the Definition 2.1, the result learned by the model $f_{s}$ is not completely deterministic and unique; it depends on the artificially chosen loss function, rather than being entirely determined by the data. This raises deeper questions: If the learning outcome depends on the artificial selection of the loss function, to what objects in the real physical world does the learned result correspond?
2.2 Deep learning theory
Deep learning poses several mysteries within the conventional learning theory framework, including the outstanding performance of overparameterized neural networks, the impact of depth in deep architectures, the apparent absence of the curse of dimensionality, the surprisingly successful optimization performance despite the non-convexity of the problem. Theoretical insights into these problems from prior research have predominantly followed two main avenues: expressivity (fitting ability) and generalization ability.
Essentially, expressivity is characterized by the capability of DNNs to approximate any function. The universal approximation theorem [35, 43] states that a feedforward network with any ”squashing” activation function $\sigma(\cdot)$, such as the logistic sigmoid function, can approximate any Borel measurable function $f(x)$ with any desired non-zero amount of error, provided that the network is given enough hidden layer size $n$. Universal approximation theorems have also been proved for a wider class of activation functions, which includes the now commonly used ReLU [1]. Generally speaking, the greater the number of parameters a model possesses, the stronger its fitting capability. Moreover, research has demonstrated that extremely narrow neural networks can retain universal approximativity with an appropriate choice of depth [44, 45]. More precisely, enhancing the depth of a neural network architecture exponentially increases its expressiveness [46, 47].
Generalization ability represents a model’s predictive ability on unseen data. Increasing the depth of a neural network architecture naturally yields a highly overparameterized model, whose loss landscape is known for a proliferation of local minima and saddle points [48]. The fact that sometimes (for some specific models and datasets), the generalization ability of a model paradoxically increases with the number of parameters, even surpassing the interpolation threshold—where the model fits the training data perfectly. This observation is highly counterintuitive since this regimen usually corresponds to over-fitting. DNNs show empirically that after the threshold, sometimes the generalization error tends to descend again, showing the so called Double-Descent generalization curve, and reaching a better minimum. In order to address this question, researchers have tried to focus on different methods and studied the phenomena under different names, e.g., over-parametrization [49, 50, 51, 52], double descent [53, 54, 55, 56], sharp/flat minima [57, 58, 59, 60, 61, 62], and benign overfit [63, 64, 65, 66]. In this paper, we review the related literature and organize it into the following six categories.
  1. 1.
    Complexity-based methods. Conventional statistical learning theory has established a series of upper bounds on the generalization error (generalization bounds) based on the complexity of the hypothesis space, such as VC-dimension [67, 4] and Rademacher complexity [5]. Usually, these generalization bounds suggest that controlling the model size can help models generalize better. However, overparameterized DNNs make the generalization bounds vacuous.
  2. 2.
    Algorithmic Stability. Algorithmic Stability (e.g., uniform stability, hypothesis stability) is straightforward: if an algorithm fits the dataset without being overly dependent on particular data points, then achieving small empirical error implies better generalization [6, 7, 8, 9, 10, 11]. Algorithmic Stability simply cares about the fact that the algorithm choice should not change too much and is stable enough if we slightly change the dataset.
  3. 3.
    PAC-Bayesian. PAC-Bayesian (or simply ”PAC-Bayes”) theory, which integrates Bayesian statistic theories and PAC theory for stochastic classifiers, has emerged as a method for obtaining some of the tightest generalization bounds [68, 69]. It emerged in the late 1990s, catalyzed by an early paper [70], and shepherded forward by McAllester [71, 72]. McAllester’s PAC-Bayesian bounds are empirical bounds, in the sense that the upper bound only depends on known computable quantities linked to the data. Over the years, many PAC-Bayes bounds were developed in many ever tighter variants, such as the ones for Bernoulli losses [73, 74, 75], exhibiting fast-rates given small loss variances [76, 77], data-dependent priors [78], and so on. We refer the reader to [79] for excellent surveys.
  4. 4.
    The stochasticity of stochastic optimization algorithms as an implicit regularization. Due to the non-linear dependencies within the loss function relative to the parameters, the training of DNNs is a quintessential non-convex optimization problem, fraught with the complexities of numerous local minima. First-order optimization methods, such as SGD, have become the workhorse behind the training of DNNs. Despite its simplicity, SGD also enables high efficiency in complex and non-convex optimization problems [80]. It is widely believed that the inherent stochasticity of stochastic optimization algorithms performs implicit regularization, guiding parameters towards flat minima with lower generalization errors [81, 82, 48, 83]. To identify the flat minima, numerous mathematical definitions of flatness exist, with a prevalent one being the largest eigenvalue of the Hessian matrix of the loss function with respect to the model parameters on the training set, denoted as $\lambda_{\max}$ [84, 85, 58].
  5. 5.
    The role of over-parameterization in non-convex optimization. As for the impact of parameter quantity on optimization algorithms, it is already known in the literature that DNNs in the infinite width limit are equivalent to a Gaussian process [86, 87, 88, 89, 90]. The work [91] elucidates that the evolution of the trainable parameters in continuous-width DNNs during training can be captured by the neural tangent kernel (NTK). Recent work has shown that with a specialized scaling and random initialization, the parameters of continuous width two-layer DNNs are restricted in an infinitesimal region around the initialization and can be regarded as a linear model with infinite dimensional features [49, 92, 93, 94]. Since the system becomes linear, the dynamics of gradient descent (GD) within this region can be tracked via properties of the associated NTK and the convergence to the global optima with a linear rate can be proved. Later, the NTK analysis of global convergence is extended to multi-layer neural nets [95, 96, 97]. An alternative research direction attempts to examine the infinite-width neural network from a mean-field perspective [98, 99, 100]. The key idea is to characterize the learning dynamics of noisy SGD as the gradient flows over the space of probability distributions of neural network parameters. When the time goes to infinity, the noisy SGD converges to the unique global optimum of the convex objective function for the two-layer neural network with continuous width. NTK and mean-field methods provide a fundamental understanding on training dynamics and convergence properties of non-convex optimization in infinite-width neural networks.
  6. 6.
    Information-based methods. Information-theoretic generalization bounds have been empirically demonstrated to be effective in capturing the generalization behavior of Deep Neural Networks (DNNs) [101, 102]. The original information-theoretic bound introduced by [103] has been extended or improved in various ways, such as the chaining method [104, 105, 106], the random subset and individual technique [101, 107, 108, 109, 110]. The information bottleneck principle [111, 112, 113], from the perspective of the information change in neural networks, posits that for better prediction, the neural network learns useful information and filters out redundant information. Beyond supervised learning, information-theoretic bounds are also useful in a variety of learning scenarios, including meta-learning [114], semi-supervised learning [115], transfer learning [116], among others.
While there is yet no clarity in understanding why, for some algorithms and datasets, increasing the number of parameters actually improves generalization [117]. The accumulating evidence and existing theories indicate that the exceptional performance of DNNs is attributed to a confluence of factors, including their architectural design, optimization algorithms, and even training data. Elucidating the mechanisms underlying deep learning remains an unresolved issue at present.
2.3 Distribution estimation
In statistical data analysis, determining the probability density function (PDF) or probability mass function (PMF) from a limited number of random samples is a common task, especially when dealing with complex systems where accurate modeling is challenging. Generally, there are three main methods for this problem: parametric, semi-parametric, and non-parametric models.
Parametric models assume that the data follows a specific distribution with unknown parameters. Examples include the Gaussian model, Gaussian Mixture models (GMMs), and distributions from the generalized exponential family. A major limitation of these methods is that they make quite simple parametric assumptions, which may not be sufficient or verifiable in practice [118]. Non-parametric models estimate the PDF/PMF based on data points without making assumptions about the underlying distribution [18]. Examples include the histogram [119, 120, 121] and kernel density estimation (KDE) [122, 123, 124]. While these methods are more flexible, they may suffer from overfitting or poor performance in high-dimensional spaces due to the curse of dimensionality. Semi-parametric models [19] strike a balance between parametric and non-parametric methods. They do not assume the a prior-shape of the PDF/PMF to estimate as non-parametric techniques. However, unlike the non-parametric methods, the complexity of the model is fixed in advance, in order to avoid a prohibitive increase of the number of parameters with the size of the dataset, and to limit the risk of overfitting. Finite mixture models are commonly used to serve this purpose.
Statistical methods for determining the parameters of a model conditional on the data, including the standard likelihood method [17, 125, 20], M-estimators [126, 127], and Bayesian methods [128, 129, 130], have been widely employed. However, traditional estimators struggle in high-dimensional space: because the required sample size or computational complexity often explodes geometrically with increasing dimensionality, they have almost no practical value for high-dimensional data.
Recently, neural network-based methods have been proposed for PDF/PMF estimation and have shown promising results in problems with high-dimensional data points such as images. There are mainly two families of such neural density estimators: auto-regressive models [131, 132] and normalizing flows [133, 134]. Autoregression-based neural density estimators decompose the density into the product of conditional densities based on probability chain rule $p(x)=\prod_{i}p(x_{i}|x_{1:i-1})$. Each conditional probability $p(x_{i}|x_{1:i-1})$ is modeled by a parametric density (e.g., Gaussian or mixture of Gaussian), and the parameters are learned by neural networks. Density estimators based on normalizing flows represent $x$ as an invertible transformation of a latent variable $z$ with a known density, where the invertible transformation is a composition of a series of simple functions whose Jacobian is easy to compute. The parameters of these component functions are then learned by neural networks.
3 Preliminaries
3.1 Notation
  • 1.
    The set of real numbers is denoted by $\mathbb{R}$, and the cardinality of set $\mathcal{Z}$ is denoted by $|\mathcal{Z}|$.
  • 2.
    Random variables are denoted using upper case letters such as $X,Y$ which take the values in set $\mathcal{X}$, and $\mathcal{Y}$, respectively.
  • 3.
    The expectation of a random variable $X$ is denoted by $\mathbb{E}{[X]}$ or $\mathbb{E}_{X}[X]$, with $\mathbb{E}_{X\sim p}[X]$ when the distribution of $X$ is specified. And we define $\mathbb{E}_{X\sim p}[f(X)]=\sum_{x\in\mathcal{X}}p(X=x)f(x)$, where $\mathcal{X}$ is the value space of $X$.
  • 4.
    The space of models, which is a set of functions endowed with some structure, is represented by $\mathcal{F}_{\Theta}=\{f_{\theta}:\theta\in\Theta\}$, where $\Theta$ denotes the parameter space. Similarly, we define the hypothesis space with feature $x$ as $\mathcal{F}_{\Theta,x}:=\{f_{\theta,x}:\theta\in\Theta\}$.
  • 5.
    The sub-differential is defined as $\partial f(x)=\{s\in\mathbb{R}^{n}|f(y)-f(x)-\langle s,y-x\rangle\geq 0,\forall y\in\mathrm{dom}(f)\}$ (see e.g., [135],Page 27). If $f(x)$ is smooth and convex, the sub-differential $\partial f(x)$ reduces to the usual gradient, i.e., $\partial f(x)=\{\nabla f(x)\}$.
  • 6.
    The Legendre-Fenchel conjugate of a function $\Omega$ is denoted by $\Omega^{*}(f):=\sup_{p\in\mathrm{dom}(\Omega)}\langle p,f\rangle-\Omega(p)$. By default, $\Omega$ is a continuous strictly convex function, and its gradient with respect to $p$ is denoted by $p_{\Omega}^{*}$. When $\Omega(\cdot)=\frac{1}{2}\|\cdot\|_{2}^{2}$, we have $p=p_{\Omega}^{*}$.
  • 7.
    The Fenchel-Young loss $d_{\Omega}\colon\mathrm{dom}(\Omega)\times\mathrm{dom}(\Omega^{*})\to\mathbb{R}^{+}$ generated by $\Omega$ is defined as [136]:
    $d_{\Omega}(q;f):=\Omega(q)+\Omega^{*}(f)-\langle q,f\rangle,$ (1)
    where $\Omega^{*}$ denotes the conjugate of $\Omega$.
  • 8.
    The maximum and minimum eigenvalues of matrix $A$ are denoted by $\lambda_{\max}(A)$ and $\lambda_{\min}(A)$, respectively.
  • 9.
    To improve clarity and conciseness, we provide the following definitions of symbol:
    $q_{Y|x}(y):=q_{Y|X}(y|x),$ (2)
    $q_{Y|x}:=(q_{Y|x}(y^{(1)}),\cdots,q_{Y|x}(y^{(|\mathcal{Y}|)}))^{\top},$
    $\bar{q}:=\mathbb{E}_{q\sim\mathcal{Q}}[q],$
    $\bar{q}_{X}:=\mathbb{E}_{q_{X}\sim\mathcal{Q}_{X}}[q_{X}],$
    $\bar{q}_{Y|x}:=\mathbb{E}_{q_{Y|x}\sim\mathcal{Q}_{Y|x}}[q_{Y|x}],$
    $\mathrm{Var}(q_{Y|x}):=\mathbb{E}_{q_{Y|x}\sim\mathcal{Q}_{Y|x}}[\|q_{Y|x}-\bar{q}_{Y|x}\|_{2}^{2}],$
    where $q_{X}(x)$ and $q_{Y|X}(y|x)$ represent the marginal and conditional PMFs/PDFs, respectively, the conditional distribution $q_{Y|x}$ is also modeled as a random variable.
3.2 Lemmas
The theorems in the subsequent sections rely on the following lemmas.
Lemma 3.1 (Properties of Fenchel-Young losses)
The following are the properties of Fenchel-Young losses [136].
  1. 1.
    $d_{\Omega}(q,f)\geq 0$ for any $q\in\mathrm{dom}(\Omega)$ and $f\in\mathrm{dom}(\Omega^{*})$. If $\Omega$ is a lower semi-continuous proper convex function, then the loss is zero iff $f\in\partial\Omega(q)$. Furthermore, when $\Omega$ is strictly convex, the loss is zero iff $f=q_{\Omega}^{*}$.
  2. 2.
    $d_{\Omega}(q,f)$ is convex with respect to $f$, and its subgradients include the residual vectors: $f_{\Omega^{*}}^{*}-q\in\partial_{f}d_{\Omega}(q;f)$. If $\Omega$ is strictly convex, then $d_{\Omega}(q,f)$ is differentiable and $\nabla_{f}d_{\Omega}(q,f)=f_{\Omega^{*}}^{*}-q$. If $\Omega$ is strongly convex, then $d_{\Omega}(q,f)$ is smooth, i.e., $\nabla_{f}d_{\Omega}(q,f)$ is Lipschitz continuous.
Throughout this paper, in line with practical applications and without affecting the conclusions, we assume that $\Omega$ is a proper, lower semi-continuous (l.s.c.), and strictly convex function or functional. Consequently, the Fenchel-Young loss attains zero if and only if $q=f_{\Omega^{*}}^{*}$.
Lemma 3.2 (Sanov’s theorem)
$\lim_{n\to\infty}-\frac{1}{n}\log\Pr(\hat{q}|q)=D_{KL}(\hat{q}\|q),$ (3)
where $D_{KL}(\hat{q}\|q)$ is Kullback-Leibler (KL) divergence [137].
Lemma 3.3 (Donsker-Varadhan representation)
Let $p$ and $q$ be distributions on a common space $\mathcal{X}$. Then
$D_{KL}(p\|q)=\sup_{g}\left\{\mathbb{E}_{X\sim p}[g(X)]-\log\mathbb{E}_{X\sim q}\left[e^{g(X)}\right]\right\},$
where the supremum is taken over measurable functions $g:\mathcal{X}\rightarrow\mathbb{R}$ such that $\mathbb{E}_{X\sim q}\left[e^{g(X)}\right]<\infty$ [138].
Lemma 3.4
Suppose that we sample $n$ points $x^{(1)},\ldots,x^{(n)}$ uniformly from the unit ball $\mathcal{B}^{m}_{1}:\{x\in\mathbb{R}^{m},\|x\|_{2}\leq 1\}$. Then with probability $1-O(1/n)$ the following holds [139, Theorem 2.8]:
$\left\|x^{(i)}\right\|_{2}\geq 1-\frac{2\log n}{m},\text{ for }i=1,\ldots,n.$ (4)
$|\langle x^{(i)},x^{(j)}\rangle|\leq\frac{\sqrt{6\log n}}{\sqrt{m-1}}\text{ for all }i,j=1,\ldots,i\neq j.$
Lemma 3.5 (Gershgorin’s circle theorem)
Let $A$ be a real symmetric $n\times n$ matrix, with entries $a_{ij}$. For $i=\{1,\cdots,n\}$ let $R_{i}$ be the sum of the absolute value of the non-diagonal entries in the $i$-th row: $R_{i}=\sum_{1\leq j\leq n,j\neq i}|a_{ij}|$. For any eigenvalue $\lambda$ of $A$, there exists $i\in\{1,\cdots,n\}$ such that
$|\lambda-a_{ii}|\leq R_{i}.$ (5)
Lemma 3.6 (Pinsker’s inequality)
$D_{KL}(p\|q)\geq\frac{1}{2\ln 2}\|p-q\|_{1}^{2},$ (6)
where $p$ and $q$ be two distribution [140].
3.3 Setting and basics
The random pair $Z=(X,Y)$ follows the distribution $q$ and takes values in the product space $\mathcal{Z}=\mathcal{X}\times\mathcal{Y}$, where $\mathcal{X}$ represents the input feature space and $\mathcal{Y}$ denotes a finite set of labels. The training data, denoted by an $n$-tuple $s^{n}:=\{z^{(i)}\}_{i=1}^{n}=\{(x^{(i)},y^{(i)})\}_{i=1}^{n}$, is composed of i.i.d. samples drawn from the unknown true distribution $q$. Given the available information, the true underlying distribution is undetermined. Consequently, we model $q$ as a random variable in the space $\Delta$. Let $\mathcal{Q}^{(s^{n},\mathcal{P})}$ (abbreviated as $\mathcal{Q}$) be the posterior distribution of $q$, conditional on the i.i.d. sample set $s^{n}$ and the prior distribution $\mathcal{P}$. We define $\bar{q}=\mathbb{E}_{q\sim\mathcal{Q}}q$. Furthermore, let $\tilde{q}$ denote an unbiased and consistent estimator of $\bar{q}$. In the absence of prior knowledge about $q$, the Monte Carlo method (MC) is commonly utilized. This method estimates $q$ based on the event frequencies, denoted by $\hat{q}(s^{n}):=\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\{z\}}(z^{(i)})$. In this article, we refer to $\hat{q}$ as the frequency distribution estimator to distinguish it from the general estimator $\tilde{q}$. We also assume that $\tilde{q}$ is a better estimator relative to $\hat{q}$. Real-world distributions are usually smooth, which inspires the common practice of smoothing $\hat{q}$ to construct $\tilde{q}$ to approximate the true distribution $q$. These smoothing techniques effectively integrate prior knowledge and are often advantageous, as exemplified by classical Laplacian smoothing. For simplicity, when the sample set is given, we abbreviate the results estimated by the estimator $\tilde{q}(s^{n})$ and $\hat{q}(s^{n})$ as $\tilde{q}$ and $\hat{q}$, respectively.
It is important to note that, given the sample set $s^{n}$ and the prior distribution $\mathcal{P}$, calculating $\bar{q}$ and $\mathcal{Q}$ is feasible. The specific calculation methods are detailed in Theorem 4.2. In this article, we utilize the notation $f_{\theta}$ (hereinafter abbreviated as $f$) to denote the model characterized by the parameter vector $\theta$. Additionally, $f_{\theta,x}$ (abbreviated as $f_{x}$) represents the model with a parameter vector $\theta$ and a specific input $x$. For ease of reference, we list the symbol variables defined above in Table 1.
Table 1: Definitions and abbreviations of key variables.
$q$ the true underlying distribution which is modeled as a random variable in the space $\Delta$.
$\mathcal{P}$ the prior distribution of $q$.
$s^{n}$($s$) i.i.d. samples drawn from the unknown true distribution $q$.
$\mathcal{Q}^{(s^{n},\mathcal{P})}$ ($\mathcal{Q}$) the posterior distribution of $q$, conditional on the i.i.d. sample set $s^{n}$ and the prior distribution $\mathcal{P}$.
$\hat{q}(s^{n})$($\hat{q}$) the frequency distribution estimator.
$\tilde{q}(s^{n})$( $\tilde{q}$) an unbiased and consistent estimator of $\bar{q}$.
$f_{\theta}$($f$) the model characterized by the parameter vector $\theta$.
4 Methods
This section provides a detailed introduction to the PD learning framework.
4.1 Problem formulation
This subsection introduces the optimization objectives, related definitions, and the necessary conditions that loss functions, models, and optimization algorithms must satisfy within the PD learning framework.
We refer to the tasks with learning objectives of $q$ and $q_{Y|X}$ as native PD learning and conditional PD learning, respectively. Since the learning target is modeled as a random variable, the learning errors in native PD learning and conditional PD learning are defined as follows:
$L_{n}(q,p)$ $=\mathbb{E}_{q\sim\mathcal{Q}}[\|q-p\|_{2}^{2}],$ (7)
$L_{c}(q,p)$ $=\mathbb{E}_{q_{X}\sim\mathcal{Q}_{X}}\mathbb{E}_{X\sim q_{X}}L_{n}(q_{Y|X},p_{Y|X}),$
where $q_{X}$, $q_{Y|X}$, and $q$ are modeled as random variables. We thus formulate PD learning in two situations:
  • 1.
    Native PD learning
    $\min_{\theta}L_{n}(q,p)\,\ \mathrm{s.t.}\,\ s=\{z_{i}\}_{i=1}^{n},$ (8)
    where $p=(f_{\theta})_{\Omega^{*}}^{*}$, ϵ+italic-$\epsilon\in\mathbb{R}^{+}$.
  • 2.
    Conditional PD learning
    $\min_{\theta}L_{c}(q,p)\,\ \mathrm{s.t.}\,\ s=\{z_{i}\}_{i=1}^{n},$ (9)
    where $p_{Y|x}=(f_{\theta,x})_{\Omega^{*}}^{*}$, ϵ+italic-$\epsilon\in\mathbb{R}^{+}$.
In PD learning, we utilize the Legendre-Fenchel conjugates of $f_{\theta}$ and $f_{\theta,x}$, denoted as $p$ and $p_{Y|x}$, to approximate the $q$ and $q_{Y|x}$. Native PD learning can be considered as a specific case of conditional PD learning when the feature variable $X$ is fixed, i.e., $X\sim\delta_{x}$.
We list the definitions used in the PD learning framework as follows:
  1. 1.
    Fitting error: This error measures the performance of the model in fitting the target $\tilde{q}$, denoted by $\mathrm{Fit}(\tilde{q},p)=\|p-\tilde{q}\|_{2}^{2}$.
  2. 2.
    Uncertainty: We refer the variances $\mathrm{Var}(q)$ and $\mathrm{Var}(q_{Y|x})$ as the uncertainties of $q$ and $q_{Y|x}$ respectively. They are independent of the models $f$ and $f_{x}$ and depend solely on the sample set and prior knowledge.
  3. 3.
    Estimation error: This error quantifies the distance between $\bar{q}$ and the corresponding provided estimate, given by $\mathrm{Est}(\tilde{q})=\|\tilde{q}-\bar{q}\|_{2}^{2}$. The estimation error arises from the finite sample size and serves as a model-independent metric that characterizes the estimator. Since conditional PD learning involves estimation errors in both the feature and label dimensions, we formally define the label estimation error as $\mathrm{Est}(\tilde{q}_{Y|x})$ and the feature estimation error as $\mathrm{Est}(\tilde{q}_{X})$.
  4. 4.
    Standard loss function: A loss function $\ell(\tilde{q},p):\Delta\times\Delta\to\mathbb{R}^{+}$ is regarded as a standard loss if it meets the following conditions :
    • (a)
      Unique optimum. The loss function is strictly convex, ensuring that it has a unique minimum when predicting the correct probability $\tilde{q}$, i.e., $p=\tilde{q}$. The unique optimum condition guarantees that the standard loss function is effective in directing the model towards the desired fitting target.
    • (b)
      Sampling stability. The expected loss, $\mathbb{E}_{q}\ell(q,p)$, achieves its minimum value when $p=\mathbb{E}_{q}[q]$, expressed as $\mathbb{E}_{q}[q]=\arg\min_{p}\mathbb{E}_{q}[\ell(q,p)]$. The sampling stability condition ensures the consistency of optimization results across different optimization methods, such as SGD, batch-GD, and GD. In other words, minimizing the loss function on the entire dataset yields results consistent with those obtained by minimizing the expected value of the loss function under random sampling.
  5. 5.
    Structural matrix: We define $A:=\nabla_{\theta}f^{\top}\nabla_{\theta}f$ and $A_{x}:=\nabla_{\theta}f_{x}^{\top}\nabla_{\theta}f_{x},$ as the structural matrices corresponding to the model $f$ and $f_{x}$.
  6. 6.
    Structural error: we define the structural error as follows:
    $S(\alpha,\beta,\gamma,A)=\alpha G(A)+\beta U(A)+\gamma L(A),$ (10)
    where $U(A)=-\log\lambda_{\min}(A)$, $L(A)=-\log\lambda_{\max}(A)$, $G(A)=U(A)-L(A)$, $\alpha$, $\beta$ and $\gamma$ are all positive real numbers and signify the weights associated with $G(A)$, $U(A)$ and $L(A)$, respectively.
  7. 7.
    GSC algorithm: We refer to the optimization algorithm capable of simultaneously reducing both the gradient norm and the structural error as the GSC algorithm . The GSC algorithm for native PD learning is mathematically formulated as follows:
    $\min_{\theta}\{\log\|\nabla_{\theta}\ell(\tilde{q},p)\|_{2}^{2}+S(\alpha,\beta,\gamma,A)\}.$ (11)
    Similarly, the GSC algorithm under conditional PD learning is formulated as
    $\min_{\theta}\mathbb{E}_{X}[\log\|\nabla_{\theta}\ell(\tilde{q}_{Y|X},p_{Y|X})\|_{2}^{2}+S(\alpha,\beta,\gamma,A_{X})].$ (12)
The loss function, model and optimization algorithm within the PD learning framework must satisfy the following conditions :
  • 1.
    Loss function. The loss function $\ell(\tilde{q},p):\Delta\times\Delta\to\mathbb{R}^{+}$ is standard loss function.
  • 2.
    Model. For any $\theta\in\Theta$, $f_{\theta}$ and $f_{\theta,x}$ possess gradients or subgradients. This condition ensures that we can search for the model’s stationary points using gradient or subgradient optimization algorithms.
  • 3.
    optimization algorithm. GSC algorithm is applied.
Based on the aforementioned definitions, employing Bias-Variance Decomposition, we obtain the following theorem.
Theorem 4.1
$L_{n}(q,p)$ $=\mathrm{Var}(q)+\mathrm{Fit}(q,p),$ (13)
$L_{c}(q,p)$ $=\mathbb{E}_{X\sim\bar{q}_{X}}[\mathrm{Var}(q_{Y|X})]+\mathbb{E}_{X\sim\bar{q}_{X}}[\mathrm{Fit}(\bar{q}_{Y|X},p_{Y|X})]$
The calculation method for $\bar{q}$ is outlined in the following theorem, and this method is equally applicable to $\bar{q}_{Y|x}$.
Theorem 4.2
With the provision of an i.i.d. sample set $s^{n}$ and a prior distribution $\mathcal{P}$, we have
$\bar{q}=\sum_{q_{i}\in\Delta}\mathcal{Q}(q_{i})q_{i},$ (14)
$\mathcal{Q}(q_{0})=\Pr(s^{n},q_{0})/\sum_{q_{i}\in\Delta}\Pr(s^{n},q_{i}),$
where $n(z)=\sum_{i=1}^{n}\mathbf{1}_{\{z\}}(z^{(i)})$, $\Pr(s^{n},q_{i})=\mathcal{P}(q_{i})n!\prod_{z\in\mathcal{Z}}\frac{q_{i}(z)^{n(z)}}{n(z)!}$.
In the absence of further information regarding the i.i.d. sample $s^{n}$ for $q$, we invoke the equal-probability hypothesis, according to which $\mathcal{P}$ is assumed to be the uniform distribution across the space $\Delta$. The equal-probability hypothesis is one of the fundamental assumptions of statistical physics, positing that all probabilities are uniformly distributed over the space $\Delta$. Assuming the equal-probability hypothesis holds, it follows that
$\mathcal{Q}(q_{0})=(n!\prod_{z\in\mathcal{Z}}\frac{q_{0}(z)^{n(z)}}{n(z)!})/\sum_{q_{i}\in\Delta}(n!\prod_{z\in\mathcal{Z}}\frac{q_{i}(z)^{n(z)}}{n(z)!}).$ (15)
Based on Theorem 4.1, optimizing the learning error is equivalent to optimizing the expected fitting error of the model for $\bar{q}$ and $\bar{q}_{Y|x}$ under $X\sim\bar{q}_{X}$, as shown in the following formula:
$\min_{\theta}\mathrm{Fit}(q,p),$ (16)
$\min_{\theta}\mathbb{E}_{X\sim\bar{q}_{X}}\mathrm{Fit}(\bar{q}_{Y|X},p_{Y|X}).$
Clearly, $p=\bar{q}$, $p_{Y|x}=\bar{q}_{Y|x}$, and $p_{X}=\bar{q}_{X}$ are the optimal solutions for the optimization objectives. However, when the prior distribution $\mathcal{P}$ is unknown or the equal-probability hypothesis clearly does not hold, Theorem 4.2 is not applicable. Therefore, in practice, we use simpler unbiased and consistent estimators based on $s^{n}$ as substitutes, as shown below:
$\min_{\theta}\mathrm{Fit}(\tilde{q},p)+\lambda\mathrm{Est}(\tilde{q}),$ (17)
$\min_{\theta}\mathbb{E}_{X\sim\tilde{q}_{X}}[\mathrm{Fit}(\tilde{q}_{Y|X},p_{Y|X})+\lambda\mathrm{Est}(\tilde{q}_{Y|X})].$
Thus, the model’s learning error comprises three parts: 1. Uncertainty that depends solely on the data, 2. The fitting error of the model for $\tilde{q}$ and $\tilde{q}_{Y|x}$, and 3. The estimation error generated by $\tilde{q}$, $\tilde{q}_{X}$ and $\tilde{q}_{Y|x}$.
4.2 Standard loss function
In this subsection, the equivalence between standard losses and the Fenchel-Young loss is established. Common loss functions are all standard loss functions, which also demonstrates the generality of the concept of standard loss functions.
Theorem 4.3
Given any two vectors $\mu,\nu$ and a standard loss $\ell(\mu,\nu)$, there exists a strictly convex function $\Omega$ such that $\ell(\mu,\nu)-\ell(\mu,\mu)=d_{\Omega}(\mu,\nu_{\Omega}^{*})$, where $d_{\Omega}(\mu,\nu_{\Omega}^{*})$ denotes the Fenchel-Young loss.
In our study, we represent the standard loss as $d_{\Omega}(q,f_{\theta})$, where $f_{\theta}$ denotes the model parameterized by $\theta$. To maintain generality without compromising clarity, we assume that $d_{\Omega}(q,f_{\theta})$ is continuous and differentiable with respect to $\theta$. The standard loss function is a broad concept that encompasses a wide range of commonly used loss functions in statistics and machine learning, as illustrated in Table 2.
Table 2: Examples of common loss functions and their corresponding standard loss forms [136]
Loss $\mathrm{dom}(\Omega)$ $\Omega(\mu)$ $\hat{y}_{\Omega}(\theta)$ $d_{\Omega}(\theta;y)$
Squared $\mathbb{R}^{d}$ $\frac{1}{2}\|\mu\|^{2}$ $\theta$ $\frac{1}{2}\|y-\theta\|^{2}$
Perceptron $\Delta^{|\mathcal{Y}|}$ 00 $\arg\max(\theta)$ $\max_{i}\theta_{i}-\theta_{k}$
Logistic $\Delta^{|\mathcal{Y}|}$ $-H(\mu)$ $\mathrm{softmax}(\theta)$ $\log\sum_{i}\exp\theta_{i}-\theta_{k}$
Hinge $\Delta^{|\mathcal{Y}|}$ $\langle{\mu},{\mathbf{e}_{k}-\mathbf{1}}\rangle$ $\arg\max(\mathbf{1}-\mathbf{e}_{k}+\theta)$ $\max_{i}~{}[[i\neq k]]+\theta_{i}-\theta_{k}$
Sparsemax $\Delta^{|\mathcal{Y}|}$ $\frac{1}{2}\|\mu\|^{2}$ $\mathrm{sparsemax}(\theta)$ $\frac{1}{2}\|y-\theta\|^{2}-\frac{1}{2}\|\hat{y}_{\Omega}(\theta)-\theta\|^{2}$
Logistic (one-vs-all) $[0,1]^{|\mathcal{Y}|}$ $-\sum_{i}H([\mu_{i},1-\mu_{i}])$ $\mathrm{sigmoid}(\theta)$ $\sum_{i}\log(1+\exp(-(2y_{i}-1)\theta_{i}))$
  • 1
    $\hat{y}_{\Omega}(\theta)\in\arg\max_{p\in\mathrm{dom}(\Omega)}\langle\theta,p\rangle-\Omega(p)$.
  • 2
    For multi-class classification, we assume that $\mathcal{Y}=\{\mathbf{e}_{i}\}_{i=1}^{d}$ and the ground-truth is $y=\mathbf{e}_{k}$, where $\mathbf{e}_{i}$ denotes a standard basis (”one-hot”) vector.
  • 3
    We denote the Shannon entropy by $H(p):=-\sum_{i}p_{i}\log p_{i}$, where $p\in\Delta^{\mathcal{Y}}$.
4.3 GSC algorithm
This subsection demonstrates the specific implementation method of the GSC algorithm when using standard loss function.
The GSC algorithm aims to simultaneously reduce both the gradient norm and the structural error. Gradient-based optimization methods, including SGD, Batch-GD, and Newton’s method, primarily focus on minimizing the squared $L_{2}$ norm of the gradient, $\|\nabla_{\theta}d(\tilde{q},f)\|_{2}^{2}$. Alternative methods, such as simulated annealing and genetic algorithms, can also be employed as complementary strategies to achieve this objective. As for the optimization of $S(\alpha,\beta,\gamma,A)$, we establish the following reasonable assumption as the foundation for our analysis on the impact of parameter count and model depth.
Definition 4.1 (Uniform Sampling Assumption)
Each column of $\nabla_{\theta}f_{x}$ is approximated as uniformly sampled from a ball $\mathcal{B}^{|\theta|}_{\epsilon}:\{x\in\mathbb{R}^{\theta},\|x\|_{2}\leq\epsilon\}$.
This assumption essentially requires that the derivative values of the loss function with respect to the parameters are finite and remain approximately independent of each other. It should be noted that the aforementioned assumption is not inherently and unconditionally valid. In section 5, we will explain how numerous existing deep learning techniques contribute to the substantiation of these assumptions. Based on the established Theorem 4.5, we deduce the following theorem to provide a theoretical basis for reducing structural error.
Theorem 4.4
  1. 1.
    If $|\theta|<|\tilde{q}|$, we have
    $\lambda_{\min}(A)$ $=0.$ (18)
  2. 2.
    If $A=aI$, it follows that:
    $\lambda_{min}(A)=\lambda_{max}(A).$ (19)
  3. 3.
    If the model satisfies the uniform sampling assumption 4.1 and $\tilde{q}\leq|\theta|-1$, then with probability at least $1-O(1/|\tilde{q}|)$ the following holds:
    $\lambda_{\max}(A)$ $\geq(1-\frac{2\log|\tilde{q}|}{|\theta|})^{2}\epsilon^{2},$ (20)
    $\lambda_{\max}(A)-$ $\lambda_{\min}(A)\leq\frac{2|\tilde{q}|\sqrt{6\log|\tilde{q}|}\epsilon^{2}}{\sqrt{|\theta|-1}}.$
The present theorem motivates the implementation of several strategies to regulate the structural error, thereby facilitating the convergence of the model to the optimal price. These strategies encompass:
  1. 1.
    Mitigating the dependencies among parameters to facilitate the validity of the uniform sampling assumption;
  2. 2.
    Increasing the number of the model parameters when the model aligns with the uniform sampling assumption.
  3. 3.
    Utilizing a specialized initialization method or model architecture to ensure that $G(A)\approx 0$ ($A\approx aI$), and this approximation is maintained throughout the training process.
In section 5, we will elucidate that numerous existing deep learning techniques essentially correspond to the above strategies.
4.4 Fitting error optimization
This subsection demonstrates that under the conditions specified in 4.1, the model can converge to the global optimal solution of the fitting error minimization problem as follows:
$\min_{\theta}\mathrm{Fit}(\tilde{q},p),$ (21)
$\min_{\theta}\mathbb{E}_{X\sim\tilde{q}_{X}}\mathrm{Fit}(\tilde{q}_{Y|X},p_{Y|X}).$
Specifically, we provide upper and lower bounds for the fitting error based on the gradient norm of the loss function and the structural error through the following theorem and corollary. These upper and lower bounds can converge to zero under the conditions specified in 4.1, thereby ensuring that the optimization objective (LABEL:eq:fitting_obj) can achieve the global optimal solution.
Theorem 4.5
If $\lambda_{\min}(A)\neq 0$, we have
$\frac{\|\nabla_{\theta}d_{\Omega}(\tilde{q},f)\|_{2}^{2}}{\lambda_{\max}(A)}\leq\mathrm{Fit}^{2}(\tilde{q},p)\leq\frac{\|\nabla_{\theta}d_{\Omega}(\tilde{q},f)\|_{2}^{2}}{\lambda_{\min}(A)}.$ (22)
This theorem leads to the following significant conclusions:
  1. 1.
    Since $\mathrm{Fit}(\tilde{q},p)$ is equivalent to $\|\nabla_{\theta}d_{\Omega}(\tilde{q},f)\|_{2}$ when $\lambda_{\min}(A)\neq 0$, in the context of the aforementioned non-convex optimization problem LABEL:eq:fitting_obj, the global optima can be approximated by concurrently minimizing $S(\alpha,\beta,\gamma,A)$ and $\|\nabla_{\theta}d_{\Omega}(\tilde{q},f)\|$.
  2. 2.
    The gradient-based/sub-gradient-based iterative algorithms, such as GD and SGD, allow the model parameters to converge to stationary points, where $\|\nabla_{\theta}d_{\Omega}(\tilde{q},f)\|_{2}=0$.
  3. 3.
    The above theorem also provides the condition under which the gradient-based iterative algorithms can converge to a globally optimal solution, i.e., $S(\alpha,\beta,\gamma,A)=0$.
Based on the established Theorem 4.5, we can derive the upper bound on the fitting error in conditional PD learning.
Corollary 4.1
If $\lambda_{\min}(A_{x})\neq 0$, we have
$\frac{\|\nabla_{\theta}d_{\Omega}(\tilde{q}_{Y|x},f_{x})\|_{2}^{2}}{\lambda_{\max}(A_{x})}\leq\mathrm{Fit}^{2}(\tilde{q}_{Y|x},p_{Y|x})\leq\frac{\|\nabla_{\theta}d_{\Omega}(\tilde{q}_{Y|x},f_{x})\|_{2}^{2}}{\lambda_{\min}(A_{x})}.$ (23)
Similarly, in conditional PD learning, the gradient-based iterative algorithms are able to converge to a globally optimal solution, provided that the following condition is satisfied:
$S(\alpha,\beta,\gamma,A_{x})\approx 0,\forall x\in\mathcal{X}.$ (24)
4.5 Estimation error and uncertainty
This subsection examines the properties of the estimation error and uncertainty. Since both $\tilde{q}$ and $\hat{q}$ serve as unbiased and consistent estimators of $\bar{q}$, we can derive the following theorem.
Theorem 4.6
$\lim_{n\to\infty}\hat{q}=\tilde{q}=\bar{q},$ (25)
$\lim_{n\to\infty}\mathrm{Var}(q)=0,$
$\lim_{n\to\infty}\mathrm{Est}(\tilde{q})=0.$
This theorem suggests that, given a sufficiently large sample size, the estimation error associated with an arbitrary unbiased and consistent estimator is negligible. It is evident that this theorem also holds true in the context of conditional PD learning. Beyond the sample set, the prior distribution $\mathcal{P}$ also exerts a significant influence on the estimation error. For a typical deterministic single-label classification with negligible noise, the label probability distribution conditioned on feature $x$ is given by $q_{Y|x}=\delta_{y}$, where $y$ is the label associated with $x$. Under these conditions, it follows that $\forall x\in\mathcal{X}$, $\mathrm{Est}(\tilde{q}_{Y|x})=0$ and $\mathrm{Var}(q_{Y|x})=0$.
4.6 Bounds of learning error
This subsection derives two types of bounds on the learning error: model-dependent bounds and model-independent bounds. Utilizing the triangle inequality of norms, the upper and lower bounds for the learning error, which encompasses both fitting and estimation errors, can be derived as follows.
Theorem 4.7
Model-dependent bounds in native PD learning:
$\mathrm{Var}(q)+\left(\sqrt{\mathrm{Fit}(\tilde{q},p)}-\sqrt{\mathrm{Est}(\tilde{q})}\right)^{2}$ (26)
$\leq L_{n}(q,p)$
$\leq\mathrm{Var}(q)+\left(\sqrt{\mathrm{Fit}(\tilde{q},p)}+\sqrt{\mathrm{Est}(\tilde{q})}\right)^{2},$
$\frac{\|\nabla_{\theta}d_{\Omega}(\tilde{q},f)\|_{2}^{2}}{\lambda_{\max}(A)}\leq\mathrm{Fit}^{2}(\tilde{q},p)\leq\frac{\|\nabla_{\theta}d_{\Omega}(\tilde{q},f)\|_{2}^{2}}{\lambda_{\min}(A)},$
where $\lambda_{\min}(A)\neq 0$.
Model-dependent bounds in conditional PD learning:
$\mathbb{E}_{X\sim\bar{q}_{X}}\mathrm{Var}(q_{Y|x})+\mathbb{E}_{X\sim\bar{q}_{X}}\left(\sqrt{\mathrm{Fit}(\tilde{q}_{Y|x},p_{Y|x})}-\sqrt{\mathrm{Est}(\tilde{q}_{Y|x})}\right)^{2}$ (27)
$\leq L_{c}(q,p)$
$\leq\mathbb{E}_{X\sim\bar{q}_{X}}\mathrm{Var}(q_{Y|x})+\mathbb{E}_{X\sim\bar{q}_{X}}\left(\sqrt{\mathrm{Fit}(\tilde{q}_{Y|x},p_{Y|x})}+\sqrt{\mathrm{Est}(\tilde{q}_{Y|x})}\right)^{2},$
$\frac{\|\nabla_{\theta}d_{\Omega}(\tilde{q}_{Y|x},f_{x})\|_{2}^{2}}{\lambda_{\max}(A_{x})}\leq\mathrm{Fit}^{2}(\tilde{q}_{Y|x},p)\leq\frac{\|\nabla_{\theta}d_{\Omega}(\tilde{q}_{Y|x},f_{x})\|_{2}^{2}}{\lambda_{\min}(A_{x})},$
where $\lambda_{\min}(A_{x})\neq 0,\forall x\in\mathcal{X}$.
The preceding theorems indicate that the learning error is bounded by the estimation error and the fitting error with respect to the estimators $\tilde{q}$, $\tilde{q}_{X}$ and $\tilde{q}_{Y|x}$. Therefore, to minimize the learning error, it is essential to reduce the estimation error and adjust the model to align with $\bar{q}$ and $\bar{q}_{Y|x}$ by fitting $\tilde{q}$ and $\tilde{q}_{Y|x}$.
For the learning error in conditional PD learning, we obtain the following model-independent upper and lower bounds.
Theorem 4.8
When there exists $x\in\mathcal{X}$ such that $(\bar{q}_{Y})_{\Omega}^{*}\in\mathcal{F}_{\Theta,x}$, we have
$\mathbb{E}_{X\sim\bar{q}_{X}}[\mathrm{Var}(q_{Y|X})]$ $\leq\min_{\theta}L_{c}(q,p)\leq I(X;Y),$ (28)
where $(X,Y)\sim\bar{q}_{XY}$.
Based on this theorem, we can draw the following two corollaries.
Corollary 4.2
  1. 1.
    Given limited sampling data, the target distribution cannot be uniquely determined, and learning error is always present, with the lower bound being the expectation of uncertainty.
  2. 2.
    The greater the mutual information between features and labels, the higher the potential generalization error.
4.7 Algorithmic logic of PD learning
In this subsection, we provide a comprehensive description of the PD learning framework and summarize the algorithm logic and components in PD learning. In addition, we compare PD learning with the traditional deep learning framework to clarify their similarities and differences, highlighting the advantages inherent in the PD learning framework.
Figure 1: Algorithm derivation logic diagram of PD learnig.
We summarize the algorithm logic of native PD learning and conditional PD learning in Figure 1, where the white box denotes the intermediate quantity of reasoning, whereas the gray box signifies the final processable components. A central aspect of the estimation error is the selection of an unbiased and consistent estimator $\tilde{q}$, which serves as a bridge between the optimal estimator $\bar{q}$ and the model output $p$. Although $\hat{q}$ is theoretically feasible, in practical scenarios, especially in the case of insufficient samples, $\hat{q}$ may not be the best choice. Based on the prior knowledge, we can generate more accurate estimators. If the underlying true distribution is known to be smooth, we can construct $\tilde{q}$ by applying Laplacian smoothing to $\hat{q}$, which often yields smaller estimation error and superior performance. In the field of machine learning, techniques such as label smoothing [141] and synthetic minority oversampling technique (SMOTE) [142], which introduce a degree of smoothness to the existing data, also yield favorable outcomes. After determining $\tilde{q}$, the subsequent step is to approximate it with the model by minimizing the fitting error $\mathrm{Fit}(\tilde{q},p)$. Optimization of the fitting error, given by $\arg\min_{\theta}\mathrm{Fit}(\tilde{q},p)$, is a typical non-convex problem. Simple gradient-based optimization algorithms do not guarantee a global optimal solution. By using the standard loss function, the GSC algorithm facilitates the approximation of global optimal solutions in fitting error minimization by minimizing both $\|\nabla_{\theta}d_{\Omega}(\tilde{q},f)\|$ and the structural error. Ultimately, based on the aforementioned algorithm design, the upper and lower bounds on the learning error in the native PD learning and conditional PD learning are derived separately.
Corresponding to the algorithm logic above, the PD learning framework encompasses five principal components: standard loss, model, prior knowledge ($\mathcal{P}$), estimator ($\tilde{q}$), and the GSC algorithm (represented by $\mathcal{A}$), as shown in Figure 2. The estimator yields an approximation of $\bar{q}$, denoted as $\tilde{q}$, leveraging both the sample set $s^{n}$ and prior knowledge $\mathcal{P}$. The standard loss function proposed is a type of loss function that possesses a unique optimal solution and sampling stability. We have demonstrated that under these conditions, it has a unique mathematical expression and can be represented as a Young–Fenchel loss function, which encompasses various common losses. The model in PD learning is utilized to fit the target distribution $\tilde{q}$ via its Legendre-Fenchel conjugate function. The GSC algorithm is designed to minimize both the gradient norm and the structural error. Within the PD learning framework, we demonstrate that the global optimal solution to non-convex optimization problems, such as fitting error minimization, can be approached using the GSC algorithm.
Figure 2: Basic components of PD learning.
To elucidate PD learning more effectively, we compare it with the conventional deep learning theory framework.
  • 1.
    Learning objective. The underlying true distribution $q$ or $q_{Y|x}$ represents the learning objectives in PD learning. Within the conventional deep learning theoretical framework, the primary objective is to ascertain the unknown mapping between features and labels.
  • 2.
    Modeling method. In PD learning, the learning objective is conceptualized as a random variable within the probability simplex. Conversely, within the conventional deep learning theoretical framework, the unknown mapping is represented as a function that processes input features to generate corresponding labels within a defined functional space $\mathcal{F}_{\Theta}$.
  • 3.
    Optimization objective and its reformulation. The optimization objective in PD learning is the minimization of learning error, as defined in (9). This objective is achieved by controlling the estimation error and the fitting error. In contrast, within the conventional deep learning theoretical framework, the optimization objective is the minimization of expected risk. Expected risk minimization is reformulated as Empirical Risk Minimization (ERM) and generalization error minimization, collectively referred to as Structural Risk Minimization (SRM).
  • 4.
    Fitting method of model. In PD learning, the target is fitted by employing the Legendre-Fenchel conjugate function of the model’s output. Conversely, in the conventional deep learning theoretical framework, neural networks are typically used to directly fit the target without the intermediary step of the conjugate function.
  • 5.
    Loss function and its role. In PD learning, the standard loss function is used to facilitate the minimization of the gradient norm. In contrast, within the conventional deep learning theoretical framework, a variety of loss functions may be employed. The minimization of the loss function, in conjunction with the regularization term corresponding to SRM, serves as the optimization objective.
  • 6.
    Optimization algorithm. The GSC algorithm is used to optimize the fitting error, which can ensure obtaining an approximate global optimal solution for fitting error minimization in the PD learning framework. Within the conventional deep learning theoretical framework, gradient-based iterative algorithms, including GD, SGD, and Adaptive Moment Estimation(Adam), are utilized to achieve SRM. While these algorithms strive to identify the global optimal solution, they do not ensure its attainment.
5 Application to Deep Learning
This section analyzes various unresolved issues in deep learning from the perspective of distribution learning.
5.1 Classic machine learning paradigms from the perspective of PD learning
In this subsection, we demonstrate that PD learning can be transformed into a classic machine learning paradigm, specifically the minimization of the sum of a loss function and an $L_{2}$ regularization term.
Theorem 5.1
Given that when $\theta=\mathbf{0}$, $\forall x\in\mathcal{X}$, $p_{Y|x}$ is the uniform distribution over $|\mathcal{Y}|$, it follows that
  • 1.
    There exists $k>0$ such that
    $\|p_{Y|x}\|_{2}^{2}-\min_{\theta\in\Theta^{\prime}}\|p_{Y|x}\|_{2}^{2}\leq k\|\theta\|_{2}^{2}.$ (29)
  • 2.
    $\min_{\theta\in\Theta^{\prime}}L_{c}(q,p)$ is equivalent to
    $\min_{\theta\in\Theta^{\prime}}\left\{\mathbb{E}_{X}\langle\bar{q}_{Y|X},-(f_{X})_{\Omega}^{*}\rangle+\lambda\|\theta\|^{2}_{2}\right\},$
    where $\lambda>0$, $\Theta^{\prime}=\{\theta|\nabla^{2}_{\theta}\|(f_{x})_{\Omega^{*}}^{*}\|_{2}^{2}\succeq 0\}$.
The proof can be found in A.8.
The condition that $f_{\mathcal{Y}|\mathbf{0},x}=\mathbf{0}$ in Theorem 5.1 holds true in the vast majority of DNNs, thereby endowing the theorem with broad applicability. According to Theorem 5.1, $\theta=\mathbf{0}$ corresponds to the uniform distribution when $\theta\in\Theta^{\prime}$. Therefore, $\|\theta\|_{2}^{2}$ is related to the uniformity of the model’s output distribution. Larger values of this norm imply a lower degree of uniformity in the output distribution $p_{Y|x}$.
On the other hand, Softmax Cross-entropy is a typical standard loss function. From the perspective of distribution learning, when Softmax Cross-entropy is employed as the loss function in supervised classification tasks using DNNs, the models ultimately learn to fit the label distribution under the given features of the training set. Moreover, the convergence of model parameters to the global optimum can be theoretically guaranteed through the PD learning framework.
5.2 Non-Convex optimization in Deep learning from the perspective of PD learning
In this subsection, we analyze and discuss various deep learning techniques within the PD learning framework. We propose that many deep learning techniques can be interpreted from the perspective of the three GSC algorithm implementations discussed in subsection 4.3. These implementations involve enhancing gradient independence to achieve the Uniform Sampling Assumption and reducing the structural error by increasing the number of parameters upon fulfilling the Uniform Sampling Assumption.
5.3 Enhancing the substantiation of the uniform sampling assumption
During model training process, numerous techniques deliberately or inadvertently introduce randomness to reduce the dependency among parameters and the enhance the independence among the columns of $\nabla_{\theta}f_{x}$. Consequently, we can interpret some existing deep learning techniques from the perspective of fostering the fulfillment of the Uniform Sampling Assumption.
Random initialization
During the parameter initialization phase, the random initialization of parameters in DNNs [143, 144] ensures that the gradient $\nabla_{\theta}f_{x}$ for an untrained model is randomly assigned and helps to reduce the dependence among the elements of $\nabla_{\theta}f_{x}$. Thus, the random initialization of parameters allows the columns of the gradient matrix $\nabla_{\theta}f_{x}$ to be approximately uniformly sampled from a ball.
Gradient-based optimization
During the model training, randomly selecting samples in each iteration of SGD or Batch Gradient Descent (Batch-GD) leads to a random update of gradient direction, which enhances independence between the columns of $\nabla_{\theta}f_{x}$. This, in turn, enhances their mutual orthogonality according to Lemma 3.4. Therefore, we believe that the stochasticity introduced by SGD facilitates the validity of the uniform sampling assumption, thereby enabling the structural error to be controlled by the model’s parameter size.
Depth of model
From a model structure perspective, if the model possesses an extremely large number of layers, it is justifiable to assume that the majority of the gradients exhibit a high degree of approximate independence. As the number of layers between parameters $\theta_{i}$ and $\theta_{j}$ increases, the dependence between $\nabla_{\theta_{i}}f$ and $\nabla_{\theta_{j}}f$ weakens [145]. Thus, an increase in model depth helps to make the Uniform Sampling Assumption 4.1 hold.
Dropout
Dropout [146] ensures that the loss remains relatively constant even when specific nodes in the DNN are temporarily discarded. This effectively reduces the influence of these nodes on other parameters, thereby weakening the dependence between $\nabla_{\theta}f_{x}$ and $\nabla_{\theta}f_{x}$ and improving the rationality of the Uniform Sampling Assumption 4.1.
5.4 Reducing structural error by increasing parameter count
over-parameterization
As analyzed above, the applications of parameter random initialization, stochasticity inherent in gradient-based iterative algorithms, parameter and feature normalization, and dropout, contribute to validating the Uniform Sampling Assumption 4.1 and reducing the sampling ball radius . Based on the conclusion of Theorem 4.4, it can be deduced that the following conclusion holds with probability $1-O(1/|\mathcal{Y}|)$:
$\lambda_{\max}(A_{x})$ $\geq(1-t^{2})^{2}\epsilon^{2},$ (30)
$\lambda_{\max}(A_{x})$ $-\lambda_{\min}(A_{x})\leq 2|\mathcal{Y}|\sqrt{3}t\epsilon^{2},$
where $t=\frac{2\log|\mathcal{Y}|}{|\theta|-1}$. Therefore, increasing the number of parameters $|\theta|$ can effectively manage the upper bound of $G(A_{x})$, and be conducive to reducing the value of $L(A_{x})$. In addition, when $|\theta|\gg 4\epsilon^{2}\log|\mathcal{Y}|$, it follows that $G(A_{x})\approx 0$. According to Theorem 4.7, in this case the fitting error is proportional to the norm of gradients with the rate being $L(A_{x})$. Thus increasing the number of parameters $|\theta|$ while decreasing $\|\nabla_{\theta}(f_{x})_{i}\|$ can reduce both the structural error and fitting error. This underlies the ability of the overparameterized model to approximate global optimal solutions for non-convex optimization through a gradient-based optimization algorithm.
6 Experiments and analysis of results
In this section, we aim to validate GSC algorithm through three experiments. Firstly, we examine the variations in structural error, gradient norm, and fitting error during the training process of the model to confirm that the deep model’s training adheres to the GSC algorithm’s description. Then, we analyze the convergence curves of models with varying structures as the number of layers increases, discussing the impact of model depth on convergence speed and results.
In our experiments, we utilize the MNIST database [147] as a case study to visualize how the structural error, gradient norm and fitting error evolve as we increase the network depth across different model architectures. This database has 60k training images and 10k testing images in 10 categories, containing 10 handwritten digits. The model architectures and configuration parameters in the experiment are depicted in Figures 3.
Figure 3: Model architectures and configuration parameters.
The experiments were executed using the following computing environment: Python 3.7, Pytorch 2.2.2, and a GeForce RTX2080Ti GPU.
Table 3: The configuration of Training.
loss function cross-entropy loss #epochs 1
batch size 64 learning rate 0.01
momentum 0.9 data size 28*28
Initially, we set $k=1$ for both Model A and Model B and employed the batch-GD algorithm to train the two models. The training parameters are detailed in Table 3, and the changes in key indicators during the training process are illustrated in Figures 4. The fitting errors of the models and their respective bounds are depicted in Figure 5.
Figure 4: Changes in indicators during the training process.
Figure 5: Changes of bounds during the training process.
Figure 6: Model c with increasing layers
Figure 7: Model d with increasing blocks
Based on Figure 4 and Figure 5, we can obtain the following conclusions:
  1. 1.
    As the training progresses, $U(A_{x})$, $L(A_{x})$ and $G(A_{x})$, apart from a brief initial increase, generally exhibit a stable downward trend. In contrast, the gradient norm $\|\nabla_{\theta}d_{\Omega}(\tilde{q}_{Y|x},f_{x})\|_{2}^{2}$ initially rises and then undergoes sharp oscillations before decreasing towards zero. Even when the training reaches a stable state, the gradient norm still exhibits significant fluctuations.
    The diminution of structural error throughout the progression of gradient iterations demonstrates that, as we explained in Paragraph 5.3, the stochasticity inherent in the SGD algorithm is conducive to reducing structural error.
  2. 2.
    As the training progresses, we observe that the variation in fitting error increasingly aligns with the trends of its upper and lower bounds. Concurrently, while the bounds themselves decrease, the gap between them narrows. This phenomenon validates our proposed GSC algorithm, demonstrating that the minimization of the fitting error is achieved by controlling its upper and lower bounds. This also empirically validates the validity of the bounds of fitting error in Theorem 4.7.
  3. 3.
    Upon comparing the fitting errors, gradient norms, and structural errors of various models, it becomes evident that Models c and d, with a larger number of parameters, exhibit superior fitting performance. Furthermore, Models a and b, which have fewer parameters, display significant discrepancies between their fitting error bounds and the actual fitting errors. In contrast, the fitting errors of Models c and d are nearly identical to their respective upper and lower bounds. Consequently, this finding aligns with our theoretical conclusions regarding over-parameterization.
The experimental results regarding the impact of model depth on fitting error optimization are illustrated in Figure 6 and Figure 7. As the number of layers $k$ increases from 0 to 3, the model’s convergence to better parameters is facilitated, resulting in a stable reduction in fitting error. However, beyond $k=3$, further increases in layers do not yield positive effects; Model d does not show significant improvements in convergence speed or final values, while Model c experiences gradient explosion at $k=5$, and the fitting error does not converge to a smaller value. This observation is consistent with our expectations: increasing the number of layers in a model can enhance its fitting capability, but the training difficulty of models tends to increase with their depth. Compared to Model c, Model d, which incorporates residual blocks, exhibits more stable convergence. Nevertheless, at the same number of layers, the fitting error values attained by Model d are higher than those of Model c, a phenomenon worthy of attention.
7 Conclusion
This study introduces a novel theoretical learning framework, PD learning, which is characterized by its focus on the true underlying distribution. In order to learning the underlying distribution, we introduce the learning error as the optimization objective. To optimize the learning error, this paper proposes the necessary conditions for loss functions, models, and optimization algorithms. Based on these conditions, the non-convex optimization mechanism corresponding to model training can be theoretically resolved. We derive both model-independent and model-dependent upper bounds for the learning error, providing new insights into understanding machine learning. We illustrate that PD learning provides clearer insights into the unresolved issues in deep learning. The consistency of our framework with existing theories and experimental results, as well as the explanatory power of existing theories and phenomena, demonstrate the correctness and effectiveness of our proposed framework.
Appendix A Appendix
In this section, we prove the results stated in the paper and provide necessary technical details and discussions.
A.1 Proof of Theorem 4.2
Assuming $q=q_{0}$, the frequency distribution $\hat{q}$, inferred from the sample $s^{n}$, adheres to a multinomial distribution: $n\hat{q}\sim\operatorname{multinomial}(n,q_{0})$. With the provision of i.i.d. sample data $s^{n}$ and a prior distribution such that $q\sim\mathcal{P}$, the application of Bayes’s theorem facilitates the derivation of the posterior distribution for $q$.
$\mathcal{Q}(q_{0})$ $=\Pr(q_{0}|s^{n})$ (31)
$=\frac{\Pr(\hat{q}|q_{0})\Pr(q_{0})}{\Pr(s)}$
$=\frac{\Pr(\hat{q}|q_{0})\mathcal{P}(q_{0})}{\Pr(\hat{q})}$
$=\mathcal{P}(q_{0})(n!\prod_{z\in\mathcal{Z}}\frac{q_{0}(z)^{n(z)}}{n(z)!})/\sum_{q_{i}\in\Delta}(\mathcal{P}(q_{i})n!\prod_{z\in\mathcal{Z}}\frac{q_{i}(z)^{n(z)}}{n(z)!}),$
where $n(z)=\sum_{i=1}^{n}\mathbf{1}_{\{z\}}(z^{(i)})$.
A.2 Proof of Theorem 4.3
Since $\ell(\mu,\nu)$ is continuous, differentiable, and has a unique minimum value of zero at $\nu=\mu$, where its gradient with respect to $\nu$ vanishes, i.e., $\nabla_{\nu}\ell(\mu,\mu)=0$, we can express this gradient as
$\nabla_{\nu}\ell(\mu,\nu)=(\nu-\mu)^{\top}g(\mu,\nu).$ (32)
Due to the uniqueness of the minimum value, the expected loss, $\mathbb{E}_{\omega\sim Q}[\ell(\omega,\nu)]$, is strictly convex, and at its minimum $\bar{\omega}$, where $\bar{\omega}=\mathbb{E}_{\omega\sim Q}[\omega]$, the gradient with respect to $\nu$ vanishes, i.e., $\nabla_{\nu}\mathbb{E}_{\omega\sim Q}[\ell(\omega,\bar{\omega})]=0$, which implies that $\sum_{\omega\in\mathcal{Q}}Q(\omega)(\bar{\omega}-\omega)^{\top}g(\omega,\bar{\omega})=0$, where $\omega\sim Q$.
Given that $Q(\omega)$ is an arbitrary distribution function, $g(\omega,\nu)$ must be independent of $\omega$, allowing us to replace $g(\omega,\nu)$ by $g(\nu)$. Since $\mathbb{E}_{\omega\sim Q}[\ell(\omega,\nu)]$ is a strictly convex function, its Hessian with respect to $\nu$ can be derived as:
$\nabla^{2}_{\nu}\mathbb{E}_{\omega\sim Q}[\ell(\omega,\nu)]=\sum_{\omega\in\mathcal{Q}}Q(\omega)[g(\nu)+(\nu-\omega)^{\top}\nabla_{\nu}g(\nu)].$ (33)
Simplifying we get:
$\nabla^{2}_{\nu}\mathbb{E}_{\omega\sim Q}[\ell(\omega,\nu)]=g(\nu)+(\nu-\bar{\omega})\nabla_{\nu}g(\nu)>0.$ (34)
The inequality above is derived from the fact that $\nabla^{2}_{\nu}\mathbb{E}_{\omega\sim Q}[\ell(\omega,\nu)]$ is strictly convex.
When $\nu$ is set to $\bar{\omega}$, we have $\nabla^{2}_{\nu}\mathbb{E}_{\omega\sim Q}[\ell(\omega,\nu)]=g(\bar{\omega})>0$. Since $\bar{\omega}$ can take any value, it follows that for any arbitrary $\nu$, we get $g(\nu)>0$.
Let us define a strictly convex function $\Omega(\nu)$ with its Hessian matrix given by $\nabla^{2}_{\nu}\Omega(\nu)=g(\nu)$. The gradient of the loss function $\ell(\mu,\nu)$ with respect to $\nu$ can be expressed as follows:
$\nabla_{\nu}\ell(\mu,\nu)$ $=\nabla^{2}_{\nu}\Omega(\nu)(\nu-\mu)$ (35)
$=\nabla_{\nu}\nu_{\Omega}^{*}\nu-\nabla_{\nu}\nu_{\Omega}^{*}\mu$
$=\nabla_{\nu}(\langle\nu,\nu_{\Omega}^{*}\rangle-\Omega(\nu))-\mu^{\top}\nabla_{\nu}\nu_{\Omega}^{*}$
$=\nabla_{\nu}(\Omega^{*}(\nu_{\Omega}^{*})-\langle\mu,\nu_{\Omega}^{*}\rangle+h(\mu)).$
Since $\nabla_{\nu}\ell(\mu,\mu)=0$, we find that $h(\mu)=\Omega^{*}(\mu_{\Omega}^{*})-\langle\mu,\mu_{\Omega}^{*}\rangle+c=\Omega(\mu)+c$, where $c$ is a constant. This leads to the following expression for the loss function:
$\ell(\mu,\nu)=\Omega^{*}(\nu_{\Omega}^{*})-\langle\mu,\nu_{\Omega}^{*}\rangle+\Omega(\mu)+c,$ (36)
$\ell(\mu,\mu)=c.$
A.3 Proof of Theorem 4.4
When $|\theta|<|\tilde{q}|$, the matrix $A$ is underdetermined, resulting in $\lambda_{\min}(A)=0$.
Since each column of $\nabla_{\theta}f$ is uniformly from the ball $\mathcal{B}^{|\theta|}_{\epsilon}:\{x\in\mathbb{R}^{|\theta|},\|x\|_{2}\leq\epsilon\}$, according to Theorem 3.4, then with probability $1-O(1/|\tilde{q}|)$ the following holds:
$A_{ii}=\left\|\nabla_{\theta}f_{i}\right\|^{2}_{2}\geq(1-\frac{2\log|\tilde{q}|}{|\theta|})^{2}\epsilon^{2},\text{ for }i=1,\ldots,|h|,$ (37)
$|A_{ij}|=|\langle\nabla_{\theta}f_{i},\nabla_{\theta}f_{j}\rangle|\leq\frac{\sqrt{6\log|\tilde{q}|}}{\sqrt{|\theta|-1}}\epsilon^{2}\quad\text{ for all }i,j=1,\ldots,|h|,i\neq j.$
Because $A=\nabla_{\theta}f^{\top}\nabla_{\theta}f$ is symmetric positive semidefinite matrix, then we have $\lambda_{\min}(A)\leq A_{ii}\leq\lambda_{\max}(A)$. Then the following holds with with probability $1-O(1/|\tilde{q}|)$ at least
$\lambda_{\max}(A)\geq(1-\frac{2\log|\tilde{q}|}{|\theta|})^{2}\epsilon^{2}.$ (38)
According to Lemma 3.5, there exist $k,k^{\prime}\in\{1,\cdots,|\tilde{q}|\}$
$\lambda_{\max}(A)-A_{kk}$ $\leq R_{k},$ (39)
$A_{k^{\prime}k^{\prime}}-\lambda_{\min}(A)$ $\leq R_{k^{\prime}},$
where $R_{k}=\sum_{1\leq j\leq n,j\neq k}|A_{kj}|$, $R_{k^{\prime}}=\sum_{1\leq j\leq n,j\neq k^{\prime}}|A_{k^{\prime}j}|$. Based on the inequalities (37), we obtain
$R_{k}+R_{k^{\prime}}\leq 2(|\tilde{q}|-1)\frac{\sqrt{6\log|\tilde{q}|}}{\sqrt{|\theta|-1}}\epsilon^{2}.$ (40)
Since $A_{kk}\leq\epsilon^{2}$ and $|\tilde{q}|\leq|\theta|-1$, then we have
$\lambda_{\max}(A)-\lambda_{\min}(A)$ $\leq R_{k}+R_{k^{\prime}}+A_{kk}-A_{k^{\prime}k^{\prime}}$ (41)
$\leq R_{k}+R_{k^{\prime}}+\epsilon^{2}-A_{k^{\prime}k^{\prime}}$
$\leq 2(|h|-1)\frac{\sqrt{6\log|\tilde{q}|}}{\sqrt{|\theta|-1}}\epsilon^{2}+\epsilon^{2}-(1-\frac{2\log|\tilde{q}|}{|\theta|})^{2}\epsilon^{2}$
$\leq 2(|\tilde{q}|-1)\epsilon^{2}\frac{\sqrt{6\log|\tilde{q}|}}{\sqrt{|\theta|-1}}+\frac{4\log|\tilde{q}|}{|\theta|}\epsilon^{2}$
$\leq\frac{2|\tilde{q}|\epsilon^{2}\sqrt{6\log|\tilde{q}|}}{\sqrt{|\theta|-1}}.$
A.4 Proof of Theorem 4.5
According to the given definition, we have
$\|\nabla_{\theta}d_{\Omega}(\tilde{q},f)\|_{2}^{2}=e^{\top}Ae,$ (42)
where $A:=\nabla_{\theta}f^{\top}\nabla_{\theta}f$,$e:=\tilde{q}-f_{\Omega}^{*}$. Let $A=UDU^{\top}$ denote the eigenvalue decomposition of $A$. Consequently, $D$ is a diagonal matrix whose entries are the eigenvalues of $A$. We then have
$\min_{\|e\|_{2}=K}e^{\top}Ae$ $=\min_{\|e\|_{2}^{2}=K}e^{\top}(UDU^{\top})e$ (43)
$=\min_{\|e\|_{2}^{2}=K}(U^{\top}e)^{\top}D(U^{\top}e)$
$=\min_{\|U^{\top}e\|_{2}^{2}=K}(U^{\top}e)^{\top}D(U^{\top}e)$
$=K\min_{x\in\mathcal{S}}\sum_{i}x_{i}D_{ii},$
where $\mathbb{S}=\{(x_{1},\dots,x_{N})\in\mathbb{R}^{N}\mid x_{i}\geq 0,~{}~{}\sum_{i}x_{i}=1\}$. The final step is equivalent to
$K\min_{x\in\mathbb{S}}\sum_{i}x_{i}D_{ii}=K\min_{i}D_{ii}=\lambda_{\min}(A)\|e\|_{2}^{2}.$ (44)
Similarly, we obtain
$\max_{\|e\|_{2}=K}e^{\top}Ae=\lambda_{\max}(A)\|e\|_{2}^{2}.$ (45)
It follows that
$\lambda_{\min}(A)\mathrm{Fit}^{2}(\tilde{q},p)\leq\|\nabla_{\theta}d_{\Omega}(\tilde{q},f)\|_{2}^{2}\leq\lambda_{\max}(A)\mathrm{Fit}^{2}(\tilde{q},p).$ (46)
A.5 Proof of Theorem 4.6
According to the Bayesian theorem and Sanov’s theorem 3.2, we have
$\Pr(q_{0}|s)={\Pr(\hat{q}|q_{0})\Pr(q_{0})}/{\Pr(\hat{q})},$ (47)
$\lim_{n\to\infty}-\frac{1}{n}\log\Pr(\hat{q}|q_{0})=D_{KL}(\hat{q}\|q_{0}).$
Therefore, we obtain
$\lim_{n\to\infty}\Pr(q_{0}|s)=\lim_{n\to\infty}e^{-nD_{KL}(\hat{q}\|q_{0})}\Pr(q_{0})/\Pr(\hat{q}).$ (48)
Since $\Pr(q_{0})/\Pr(\hat{q})$ is finite, then we have
$\lim_{n\to\infty}\mathcal{Q}(q_{0})\sim\delta_{\hat{q}}.$ (49)
Since $\hat{q}$ and $\tilde{q}$ are both the unbiased and consistent estimators of $\bar{q}$, it follows that
$\lim_{n\to\infty}\mathrm{Var}(q)=0,$ (50)
$\lim_{n\to\infty}\hat{q}=\tilde{q}=\bar{q},$
$\lim_{n\to\infty}\mathrm{Est}(\tilde{q})=0.$
A.6 Proof of Theorem 4.7
Utilizing the triangle inequality of norms, we obtain
$|\|p-\tilde{q}\|_{2}-\|\tilde{q}-\bar{q}\|_{2}|\leq\|p-\bar{q}\|_{2}\leq\|p-\tilde{q}\|_{2}+\|\tilde{q}-\bar{q}\|_{2}.$ (51)
Substituting the above inequality into equalities 13 completes the proof.
A.7 Proof of Theorem 4.8
Let $v$ be a vector of dimension $n$. According to the definition of the norm, we have
$\|v\|_{1}^{2}$ $=\sum_{i=1}^{n}\sum_{j=1}^{n}|v_{i}||v_{j}|$ (52)
$\geq\sum_{i=1}^{n}|v_{i}|^{2}$
$=\|v\|_{2}^{2}.$
Consequently, we obtain
$\|\bar{q}_{Y|x}-p_{Y|x}\|_{2}^{2}$ $\leq\|\bar{q}_{Y|x}-p_{Y|x}\|_{1}^{2}$ (53)
$\leq 2\ln 2D_{KL}(\bar{q}_{Y|x}||p_{Y|x}),$
where the last inequality is based on Pinsker’s inequality 3.6. Since there exists $x\in\mathcal{X}$ such that $(\bar{q}_{Y})_{\Omega}^{*}\in\mathcal{F}_{\Theta,x}$, it follows that
$\min_{\theta}L_{c}(q,p)$ $\leq\mathbb{E}_{X\sim\bar{q}_{X}}\|\bar{q}_{Y|x}-\bar{q}_{Y}\|_{2}^{2}$ (54)
$\leq\mathbb{E}_{X\sim\bar{q}_{X}}[2\ln 2D_{KL}(\bar{q}_{Y|x}||\bar{q}_{Y})]$
$=2\ln 2I(X;Y),$
where $(X,Y)\sim\bar{q}_{XY}$.
A.8 Proof of Theorem 5.1
Since $\theta=\mathbf{0}$ represents the minimum point of $\|p_{Y|x}\|^{2}_{2}$, where $p_{Y|x}=(f_{\theta,x})_{\Omega^{*}}^{*}$, it follows that $\forall\theta\in\Theta^{\prime}$:
$\nabla_{\theta}\|(f_{\mathbf{0},x})_{\Omega^{*}}^{*}\|_{2}^{2}=0,$ (55)
$\nabla^{2}_{\theta}\|(f_{\theta,x})_{\Omega^{*}}^{*}\|_{2}^{2}\succeq 0.$
By applying the mean value theorem, there exists $\theta^{\prime}=\alpha\mathbf{0}+(1-\alpha)\theta,\alpha\in[0,1]$ such that the following inequalities hold:
$\|(f_{\theta,x})_{\Omega^{*}}^{*}\|_{2}^{2}$ $=\|(f_{\mathbf{0},x})_{\Omega^{*}}^{*}\|_{2}^{2}+\theta^{\top}\nabla^{2}_{\theta}\|(f_{\theta^{\prime},x})_{\Omega^{*}}^{*}\|_{2}^{2}\theta,$ (56)
$\leq 1/|\mathcal{Y}|+\|\theta\|_{2}\|\nabla^{2}_{\theta}\|(f_{\theta^{\prime},x})_{\Omega^{*}}^{*}\|_{2}^{2}\theta\|_{2},$
$\leq 1/|\mathcal{Y}|+\|\theta\|_{2}^{2}\|\nabla^{2}_{\theta}\|(f_{\theta^{\prime},x})_{\Omega^{*}}^{*}\|_{2}^{2},$
where the first inequality is based on the Cauchy-Schwarz inequality, and the second inequality follows from the definition of the matrix-induced 2-norm. Consequently, within the neighborhood of $\theta=\mathbf{0}$, that is $\Theta^{\prime}$, we have
$1/|\mathcal{Y}|\leq\|(f_{\theta,x})_{\Omega^{*}}^{*}\|_{2}^{2}\leq 1/|\mathcal{Y}|+k\|\theta\|_{2}^{2},$ (57)
where $k\geq\|\nabla^{2}_{\theta}\|(f_{\theta^{\prime},x})_{\Omega^{*}}^{*}\|_{2}^{2}$. That is, there exists $k>0$ such that
$\|p_{Y|x}\|_{2}^{2}-\min_{\theta\in\Theta^{\prime}}\|p_{Y|x}\|_{2}^{2}\leq k\|\theta\|_{2}^{2}.$ (58)
It indicates that $\min_{\theta}\|\theta\|_{2}^{2}$ and $\min_{\theta}\|p_{Y|x}\|_{2}^{2}$ are equivalent. Therefore, $\min_{\theta}\mathbb{E}_{X}\|p_{Y|X}\|_{2}^{2}$ is equivalent to minimizing $\min_{\theta}\mathbb{\|\theta\|_{2}^{2}}$. Given the equation:
$L_{c}(q,p)=\mathbb{E}_{X}[\|\bar{q}_{Y|X}\|_{2}^{2}+\|p_{Y|X}\|_{2}^{2}-2\langle\bar{q}_{Y|X},p_{Y|X}\rangle]$
and since $\mathbb{E}_{X}[\|\bar{q}_{Y|X}\|_{2}^{2}]$ is model-independent, it follows that $\min_{\theta}L_{c}(q,p)$ is equivalent to
$\min_{\theta}\{\mathbb{E}_{X}\langle\bar{q}_{Y|x},-(f_{\theta,X})_{\Omega^{*}}^{*}\rangle+\lambda\|\theta\|_{2}^{2}\},$
where $\theta\in\Theta^{\prime}$.
Appendix B Declarations
B.1 Funding
This work was supported in part by the National Key Research and Development Program of China under No. 2022YFA1004700, Shanghai Municipal Science and Technology, China Major Project under grant 2021SHZDZX0100, the National Natural Science Foundation of China under Grant Nos. 62403360, 72171172, 92367101, 62088101, iF open-funds from Xinghuo Eco and China Institute of Communications.
B.2 Competing interests
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
B.3 Ethics approval and consent to participate
Not applicable.
B.4 Data availability
The datasets generated and/or analyzed in the present study are accessible via the GitHub repository at mnist and the website at mnist.
B.5 Materials availability
Not applicable.
B.6 Code availability
Code for preprocessing and analysis is provided in the GitHub repository pd_learning.
B.7 Author contribution
Binchuan Qi: Conceptualization, Methodology, Writing original draft. Wei Gong: Supervision, Review, Declarations. Li Li: Supervision, Review, Declarations.